### v1

# Statistical learnability of smooth boundaries via pairwise binary classification with deep ReLU networks

Abstract
The topic of nonparametric estimation of smooth boundaries is extensively studied in the conventional setting where pairs of single covariate and response variable are observed. However, this traditional setting often suffers from the cost of data collection. Recent years have witnessed the consistent development of learning algorithms for binary classification problems where one can instead observe paired covariates and binary variable representing the statistical relationship between the covariates. In this work, we theoretically study the question of whether multiple smooth boundaries are learnable if the pairwise binary classification setting is considered. We investigate the question with the statistical dependence of paired covariates to develop a learning algorithm using vector-valued functions. The main theorem shows that there is an empirical risk minimization algorithm in a class of deep ReLU networks such that it produces a consistent estimator for indicator functions defined with smooth boundaries. We also discuss how the pairwise binary classification setting is different from the conventional settings, focusing on the structural condition of function classes. As a by-product, we apply the main theorem to a multiclass nonparametric classification problem where the estimation performance is measured by the excess risk in terms of misclassification.
1 Introduction
Let $(\mathcal{X},\mathcal{B}(\mathcal{X}),\mu)$ be the measure space such that $\mathcal{X}=[0,1]^{d_{0}}$, $\mathcal{B}(\mathcal{X})$ is the Borel $\sigma$-algebra, and $\mu$ is the Lebesgue measure, following the common setting in the literature on nonparametric statistics (see, e.g., (Schmidt-Hieber, , 2020; Suzuki, , 2019; Kim et al., , 2021; Bos and Schmidt-Hieber, , 2022; Imaizumi and Fukumizu, , 2022)). Let $\{\mathcal{K}_{1},\cdots,\mathcal{K}_{d_{1}}\}$ be a disjoint partition of $\mathcal{X}$ for some $d_{1}\in\mathbb{N}$, namely
$\{\mathcal{K}_{1},\cdots,\mathcal{K}_{d_{1}}\}\;\textup{ such that }\;\mathcal{K}_{i}\cap\mathcal{K}_{j}=\varnothing\textup{ and }\bigcup_{i=1}^{d_{1}}\mathcal{K}_{i}=\mathcal{X}.$
In the field of nonparametric estimation and statistical learning, estimation or learning of the indicator functions $\mathds{1}_{\mathcal{K}_{1}},\cdots,\mathds{1}_{\mathcal{K}_{d_{1}}}$, which are defined with some additional smoothness condition on $\mathcal{K}_{1},\cdots,\mathcal{K}_{d_{1}}$, is extensively studied in many topics such as set estimation (Mammen and Tsybakov, , 1995), discriminant analysis (Mammen and Tsybakov, , 1999), classification (Tsybakov, , 2004; Tarigan and van de Geer, , 2008; Kim et al., , 2021; Meyer, , 2023; Caragea et al., , 2023), and regression (Imaizumi and Fukumizu, , 2019, 2022). In (Dudley, , 1974), the smoothness of boundaries is defined with Hölder continuous functions expressing sublevel sets (see also (Mammen and Tsybakov, , 1995, 1999; Petersen and Voigtlaender, , 2018; Imaizumi and Fukumizu, , 2022)).
Statistical learnability of smooth boundaries using deep neural networks in the conventional supervised learning setting has recently been studied in (Kim et al., , 2021; Meyer, , 2023; Caragea et al., , 2023; Imaizumi and Fukumizu, , 2019, 2022). Here, in the conventional supervised learning setting, a dataset is commonly defined as pairs of independently and identically distributed (i.i.d.) random variables on a probability space $(\Omega,\Sigma,Q)$, which are drawn from the joint distribution of a $\mathcal{X}$-valued covariate $X$ and a label $Z$, namely
$(X_{1},Z_{1}),\cdots,(X_{n},Z_{n})\sim_{i.i.d.}Q\circ(X,Z)^{-1}.$
Kim et al., (2021) investigate deep ReLU networks in terms of the convergence rates in classification problems. The convergence rate in a binary classification problem, which has a decision boundary represented by some functions in a ball of Hölder space, is studied in (Kim et al., , 2021; Meyer, , 2023). Caragea et al., (2023) study binary classification in a setting where the decision boundary is defined with Barron functions introduced by Barron, (1993). Imaizumi and Fukumizu, (2019, 2022) show the convergence rates of deep neural networks in nonparametric regression, where they consider to estimate regression functions characterized by smooth boundaries defined with Hölder continuous functions.
However, in practice it is often costly to collect labels $Z_{1},\cdots,Z_{n}$ corresponding to the given covariates $X_{1},\cdots,X_{n}$. In the present work, we investigate the topic in another setting where some statistical relationship between covariates $X$ and $X^{\prime}$ is instead available. More precisely, we consider a dataset
$(X_{1},X_{1}^{\prime},Y_{1}),\cdots,(X_{n},X_{n}^{\prime},Y_{n})\sim_{i.i.d.}P,$ (1)
which consists of pairs of $\mathcal{X}$-valued random variables $X_{i},X_{i}^{\prime}:\Omega\to\mathcal{X}$ and a binary variable $Y_{i}:\Omega\to\mathcal{Y}=\{1,-1\}$, where $\mathcal{Y}$ is equipped with the counting measure, and $P$ is a Borel probability measure in $\mathcal{X}^{2}\times\mathcal{Y}$. This problem setting is extensively studied in the context of ranking problems (see, e.g., (Robbiano, , 2013)), similarity learning (see, e.g., (Chen et al., , 2009; Jin et al., , 2009; Cao et al., , 2016; Bao et al., 2022b, )), and self-supervised learning (see, e.g., (Arora et al., , 2019; Tosh et al., 2021a, ; Tosh et al., 2021b, ; Chen et al., , 2020; Tsai et al., , 2020; Chuang et al., , 2022; Zhai et al., , 2023; Balestriero et al., , 2023)), to improve the efficiency of learning procedures. Indeed, $Y$ always takes either $Y=1$ (i.e., $X$ and $X^{\prime}$ are similar) or $Y=-1$ (i.e., $X$ and $X^{\prime}$ are dissimilar).
Besides the aspect on data utilization, estimation of smooth boundaries based on pairwise relationship is also reasonable from some statistical learning viewpoints, provided that some additional assumption on subsets is introduced in the data space. For instance, the standard goal of binary classification is to obtain a hypersurface that classifies the observed covariates consistently in terms of the similarity determined by the Bayes classifier (see, e.g., (Hastie et al., , 2009)). Similarly in the case where one can observe paired covariates $(X,X^{\prime})$ and some binary variable $Y$ indicating the similarity, an analogous intuition may be still valid if one can divide the domain $\mathcal{X}$ into some subsets accordingly to the pairwise similarity. Note that in the present work, the notion is defined with the conditional probability of $Y=1$ given $(X,X^{\prime})$, similarly to the conventional binary classification (see Definition 1).
The above intuition is partially shown to be true by Bao et al., 2022b , where they theoretically showed the equivalence between conventional binary classification and a pairwise counterpart that uses a dataset of the form (1). However, conventional binary classification can deal with only one boundary. A natural question is about the statistical learnability of multiple smooth boundaries using paired covariates.
In the field of self-supervised learning, it is shown by HaoChen et al., (2021) that the learnability in a multiclass classification problem is guaranteed when both paired covariates and supervised data are used to define estimators. Nevertheless, to the best of our knowledge, the learnability of smooth boundaries is not proven when only data obtained in a pairwise binary classification setting is used. In the context of nonparametric statistics, it is shown by Kim et al., (2021) and Imaizumi and Fukumizu, (2019, 2022) that statistical learning of multiple boundaries using deep neural networks is possible if the observed data is a pair of single covariate and response variable. However, as we will see in Section 3.3, the arguments developed in (Kim et al., , 2021; Imaizumi and Fukumizu, , 2019, 2022) cannot apply to our problem setting because of a gap between the single-covariate setting and paired-covariate setting.
Therefore, in the present work we explore the problem under a general setting where nonparametric estimation of multiple smooth boundaries is considered. Specifically, in this work we investigate the following question:
Is it possible to estimate multiple smooth boundaries by using some learning algorithm that requires only samples generated in a pairwise binary classification setting (1)?
1.1 Approach
To investigate the problem, we need to ask what kind of pairwise relation is reasonable to be used in boundary estimation. In essence, our approach builds on the pairwise relation introduced by Tsai et al., (2020). More precisely, Tsai et al., (2020) introduce the density function of a probability measure $P$ in $\mathcal{X}^{2}\times\mathcal{Y}$ satisfying
$\begin{cases}\;\;p(x,x^{\prime},1)\;\;=p_{Y}(1)q(x,x^{\prime}),\\ \;p(x,x^{\prime},-1)=p_{Y}(-1)p_{X}(x)p_{X^{\prime}}(x^{\prime}),\end{cases}$ (2)
where $q(x,x^{\prime})$ denotes a probability density function on $\mathcal{X}^{2}$, $p_{X}(x)$ and $p_{X^{\prime}}(x^{\prime})$ are the marginal distributions of $q(x,x^{\prime})$, and $p_{Y}(y)$ is a probability function on $\mathcal{Y}$. This relation is useful in the following two points.
First, the pairwise relation introduced in (Tsai et al., , 2020) uses the statistical independence, which is simple and flexible. Also, the statistical independence can be checked by applying statistical independence tests (see, e.g., (Albert et al., , 2022)), which is reasonable from practical aspects. Second, it is possible to use algorithms that can handle large-scaled problems through the power of contrastive learning (see, e.g., (Gutmann and Hyvärinen, , 2010; van den Oord et al., , 2018; Arora et al., , 2019; Hénaff et al., , 2020; He et al., , 2020; Chen et al., , 2020; HaoChen et al., , 2021; Dwibedi et al., , 2021)).
Note that Tsai et al., (2020) use the pairwise relation (2) to study mutual information estimation and some application to contrastive learning, though nonparametric estimation of smooth boundaries using (2) is not investigated. Therefore, in the present work we will start by checking the validity of the problem setting (see Section 2.4).
As in (Kim et al., , 2021; Meyer, , 2023; Imaizumi and Fukumizu, , 2019, 2022), we formalize the problem using the indicator functions $\mathds{1}_{\mathcal{K}_{1}},\cdots,\mathds{1}_{\mathcal{K}_{d_{1}}}$. Following Imaizumi and Fukumizu, (2019, 2022), we analyze the $L^{2}$-risk between the indicator function and an estimator. The formal statement of the problem we address can be found in Section 2.3.
1.2 Main Contributions
The main contributions can be summarized in the following points.
Methodology:
In Section 3, we construct a methodology to learn smooth boundaries from the pairwise dataset (1), in the spirit of similarity learning (Jin et al., , 2009; Chen et al., , 2009; Cao et al., , 2016; Bao et al., 2022b, ) and contrastive learning (Arora et al., , 2019; Awasthi et al., , 2022; Wang and Isola, , 2020; Chen et al., 2021a, ). The learning algorithm produces an Empirical Risk Minimizer (ERM) in a localized class of functions (Algorithm 1). We give the theoretical justifications of the methodology in detail (Proposition 3, Proposition 6, and Proposition 8).
Results:
We show that the ERM $\widehat{g}_{n}^{\textup{LERM}}$ generated by Algorithm 1 is a consistent estimator of the indicator functions, where the $L^{2}$-risk is upper bounded by $n^{-\frac{\alpha}{\alpha+d_{0}-1}}\log^{4}{n}$ up to some constant independent of $n$ (Theorem 12). Here, $\alpha$ represents the smoothness of Hölder continuous functions characterizing the boundaries of subsets, following (Imaizumi and Fukumizu, , 2022). We also show an extension of the main theorem for the global ERM (Corollary 19), where the results indicate the statistical learnability of smooth boundaries using the global ERM under some additional conditions. Note that the convergence rate $n^{-\frac{\alpha}{\alpha+d_{0}-1}}$ is known as the minimax optimal rate for nonparametric boundary estimation in the literature (Mammen and Tsybakov, , 1999; Tsybakov, , 2004; Meyer, , 2023; Imaizumi and Fukumizu, , 2019, 2022). As far as we know, however, this is the first time that the convergence rate $n^{-\frac{\alpha}{\alpha+d_{0}-1}}$ up to a logarithmic factor is shown in a pairwise binary classification setting (1).
Technical contributions:
The proof of Theorem 12 is based on an estimation bound presented in Theorem 17. We find that in our estimation problem we need to conduct some delicate analyses using a localization property of function classes. This finding shares some similarity with the theory of empirical risk minimization in conventional supervised learning (Mendelson, , 2015, 2017), though our finding has a different mechanism due to the pairwise binary classification setting.
Application:
As a by-product of the main theorem on the $L^{2}$-risk, we obtain a theoretical result for multiclass nonparametric classification problems (Theorem 24). The result is stated in terms of the excess misclassification risk and is consistent with the related work by Kim et al., (2021).
1.3 Organization of the Paper
The rest of this paper consists of the following sections. In Section 2, we give the formal problem setting. In Section 3, we define an empirical risk minimizer, where the estimator is obtained by a contrastive learning algorithm. In Section 4, we show the main results of this work. In Section 5 we present additional study of the estimation problem, focusing on application to multiclass classification. Finally, in Section 6 we discuss the future directions where the $L^{2}$-risk bound might be utilized. Most of the proofs are deferred to Appendix A.
2 Problem Setting
In Section 2.1 and Section 2.2, we introduce additional notations and several classes of Borel probability measures, respectively. In Section 2.3 we present the formal statement of the estimation problem. In Section 2.4 we show that the problem setting is well-conditioned for solving the estimation problem.
2.1 Notations
In addition to the symbols defined in the introduction, we need some more notations. For some basic mathematical notions, we refer the readers to (Steinwart and Christmann, , 2008).
For a topological space $\mathcal{A}$, let $\mathcal{B}(\mathcal{A})$ denote the Borel $\sigma$-algebra. Given a finite non-negative measure $\nu$ on measurable space $(\mathcal{A},\mathcal{B}(\mathcal{A}))$, we define the semi-norms
$\|g\|_{L^{s}(\mathcal{A},\nu)}$ $=\left(\int_{\mathcal{A}}|g(x)|^{s}\nu(dx)\right)^{1/s},$
$\|g\|_{L^{\infty}(\mathcal{A},\nu)}$ $=\inf\{t\geq 0\;|\;\nu(\{x\in\mathcal{A}\;|\;|g(x)|>t\})=0\}.$
For $s\in[1,\infty]$, let $\mathcal{L}^{s}(\mathcal{A},\nu)$ be the set of all $s$-integrable real-valued functions on $(\mathcal{A},\mathcal{B}(\mathcal{A}),\nu)$. Given a measurable, vector-valued function $f$ on $(\mathcal{A},\mathcal{B}(\mathcal{A}),\nu)$, let $\|f\|_{\mathcal{A},\nu,s}:=\|\|f\|_{s}\|_{L^{s}(\mathcal{A},\nu)}$, where $\|\cdot\|_{s}$ denotes the $s$-norm in the Euclidean space. Note that when the Lebesgue measure $\mu$ on $\mathcal{X}$ is used, we often abbreviate as $\mathcal{L}^{s}(\mathcal{X}):=\mathcal{L}^{s}(\mathcal{X},\mu)$ and $\|f\|_{\mathcal{X},s}:=\|f\|_{\mathcal{X},\mu,s}$ for given $s\in[1,\infty]$ and vector-valued function $f$ on $(\mathcal{X},\mathcal{B}(\mathcal{X}),\mu)$.
Throughout the paper, we endow any finite set with the discrete topology and consider the measurable space equipped with the Borel $\sigma$-algebra. Also, the cardinality of a finite set is denoted by ||$\cdot$|| ⋅ |.
Let $p(x,x^{\prime},y)$ be a probability density function on $\mathcal{X}^{2}\times\mathcal{Y}$. Define the function $p_{X,X^{\prime}}(x,x^{\prime})=p(x,x^{\prime},1)+p(x,x^{\prime},-1)$. Note that
$p_{X}(x)=\int_{\mathcal{X}}p_{X,X^{\prime}}(x,x^{\prime})\mu(dx^{\prime}),\;\;p_{X^{\prime}}(x^{\prime})=\int_{\mathcal{X}}p_{X,X^{\prime}}(x,x^{\prime})\mu(dx).$
Denote by $P_{X}$ and $P_{X^{\prime}}$, the probability measures whose Lebesgue densities are $p_{X}$ and $p_{X^{\prime}}$, respectively. Define
$\eta(x,x^{\prime}):=p(y=1|x,x^{\prime})=\frac{p(x,x^{\prime},1)}{p_{X,X^{\prime}}(x,x^{\prime})}.$
Denote by $P_{X,X^{\prime}}$, the probability measure whose Lebesgue density is $p_{X,X^{\prime}}$. We also define the sign function as $\textup{sign}(s)=1$ if $s\geq 0$ and $\textup{sign}(s)=-1$ if $s<0$.
The domain of any random variable defined in this work is the probability space $(\Omega,\Sigma,Q)$.
For any $s_{1},s_{2}\in\mathbb{R}$, the notation $s_{1}\lesssim s_{2}$ means that there is a constant $C>0$ independent of the sample size $n$ such that $s_{1}\leq Cs_{2}$, unless otherwise specified.
Several lists summarizing the main notations used in the paper can be found in Appendix B.
2.2 Assumptions
Noise condition (NC).
In this paragraph, let $\mathcal{X}_{0}$ be a measurable space, and $\nu$ be a non-negative measure in $\mathcal{X}_{0}$. Now, consider a probability measure $P_{0}$ in $\mathcal{X}_{0}\times\mathcal{Y}$ such that it has a probability density function on $\mathcal{X}_{0}\times\mathcal{Y}$. Note that the marginal distribution of $P_{0}$ in $\mathcal{X}_{0}$ is denoted by $P_{X_{0}}$.
The Tsybakov noise condition introduced by Tsybakov, (2004) is an assumption requiring that there exists some $s_{0}\in(0,1]$ such that for every $s\in(0,s_{0}]$ it holds that
$P_{X_{0}}(\{x_{0}\in\mathcal{X}_{0}\;|\;|2p(y=1|x_{0})-1|\leq s\})\leq c\cdot s^{\frac{1}{\kappa-1}},$ (3)
where $\kappa\geq 1$ and $c\geq 0$ are constants. See also (Mammen and Tsybakov, , 1999) for a more general condition.
The Tsybakov noise condition is commonly used to establish convergence rates of statistical learning algorithms (see, e.g., (Bartlett et al., , 2006; Audibert and Tsybakov, , 2007; Lecué, , 2007)). For convenience, we say that a probability measure in $\mathcal{X}_{0}\times\mathcal{Y}_{0}$ satisfies $\kappa$-(NC) if it satisfies the Tsybakov noise condition (3) with parameter $\kappa$.
In the case $\kappa=1$, the Tsybakov noise condition (3) becomes the noise condition of Massart and Nédélec, (2006)
$\exists s_{0}\in(0,1]\quad\textup{such that}\quad P_{X_{0}}(\{x_{0}\in\mathcal{X}_{0}\;|\;|2p(y_{0}=1|x_{0})-1|\leq s_{0}\})=0.$
It is also shown by Alquier et al., (2019) that several learning algorithms attain the best convergence rates when $\kappa=1$.
Smooth boundaries.
We next recall the classes of smooth boundaries introduced by Imaizumi and Fukumizu, (2022). For a given $\alpha>0$ and $d\in\mathbb{N}$, denote by $\mathcal{C}^{\alpha,d}$, the unit ball of the $\alpha$-Hölder space on $[0,1]^{d}$, namely
$\mathcal{C}^{\alpha,d}=\left\{g:[0,1]^{d}\to\mathbb{R}\;|\;g\textup{ is }\alpha_{*}\textup{-times differentiable, }\|g\|_{\mathcal{C}^{\alpha,d}}\leq 1\right\},$
where $\alpha_{*}$ is defined as
$\alpha_{*}=\begin{cases}\lfloor\alpha\rfloor&\quad\textup{if }\alpha\notin\mathbb{N},\\ \alpha-1&\quad\textup{if }\alpha\in\mathbb{N}.\end{cases}$
Here, $\lfloor\alpha\rfloor$ denotes the largest integer smaller than or equal to $\alpha$. In addition, let $\bm{s}\in(\mathbb{N}\cup\{0\})^{d}$ denote the multi-index, $\partial^{\bm{s}}$ be the differential operator, and $\|g\|_{\infty}$ and $\|x\|_{\infty}$ be respectively the uniform norms for real-valued functions and vectors. Then, the functional $\|\cdot\|_{\mathcal{C}^{\alpha,d}}$ is defined as
$\|g\|_{\mathcal{C}^{\alpha,d}}=\sum_{\bm{s}:\|\bm{s}\|_{1}\leq\alpha_{*}}\|\partial^{\bm{s}}g\|_{\infty}+\sum_{\bm{s}:\|\bm{s}\|_{1}=\alpha_{*}}\sup_{\begin{subarray}{c}x,x^{\prime}\in\mathcal{X}\\ x\neq x^{\prime}\end{subarray}}\frac{|\partial^{\bm{s}}g(x)-\partial^{\bm{s}}g(x^{\prime})|}{\|x-x^{\prime}\|_{\infty}^{\alpha-\alpha_{*}}}.$
Then, the following conditions are introduced by Imaizumi and Fukumizu, (2022):
  1. (P1)
    Given $\mathscr{S}=\{\mathcal{J}_{i}\}_{i=1}^{d_{1}}$, there are disjoint subsets $\mathcal{I}_{1},\cdots,\mathcal{I}_{d_{1}}\subset\{1,-1\}^{d_{*}}$, functions $h_{1},\cdots,h_{d_{*}}\in\mathcal{C}^{\alpha,d_{0}-1}$, and indices $j_{1},\cdots,j_{d_{*}}\in\{1,\cdots,d_{0}\}$, such that for every $i=1,\cdots,d_{1}$,
    $\mathcal{J}_{i}=\bigcup_{(s_{1},\cdots,s_{d_{*}})\in\mathcal{I}_{i}}\bigcap_{k=1}^{d_{*}}\{x\in\mathcal{X}\;|\;s_{k}x_{j_{k}}\geq s_{k}h_{k}(x_{\setminus j_{k}})\},$
    where $x_{\setminus j}=(x_{1},\cdots,x_{j-1},x_{j+1},\cdots,x_{d_{0}})$.
  2. (P2)
    Given a sequence $\mathscr{S}=\{\mathcal{A}_{i}\}_{i=1}^{d_{1}}$ in $\mathcal{X}$, for every $i,j\in\{1,\cdots,d_{1}\}$ such that $i\neq j$, it holds that $\mu(\overline{\mathcal{A}_{i}}\cap\overline{\mathcal{A}_{j}})=0$, where $\overline{\mathcal{A}}_{i}$ denotes the closure of $\mathcal{A}_{i}$ in $\mathcal{X}$.
The following class $\widetilde{\mathscr{P}}_{\alpha}$ with fixed numbers $d_{1},d_{*}\in\mathbb{N}$ is introduced by Imaizumi and Fukumizu, (2022),
[…]
Here, the following variant of the condition (P1) for disjoint partitions is also covered by the theory of Imaizumi and Fukumizu, (2022):
  1. (P1’)
    Given $\mathscr{S}^{\prime}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$, there are disjoint subsets $\mathcal{I}_{1},\cdots,\mathcal{I}_{d_{1}}\subset\{1,-1\}^{d_{*}}$, functions $h_{1},\cdots,h_{d_{*}}\in\mathcal{C}^{\alpha,d_{0}-1}$, and indices $j_{1},\cdots,j_{d_{*}}\in\{1,\cdots,d_{0}\}$ such that for every $i=1,\cdots,d_{1}$,
    $\mathcal{K}_{i}$ $=\bigcup_{(s_{1},\cdots,s_{d_{*}})\in\mathcal{I}_{i}}\bigcap_{k=1}^{d_{*}}\widetilde{\mathcal{K}}_{s_{k},h_{k},j_{k}},$
    $\widetilde{\mathcal{K}}_{s_{k},h_{k},j_{k}}$ =$=\begin{cases}\{x\in\mathcal{X}\;|\;x_{j_{k}}\geq h_{k}(x_{\setminus j_{k}})\}&\quad\textup{ if }s_{k}=1,\\ \{x\in\mathcal{X}\;|\;x_{j_{k}}<h_{k}(x_{\setminus j_{k}})\}&\quad\textup{ if }s_{k}=-1.\end{cases}$
Note that a similar condition on multiple disjoint subsets is also used in (Kim et al., , 2021). The following class $\mathscr{P}_{\alpha}$ is a simplified class studied in (Imaizumi and Fukumizu, , 2022):
[…]
We use the class $\mathscr{P}_{\alpha}$ to simplify the proofs of the main theorem. Nevertheless, as in (Imaizumi and Fukumizu, , 2022), by the condition (P2), the approaches used in the proof of the main theorem are applicable when the risk function is defined with the class $\widetilde{\mathscr{P}}_{\alpha}$ instead of $\mathscr{P}_{\alpha}$.
Classically, classes of smooth boundaries are studied in (Dudley, , 1974; Mammen and Tsybakov, , 1995, 1999). Also, Hölder continuous boundaries are studied in the literature (Petersen and Voigtlaender, , 2018; Imaizumi and Fukumizu, , 2019; Kim et al., , 2021; Meyer, , 2023). Furthermore, the arguments developed in this work can be modified so that other smooth functions are employed. For instance, the approximation theorems for boundaries defined with Barron functions (Barron, , 1993) are studied in (Caragea et al., , 2023). In addition, the Besov space can be also employed if we instead apply the analyses of Suzuki, (2019), following the proofs of approximation theorems for indicator functions developed in (Petersen and Voigtlaender, , 2018; Imaizumi and Fukumizu, , 2019, 2022).
Main assumptions.
We introduce the following class of probability distributions.
Definition 1 (Class $\mathcal{P}_{\alpha}$).
Let $\alpha>0$. We define the class
[…]
where the conditions (A1) – (A4) are defined as follows:
  1. (A1)
    $P$ is absolutely continuous for the product measure $\mu\otimes\mu\otimes(\chi_{1}+\chi_{-1})$, where $\chi_{s}$ denotes the Dirac measure centered at $s\in\mathbb{R}$. The density $p(x,x^{\prime},y)$ of $P$ on $\mathcal{X}^{2}\times\mathcal{Y}$ satisfies the property (2), which is introduced by Tsai et al., (2020).
  2. (A2)
    $P$ satisfies $1$-(NC), i.e., the noise condition of Massart and Nédélec, (2006).
  3. (A3)
    In the condition (A1), $q(x,x^{\prime})$ is a symmetric function. Also, $p_{X}(x),p_{X^{\prime}}(x^{\prime})$ are positive and continuous at every $x,x^{\prime}\in\mathcal{X}$. In addition, $p_{Y}(1)\in(0,1)$ holds.
  4. (A4)
    There is a sequence $\mathscr{S}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}\in\mathscr{P}_{\alpha}$ such that $P_{X}(\mathcal{K}_{i})<1$ for every $i\in\{1,\cdots,d_{1}\}$, and we have
    $\left\{(x,x^{\prime})\in\mathcal{X}^{2}\;\Big{|}\;\eta(x,x^{\prime})\geq\frac{1}{2}\right\}=\bigcup_{i=1}^{d_{1}}\mathcal{K}_{i}\times\mathcal{K}_{i}.$ (4)
We also define $\mathcal{P}=\bigcup_{\alpha>0}\mathcal{P}_{\alpha}$.
Note that the condition (A3) summarizes several technical conditions used in the proofs. In the condition (4), the boundaries are consistently defined with the sign of Bayes classifier, which is analogous to the conventional binary classification.
In the context of contrastive learning theory, the condition (4) is weaker than the formulation of Awasthi et al., (2022). Indeed, Awasthi et al., (2022) use a joint distribution introduced by Arora et al., (2019) and assume that the joint probability at any two data points belonging to the different disjoint subsets is zero. On the other hand, the condition is stronger than the settings studied by Waida et al., (2023) and Parulekar et al., (2023). The main difference to the formulation of Waida et al., (2023) is that we consider a setting of binary classification rather than contrastive learning, which requires to use disjoint partitions of the space $\mathcal{X}$. Also, (4) is understood as an example of the formulation introduced by Parulekar et al., (2023) since it defines an equivalence relation in $\mathcal{X}$. Meanwhile Parulekar et al., (2023) use a general equivalence relation using latent variables introduced by von Kügelgen et al., (2021).
Since any partition $\mathscr{S}$ in $\mathscr{P}_{\alpha}$ is expressed as a sequence, we need to introduce an equivalence relation that identifies the sequences containing the same subsets: namely, given two sequences $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ and $\{\mathcal{K}_{i}^{\prime}\}_{i=1}^{d_{1}}$ of subsets in $\mathcal{X}$, we define the equivalence relation $\sim$ as $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}\sim\{\mathcal{K}_{i}^{\prime}\}_{i=1}^{d_{1}}$ if and only if there is a permutation $\pi$ of the set $\{1,\cdots,d_{1}\}$ such that $\mathcal{K}_{\pi(i)}=\mathcal{K}_{i}^{\prime}$ for any $i\in\{1,\cdots,d_{1}\}$. Denote the equivalence class where the sequence $\mathscr{S}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ belongs to, by $[\mathscr{S}]$.
Definition 2.
Given $\alpha>0$ and $P\in\mathcal{P}_{\alpha}$, let $\mathscr{S}\in\mathscr{P}_{\alpha}$ be a partition satisfying the condition (A4) in Definition 1. Then, we define as $\mathscr{S}_{P}=[\mathscr{S}]$.
Note that by the condition (4), the map $P\mapsto\mathscr{S}_{P}$ on $\mathcal{P}_{\alpha}$ is well-defined. Also, given an equivalence class $[\mathscr{S}]\in\mathscr{P}_{\alpha}/\sim$, without loss of generality we can fix an arbitrary partition $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}\in[\mathscr{S}]$ in the proof of the main results of this work, by the definition of the equivalence relation.
2.3 Formal Statement of the Problem
Let $\widehat{g}_{n}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{G}_{0}$ be an estimator, where
$\mathcal{G}_{0}=\{g:\mathcal{X}\to[0,1]^{d_{1}}\;|\;g_{1},\cdots,g_{d_{1}}\textup{ are measurable},\;\|g(\cdot)\|_{1}=1\;\textup{on }\mathcal{X}\}.$
Here, $\|\cdot\|_{1}$ denotes the 1-norm in the Euclidean space. Given $\alpha>0$, let $P\in\mathcal{P}_{\alpha}$. Then, we aim at estimating the sets in the sequence $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}\in\mathscr{S}_{P}$. Similarly to Imaizumi and Fukumizu, (2019, 2022), we analyze the convergence rates of estimators in terms of the $L^{2}$-risk. Formally, the estimation error of an estimator $\widehat{g}_{n}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{G}_{0}$ is measured by the sum of $L^{2}$-risks
$\mathcal{R}(\widehat{g}_{n};P,U_{1}^{n})=\mathbb{E}\left[\sum_{i=1}^{d_{1}}\|\widehat{g}_{n,i}(U_{1},\cdots,U_{n})-\mathds{1}_{\mathcal{K}_{i}}\|_{L^{2}(\mathcal{X},P_{X})}^{2}\right],$ (5)
where $U_{1}^{n}=(U_{1},\cdots,U_{n})$ is any sequence of i.i.d. $\mathcal{X}^{2}\times\mathcal{Y}$-valued random variables $U_{1}=(X_{1},X_{1}^{\prime},Y_{1}),\cdots,U_{n}=(X_{n},X_{n}^{\prime},Y_{n})$ drawn from the distribution $P$, the expectation in the right-hand side of (5) is taken in terms of the distribution of $U_{1}^{n}$, and $P_{X}$ is the marginal distribution of $P$ with respect to the variable $X$.
We now introduce the estimation problem we shall investigate in this work.
Problem.
Given $P\in\mathcal{P}_{\alpha}$, is there an estimator $\widehat{g}_{n}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{G}_{0}$ such that the $L^{2}$-risk (5) defined with any sequence $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ in $\mathscr{S}_{P}$ tends to zero as $n\to\infty$? In particular:
  1. Q1.
    Is it reasonable to use the pairwise dependency to estimate smooth boundaries?
  2. Q2.
    How can one construct an estimator so that its statistical performance is guaranteed?
  3. Q3.
    How fast does the estimator converge to the true function?
Roughly speaking, we address these problems as follows: The problem (Q1) is resolved in the subsequent Section 2.4. Regarding the problem (Q2), we construct an algorithm based on contrastive learning and define an empirical risk minimizer in Section 3. The problem (Q3) is investigated in Section 4.
2.4 Validity of the Estimation Problem
It is natural to ask whether (A1) – (A4) are well-conditioned for estimating smooth partitions introduced by Imaizumi and Fukumizu, (2022). In what follows, we show that the range of the map $P\mapsto\mathscr{S}_{P}$ on $\mathcal{P}_{\alpha}$ is expressed in terms of the class
$\mathscr{P}_{\alpha,+}=\{\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}\in\mathscr{P}_{\alpha}\;|\;\mu(\mathcal{K}_{i})>0\textup{ for any }i=1,\cdots,d_{1}\}.$
Proposition 3.
For every $\mathscr{S}\in\mathscr{P}_{\alpha,+}$, there is a Borel probability measure $P\in\mathcal{P}_{\alpha}$ such that we have $[\mathscr{S}]=\mathscr{S}_{P}$.
The proof is deferred to Appendix A.1. To show the existence, we follow the approach developed in the literature (Arora et al., , 2019; Awasthi et al., , 2022). However, the above property has not been studied.
In other words, this proposition implies that statistical learning with samples drawn from any distribution in $\mathcal{P}_{\alpha}$ covers every smooth partition belonging to the class $\mathscr{P}_{\alpha,+}$. Thus, it is reasonable to focus on the class $\mathcal{P}_{\alpha}$.
3 Learning Procedure
In the estimation procedure we execute an algorithm based on contrastive learning, which is known as an efficient, tractable methodology to learn pairwise relation (see, e.g., (Gutmann and Hyvärinen, , 2010; van den Oord et al., , 2018; Arora et al., , 2019; Chen et al., , 2020)). In Section 3.1 and Section 3.2 we define a binary classification loss in the setting of contrastive learning. In Section 3.3 and Section 3.4 we introduce a theoretical estimator used in the main results. The notations of deep ReLU networks can be found in Section 3.5.
3.1 Preliminary
In order to estimate the indicator functions $\mathds{1}_{\mathcal{K}_{1}},\cdots,\mathds{1}_{\mathcal{K}_{d_{1}}}$, it is natural to use a loss function whose population risk minimizer can distinguish these subsets. To do so, we follow the theory developed by Awasthi et al., (2022) to use vector-valued functions in the algorithm. Before expositing the details of contrastive learning in the next subsection, we shall explain why estimation of vector-valued functions make sense to study the main problem.
Similarly to (Awasthi et al., , 2022), we make use of regular simplices in the algorithm. Let $d\in\mathbb{N}$. Denote by $\mathcal{S}^{d-1}$, the unit hypersphere in $\mathbb{R}^{d}$. Let $v_{1},\cdots,v_{d+1}\in\mathcal{S}^{d-1}$ be vectors satisfying $\sum_{i=1}^{d+1}v_{i}=\bm{0}$ and the condition that $\|v_{i}-v_{j}\|_{2}=\|v_{i^{\prime}}-v_{j^{\prime}}\|_{2}$ for any $i,j,i^{\prime},j^{\prime}\in\{1,\cdots,d+1\}$ such that $i\neq j$ and $i^{\prime}\neq j^{\prime}$ (see e.g., (Conn et al., , 2009, Corollary 2.6)). Then, a regular simplex $\Delta^{d}$ is commonly defined as
$\Delta^{d}=\left\{c_{1}v_{1}+\cdots+c_{d+1}v_{d+1}\;|\;\bm{c}=(c_{1},\cdots,c_{d+1})\in[0,1]^{d+1},\|\bm{c}\|_{1}=1\right\}.$
We endow $\Delta^{d}$ with the subspace topology from $\mathbb{R}^{d}$. We also define
$D_{\Delta^{d}}=\max_{z,z^{\prime}\in\Delta^{d}}\|z-z^{\prime}\|_{2}.$
It is shown by Alexander, (1977, Theorem 3) that $D_{\Delta^{d}}=\|v_{i}-v_{j}\|_{2}$ holds for any distinct vertices $v_{i}$ and $v_{j}$.
Remark 4.
Hereafter, we only consider the case $d=d_{1}-1$. Namely, in this section we develop an algorithm that requires the value of $d_{1}$ to be known in advance. This condition is mild both in the contexts of nonparametric boundary estimation and machine learning: Firstly, the definition of partitions introduced by Imaizumi and Fukumizu, (2022), which is also used in this work, implicitly assumes that the number of subsets should be less than or equal to $2^{d_{*}}$. Besides, in the context of contrastive learning some similar assumptions on the number of subsets are considered in (HaoChen et al., , 2021; Awasthi et al., , 2022; Parulekar et al., , 2023).
Define a function $f:\mathcal{X}\to\Delta^{d}$. By the definition of $\Delta^{d}$, there are some functions $g_{1},\cdots,g_{d_{1}}:\mathcal{X}\to\mathbb{R}$ such that $f=\sum_{i=1}^{d_{1}}g_{i}v_{i}$. Note that such functions $g_{1},\cdots,g_{d_{1}}$ are uniquely determined for each $f$, since $v_{1},\cdots,v_{d_{1}}$ are affinely independent (see e.g., (Hatcher, , 2002, pp.102–104)). Define
$\mathcal{F}_{0}=\left\{f:\mathcal{X}\to\Delta^{d}\;\Big{|}\;f=\sum_{i=1}^{d_{1}}g_{i}v_{i},\;g_{1},\cdots,g_{d_{1}}\textup{ are measurable}\right\}.$
We define a contrastive function, which is slightly generalized from Definition 3.7 in (Awasthi et al., , 2022).
Definition 5 (Contrastive function).
Let $P\in\mathcal{P}$. The contrastive function $f^{*}:\mathcal{X}\to\Delta^{d}$ of $P$ is defined as
$f^{*}(x)=\sum_{i=1}^{d_{1}}\mathds{1}_{\mathcal{K}_{i}}(x)v_{i}.$
We give several remarks on Definition 5.
  • Awasthi et al., (2022, Definition 3.7) also introduce a simplex-valued function defined with indicator functions. The difference is that the definition of subsets $\mathcal{K}_{1},\cdots,\mathcal{K}_{d_{1}}$ requires a weaker condition than the setting considered in (Awasthi et al., , 2022).
  • Also, some other vector-valued functions related to simplices are studied in (HaoChen et al., , 2021). However, HaoChen et al., (2021) mainly consider any function that is uniformly bounded with respect to the uniform norm. Meanwhile, in our analysis the range of any $f$ is restricted to $\Delta^{d}$, which is due to several useful properties (see Proposition 6 and Proposition 8).
  • Besides, in a supervised metric learning scenario (Khosla et al., , 2020), some similar vector-valued functions are also studied in (Graf et al., , 2021; Zhu et al., , 2022; Chen et al., , 2022). In our setting, however, no label showing the category of any covariate is used.
  • In Definition 5 the contrastive function $f^{*}$ is defined as the sum of vector-valued functions $\mathds{1}_{\mathcal{K}_{i}}v_{i}$, $i=1,\cdots,d_{1}$, for simplicity, although it is straightforward to extend the subsequent analyses to the case that $f^{*}$ is defined as a function that equals the sum for $P_{X}$-almost every points in $\mathcal{X}$.
The utility of the contrastive function is justified by the following simple fact:
Proposition 6.
Let $f,f^{\prime}\in\mathcal{F}_{0}$, and denote by $f=\sum_{i=1}^{d_{1}}g_{i}v_{i}$ and $f^{\prime}=\sum_{i=1}^{d_{1}}g_{i}^{\prime}v_{i}$. Then, we have
$\|f-f^{\prime}\|_{\mathcal{X},P_{X},2}^{2}=\frac{d_{1}}{d}\sum_{i=1}^{d_{1}}\|g_{i}-g_{i}^{\prime}\|_{L^{2}(\mathcal{X},P_{X})}^{2}.$
The proof can be found in Appendix A.2. By Proposition 6, for every $P\in\mathcal{P}$ we have
$\mathcal{R}(\widehat{g}_{n};P,U_{1}^{n})=d_{1}^{-1}d\cdot\mathbb{E}[\|\widehat{f}_{n}(U_{1},\cdots,U_{n})-f^{*}\|_{\mathcal{X},P_{X},2}^{2}],$
where $\widehat{f}_{n}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{F}_{0}$ is an estimator such that $\widehat{f}_{n}=\sum_{i=1}^{d_{1}}\widehat{g}_{n,i}v_{i}$, and $f^{*}$ is the contrastive function of $P$. Thus, it suffices to show an upper bound of the $L^{2}$-risk $\mathbb{E}[\|\widehat{f}_{n}(U_{1},\cdots,U_{n})-f^{*}\|_{\mathcal{X},P_{X},2}^{2}]$.
3.2 Loss Function
We next define the loss function used in the learning algorithm. As presented in Section 1.1, we use contrastive learning. Note that contrastive learning is originally developed by Gutmann and Hyvärinen, (2010) as a parametric estimation method, where they consider a binary classification problem for distinguishing the true distribution from a noise distribution. This approach has recently been investigated as a methodology to learn the statistical relationship between covariates in pairwise binary classification settings (Tsai et al., , 2020; Tosh et al., 2021a, ; Tosh et al., 2021b, ; Bao et al., 2022b, ; Chuang et al., , 2022; Zhai et al., , 2023). Hence, we shall define an algorithm using some binary classification loss.
Following (Chen et al., 2021a, ; Wang and Isola, , 2020), we consider a loss function defined with the squared Euclidean distance, namely, $\rho_{f}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ defined as
$\rho_{f}(x,x^{\prime})=\|f(x)-f(x^{\prime})\|_{2}^{2},$
for each $f\in\mathcal{F}_{0}$. Note that the Euclidean distance is used for simplicity. It might be possible to use other distance functions (see, e.g., (Dovgoshey and Petrov, , 2013)); we leave further investigation of this point as a future work.
Following the standard approach in similarity learning and metric learning (Jin et al., , 2009; Chen et al., , 2009; Cao et al., , 2016; Bao et al., 2022b, ), we define a hinge loss as follows:
Definition 7 (Loss function).
Let $\psi:\mathbb{R}\to\mathbb{R}$ be the function defined as
$\psi(s)=1-2D_{\Delta^{d}}^{-2}s.$
Then, for every $f\in\mathcal{F}_{0}$, we define the hinge loss $\ell_{f}:\mathcal{X}^{2}\times\mathcal{Y}\to\mathbb{R}$ as
$\ell_{f}(x,x^{\prime},y)=\max\{0,1-y\psi\circ\rho_{f}(x,x^{\prime})\}.$
The function $\psi\circ\rho_{f}$ is a binary classifier, since $\psi$ controls the range of $\rho_{f}$. In fact, the classifier returns the prediction equal to $+1$ if the covariates $x$ and $x^{\prime}$ coincide. In addition, by Theorem 3 in (Alexander, , 1977), the prediction equals $-1$ if and only if these points are embedded in different vertices on $\Delta^{d}$.
Note that a similar classifier defined with the Mahalanobis distance is proposed in (Jin et al., , 2009). On the other hand, in Definition 7, we employ the Euclidean distance and additionally embed the variables $x$ and $x^{\prime}$ in the regular simplex. Furthermore, Jin et al., (2009) define the binary variable by the information showing whether the supervised labels of two covariates coincide or not. Meanwhile, we use the statistical dependence to define the distribution of $(X,X^{\prime},Y)$, as in (Tsai et al., , 2020).
Hinge loss is often studied in the literature on statistical learning (e.g., (Boser et al., , 1992; Lin, , 2002; Zhang, , 2004; Bartlett et al., , 2006; Lecué, , 2007; Kim et al., , 2021)). In particular, in a setting of conventional binary classification, Kim et al., (2021, Theorem 3.2) study nonparametric estimation of smooth boundaries using empirical risk minimization with hinge loss. Note that it is common in the literature (Kim et al., , 2021; Meyer, , 2023; Imaizumi and Fukumizu, , 2019, 2022) that nonparametric estimation of smooth boundaries is studied with some specific loss functions, such as 0-1 loss in (Meyer, , 2023) and squared loss in (Imaizumi and Fukumizu, , 2019, 2022). Furthermore, hinge loss and its variants are also studied in the context of self-supervised learning (Arora et al., , 2019; Li et al., , 2021; Shah et al., , 2022; Waida et al., , 2023; Ji et al., , 2023) and similarity learning (Jin et al., , 2009; Cao et al., , 2016; Zhou et al., , 2024). We use this loss function in a contrastive learning algorithm, as it belongs to a general class of loss functions of contrastive learning proposed in (Chen et al., 2021b, ).
The motivation of using the above loss function is due to the following property, which utilizes a fact shown by Lin, (2002).
Proposition 8.
Let $P\in\mathcal{P}$. If $f^{*}$ is the contrastive function of $P$, then we have
$\mathbb{E}_{P}[\ell_{f^{*}}]=\inf_{f\in\mathcal{F}_{0}}\mathbb{E}_{P}[\ell_{f}].$
The proof is deferred to Appendix A.3. The key point is that the range of vector-valued functions in $\mathcal{F}_{0}$ is restricted to $\Delta^{d}$. There are several remarks on Proposition 8:
  • This proposition slightly generalizes some part of Theorem 3.8 in (Awasthi et al., , 2022) for hinge loss since we investigate in a more general setting.
  • Note that the relationship between regular simplices and empirical risk minimizers in representation learning is shown in (Liu et al., , 2021; Lee et al., , 2024; Koromilas et al., , 2024). On the other hand, in the proposition above we are mainly interested in the population risk minimizers.
  • Zhou et al., (2024, Theorem 1) show that the minimizer of the population risk with respect to a hinge loss defined in a similarity learning setting embeds covariates in the probability simplex. Since Proposition 8 deals with a different problem setting, the assumptions used in the present work are distinct from the ones considered in (Zhou et al., , 2024).
3.3 Classes of Vector-Valued Functions
Next, we introduce function classes to define estimators. We shall recall that the goal of the estimation problem is to generate an estimator that is pointwisely close to the true function under the $L^{2}$-risk.
Since the dataset available in our setting provides only the information derived from the pairwise relation between $X$ and $X^{\prime}$, it is impossible to apply the approaches developed in the literature (Kim et al., , 2021; Imaizumi and Fukumizu, , 2019, 2022) directly to our setting. This restriction also yields a technical issue: specifically, it is not always true that for every $f:\mathcal{X}\to\Delta^{d}$, there is a constant $c=c(f)>0$ such that $\|f-f^{*}\|_{2}^{2}\leq c\cdot|\psi\circ\rho_{f}(\cdot,x^{\prime})-\psi\circ\rho_{f^{*}}(\cdot,x^{\prime})|$ for any $x^{\prime}\in\mathcal{X}$, because these quantities measure distinct types of relations between vectors (this argument is motivated by considering how to utilize a useful inequality shown in Proposition 1 of (Lecué, , 2007), in a pairwise binary classification setting; the details will be shown in Section 4.2.1). A counterexample to the inequality is given below.
Example 9.
Let $P\in\mathcal{P}$. Consider the function $\widetilde{f}^{*}=T\circ f^{*}$, where $f^{*}$ is the contrastive function of $P$, and $T:\mathbb{R}^{d}\to\{v_{1},\cdots,v_{d_{1}}\}\cup\{\bm{0}\}$ is defined as
$T(z)=\begin{cases}v_{j+1}&\textup{if }z=v_{j},\;j\in\{1,\cdots,d_{1}-1\},\\ v_{1}&\textup{if }z=v_{d_{1}},\\ \bm{0}&\textup{otherwise}.\end{cases}$
The map $T$ permutes the order of vertices. Then, we have $\|\widetilde{f}^{*}(x)-f^{*}(x)\|_{2}^{2}>0$ for every $x\in\mathcal{X}$. In contrast, by Theorem 3 in (Alexander, , 1977), for any $x^{\prime}\in\mathcal{X}$ we have $|\psi\circ\rho_{\widetilde{f}^{*}}(\cdot,x^{\prime})-\psi\circ\rho_{f^{*}}(\cdot,x^{\prime})|=0$ on $\mathcal{X}$.
We will discuss this point in detail in Section 4.3.1. Alternatively, we introduce the following notion:
Definition 10 ($(\beta,\beta_{0},P)$-localized subclass).
Let $\beta>0$, $\beta_{0}\geq 0$, and $P\in\mathcal{P}$. Then, the $(\beta,\beta_{0},P)$-localized subclass of a set $\mathcal{F}\subset\mathcal{F}_{0}$, which is denoted by $\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$, is defined as
$\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})=\left\{f\in\mathcal{F}\;\Big{|}\;P_{X}\left(\{x\in\mathcal{X}\;|\;\|f(x)-f^{*}(x)\|_{2}<\beta\}\right)\geq 1-\beta_{0}\right\}.$
Thanks to the structural condition on the model class, the exceptional case in Example 9 is avoidable. Here, let $D_{\textup{proj}}$ denote the distance between $v_{1}$ and the simplex that does not contain $v_{1}$, namely
$D_{\textup{proj}}=\inf\left\{\|z-v_{1}\|_{2}\;\Big{|}\;z=0\cdot v_{1}+\sum_{i=2}^{d_{1}}c_{i}v_{i},\;\|\bm{c}\|_{1}=1\right\}.$
In the subsequent analyses we are particularly interested in the parametrization $\mathscr{F}_{\beta,\beta^{-1}\varepsilon,P}$, where $\beta\in(0,D_{\textup{proj}})$, $\varepsilon>0$, and $P\in\mathcal{P}$. One can verify that for a fixed $\beta\in(0,D_{\textup{proj}})$, $\mathscr{F}_{\beta,\beta^{-1}\varepsilon,P}(\mathcal{F})\subseteq\mathscr{F}_{\beta,\beta^{-1}\varepsilon^{\prime},P}(\mathcal{F})$ for any $0\leq\varepsilon\leq\varepsilon^{\prime}$ and $\mathcal{F}\subset\mathcal{F}_{0}$. Moreover, for any sequence $\varepsilon_{n}>0$ such that $\varepsilon_{n}\to 0$ as $n\to\infty$, we have
$\liminf_{n\to\infty}\mathscr{F}_{\beta,\beta^{-1}\varepsilon_{n},P}(\mathcal{F})=\{f\in\mathcal{F}\;|\;\|f-f^{*}\|_{2}<\beta\textup{ almost surely}\}.$
Intuitively, we can use $\mathscr{F}_{\beta,\beta^{-1}\varepsilon,P}(\mathcal{F})$ to approximate the contrastive function $f^{*}$ if $\varepsilon$ is sufficiently small. The localized subclass also contains vector-valued functions that are pointwisely far apart from the true function $f^{*}$ since $\beta$ can take any value in $(0,D_{\textup{proj}})$. In other words, the localized subclass excludes functions that map $x\in\mathcal{K}_{i}$ to some wrong vertex $v_{j}$ with non-small probability.
Note that some related concepts are introduced in (Schiebinger et al., , 2015; Trillos et al., , 2021) in the context of clustering theory. Indeed, if $f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$ for some $\mathcal{F}\subseteq\mathcal{F}_{0}$, $\beta>0$, $\beta_{0}\geq 0$, and $P\in\mathcal{P}$, then it holds that
$P_{X}\circ f^{-1}\left(\bigcup_{i=1}^{d_{1}}\{z\in\Delta^{d}\;|\;\|z-v_{i}\|_{2}<\beta\}\right)\geq 1-\beta_{0}.$
Schiebinger et al., (2015, Definition 1) introduce an embedding condition based on finite samples. Trillos et al., (2021, Definition 8) consider a population setting and use the angles between a vector $z$ and orthonormal bases in the Euclidean space. Meanwhile, we use the distance from each vertex.
The notion of localized subclasses is also related to the condition referred to as “small-ball condition” by Mendelson, (2015, p.10). Specifically, Mendelson, (2015, Assumption 3.1) additionally normalizes the difference $g-g^{\prime}$ ($g,g^{\prime}\in L^{2}(\mathcal{A})$ for a measure space $\mathcal{A}$) and considers a condition on the tail probability of the normalized gap. Also, Mendelson, (2017, Definition 2.1) employs a sub-Gaussian condition. On the other hand, in our analysis it suffices to use an unnormalized gap as defined in Definition 10.
3.4 Estimators
Although the localized subclass has a useful property, it is not guaranteed in general that the $L^{2}$-risk $\mathbb{E}[\|f-f^{*}\|_{\mathcal{X},P_{X},2}^{2}]$ becomes sufficiently small for any $f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$ because $\beta$ is a constant. To select good functions, we consider the ERM estimator within a local function class.
Definition 11 ($(\beta,\varepsilon,P,n,\mathcal{F})$-local ERM).
Let $\beta\in(0,D_{\textup{proj}})$, $n\in\mathbb{N}\setminus\{1,2\}$, $\varepsilon>0$, $P\in\mathcal{P}$, and $\mathcal{F}\subset\mathcal{F}_{0}$. Consider the $(\beta,\beta^{-1}\varepsilon,P)$-localized subclass $\mathscr{F}_{\beta,\beta^{-1}\varepsilon,P}(\mathcal{F})$ of $\mathcal{F}$. In addition, let $\ell_{f}$ be the hinge loss in Definition 7. Here, define $\widehat{f}_{n}^{\textup{LERM}}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathscr{F}_{\beta,\beta^{-1}\varepsilon,P}(\mathcal{F})$ as
$\widehat{f}_{n}^{\textup{LERM}}(u_{1},\cdots,u_{n})=\operatorname*{arg\;min}_{f\in\mathscr{F}_{\beta,\beta^{-1}\varepsilon,P}(\mathcal{F})}\frac{1}{n}\sum_{i=1}^{n}\ell_{f}(u_{i}),$
where $u_{i}=(x_{i},x_{i}^{\prime},y_{i})\in\mathcal{X}^{2}\times\mathcal{Y}$, $i=1,\cdots,n$. Then, we define the $(\beta,\varepsilon,P,n,\mathcal{F})$-local ERM $\widehat{g}_{n}^{\textup{LERM}}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{G}_{0}$ as
$\widehat{g}_{n}^{\textup{LERM}}=(\widehat{g}_{n,1}^{\textup{LERM}},\cdots,\widehat{g}_{n,d_{1}}^{\textup{LERM}})\;\;\textup{such that}\;\;\widehat{f}_{n}^{\textup{LERM}}=\sum_{i=1}^{d_{1}}\widehat{g}_{n,i}^{\textup{LERM}}v_{i}.$
Note that if $\widehat{f}_{n}^{\textup{LERM}}$ is well-defined for every $(u_{1},\cdots,u_{n})\in(\mathcal{X}^{2}\times\mathcal{Y})^{n}$, then $\widehat{g}_{n}^{\textup{LERM}}$ is uniquely determined, since $v_{1},\cdots,v_{d_{1}}$ are affinely independent. For simplicity, we assume that the estimator $\widehat{f}_{n}^{\textup{LERM}}$ is always well-defined for the given subclass $\mathscr{F}_{\beta,\beta^{-1}\varepsilon,P}(\mathcal{F})$. Nevertheless, the generalization can be made straightforwardly by considering some other model belonging to a sufficiently small neighborhood of ERM.
The ERM depends on the structural condition defined in Definition 10. The condition is regarded as a regularization for the function class that enforces it to concentrate around the true function $f^{*}$. Notably, the relation between empirical risk minimization in a convex class of functions and its localization is studied by Mendelson, (2015, 2017). However, we need the localization due to some mechanism different from the one used in the theory of Mendelson, (2015, 2017). We discuss this point in detail in Section 4.4, where we also introduce the global ERM and some generalized analyses.
3.5 Deep ReLU Networks
We introduce notations of deep ReLU networks. Let $L\in\mathbb{N}$, $d_{\textup{NN},0},\cdots,d_{\textup{NN},L}\in\mathbb{N}$, and $\bm{d}=(d_{\textup{NN},0},\cdots,d_{\textup{NN},L})$. Denote the ReLU activation by $\sigma:\mathbb{R}\to\mathbb{R}$,
$\sigma(s)=s\vee 0=\max\{s,0\}.$
Then, we define ReLU networks on $\mathcal{X}$ as
$g_{\bm{W},\bm{b}}(x)=A_{L}\circ\bm{\sigma}_{L-1}\circ A_{L-1}\circ\cdots\circ\bm{\sigma}_{1}\circ A_{1}(x),$
where $\bm{W}\in\mathcal{W}_{L,\bm{d}}$ and $\bm{b}\in\mathcal{V}_{L,\bm{d}}$ are
$\mathcal{W}_{L,\bm{d}}$ =$=\{(W_{1},\cdots,W_{L})|W_{i}\in\mathbb{R}^{d_{\textup{NN},i}\times d_{\textup{NN},i-1}},i=1,\cdots,L\},$
$\mathcal{V}_{L,\bm{d}}$ =$=\{(b_{1},\cdots,b_{L})\;|\;b_{i}\in\mathbb{R}^{d_{\textup{NN},i}},i=1,\cdots,L\},$
and we define for each $i=1,\cdots,L$, that
$A_{i}(z)=W_{i}z+b_{i},\;\;\bm{\sigma}_{i}(z)=(\sigma(z_{1}),\cdots,\sigma(z_{d_{\textup{NN},i}}))^{\top}\textup{ for every }z\in\mathbb{R}^{d_{\textup{NN},i}}.$
In (Nakada and Imaizumi, , 2020; Imaizumi and Fukumizu, , 2019, 2022), for any $L_{0}\in\mathbb{N}$, $J,S,M\geq 0$, and $\bm{d}=(d_{\textup{NN},0},\cdots,d_{\textup{NN},L})$, the class $\bigcup_{L\leq L_{0}}\mathcal{F}_{L,J,S,M,\bm{d}}^{\textup{NN}}$ is studied, where the subclass $\mathcal{F}_{L,J,S,M,\bm{d}}^{\textup{NN}}$ is defined as
$\mathcal{F}_{L,J,S,M,\bm{d}}^{\textup{NN}}$
[…]
and for $W_{i}=(W_{i,j_{1},j_{2}})\in\mathbb{R}^{d_{\textup{NN},i}\times d_{\textup{NN},i-1}}$ and $b_{i}=(b_{i,j})\in\mathbb{R}^{d_{\textup{NN},i}}$,
$\|\bm{W}\|_{\max}=\max_{i=1,\cdots,L}\max_{j_{1},j_{2}}|W_{i,j_{1},j_{2}}|,\quad$ $\|\bm{b}\|_{\max}=\max_{i=1,\cdots,L}\max_{j}|b_{i,j}|,$
$\|\bm{W}\|_{0}=\sum_{i=1}^{L}|\{(j_{1},j_{2})\;|\;W_{i,j_{1},j_{2}}\neq 0\}|,\quad$ $\|\bm{b}\|_{0}=\sum_{i=1}^{L}|\{j\;|\;b_{i,j}\neq 0\}|.$
The class $\mathcal{F}_{L,J,S,M,\bm{d}}^{\textup{NN}}$ is also commonly-used in the literature (Schmidt-Hieber, , 2020; Petersen and Voigtlaender, , 2018; Bos and Schmidt-Hieber, , 2022). Note that the class of ReLU networks defined in the same way as $\mathcal{F}_{L,J,S,M,\bm{d}}^{\textup{NN}}$ except for the domain $\mathcal{A}$ of networks is denoted by $\mathcal{F}_{L,J,S,M,\bm{d}}^{\textup{NN}}(\mathcal{A})$.
We need to select the subclass to evaluate a covering number, as in the literature on nonparametric regression (Schmidt-Hieber, , 2020; Nakada and Imaizumi, , 2020; Imaizumi and Fukumizu, , 2019, 2022). In the main result, we apply an estimate shown by Nakada and Imaizumi, (2020).
To derive upper bounds of the approximation errors, we need to approximate the indicator functions $\mathds{1}_{\mathcal{K}_{1}},\cdots,\mathds{1}_{\mathcal{K}_{d_{1}}}$. To do so, we apply the useful facts shown in (Petersen and Voigtlaender, , 2018; Imaizumi and Fukumizu, , 2019, 2022). However, direct application is impossible since the range of vector-valued functions in $\mathcal{F}_{0}$ is restricted to $\Delta^{d}$. Hence, we need to introduce additional notations of $\Delta^{d}$-valued ReLU networks. Let $L\in\mathbb{N}$, $J,S,M\geq 0$, and $\bm{d}=(d_{0},d_{\textup{NN},1},\cdots,d_{\textup{NN},L-1},d_{1})$. Let $H:\mathbb{R}^{d_{1}}\to\mathbb{R}^{d_{1}}$ be the softmax function,
$H(z)=(H_{1}(z),\cdots,H_{d_{1}}(z)),\quad H_{j}(z)=\frac{\exp(z_{j})}{\sum_{i=1}^{d_{1}}\exp(z_{i})}.$
Then, we define the class of $\Delta^{d}$-valued ReLU networks as
$\mathcal{F}_{L,J,S,M,\bm{d}}^{\Delta^{d}\textup{-NN}}=\left\{f_{\bm{W},\bm{b}}=\sum_{i=1}^{d_{1}}(H_{i}\circ g_{\bm{W},\bm{b}})v_{i}\;\bigg{|}\;g_{\bm{W},\bm{b}}\in\mathcal{F}_{L,J,S,M,\bm{d}}^{\textup{NN}}\right\}.$
In the main results, we follow (Petersen and Voigtlaender, , 2018; Imaizumi and Fukumizu, , 2019, 2022; Bos and Schmidt-Hieber, , 2022) to determine $L,J,S,M$, and $\bm{d}$.
4 Main Results
In Section 4.1 we show the main result on the $L^{2}$-risk. We give the highlight of the proof of the main result in Section 4.2. In Section 4.3 we give discussion of the main theorem and related work. In Section 4.4, we discuss to what extent the analysis for the local ERM can be generalized when the global ERM is used.
4.1 Main Theorem
In Algorithm 1, we give a summary of the algorithm shown in Section 3.
Algorithm 1 Local empirical risk minimization using contrastive learning
0:  Parameter $\beta>0$, admissible error $\varepsilon>0$, probability measure $P\in\mathcal{P}$,      number of samples $n\in\mathbb{N}$, number of layers in ReLU networks $L\in\mathbb{N}$,      number of parameters for ReLU networks $J,S,M\geq 0$, dimensions of      layers $\bm{d}=(d_{0},d_{\textup{NN},1},\cdots,d_{\textup{NN},L-1},d_{1})$ in the ReLU networks.
0:  $(\beta,\varepsilon,P,n,\mathcal{F}_{L,J,S,M,\bm{d}}^{\Delta^{d}\textup{-NN}})$-local ERM $\widehat{g}_{n}^{\textup{LERM}}$
1:  Observe $n$ samples $u_{1}=(x_{1},x_{1}^{\prime},y_{1}),\cdots,u_{n}=(x_{n},x_{n}^{\prime},y_{n})$ from $P$.
2:  Define the empirical risk $\frac{1}{n}\sum_{i=1}^{n}\ell_{f}(u_{i})$.
3:  Find the minimizer $\widehat{f}_{n}^{\textup{LERM}}$ of the empirical risk in the localized subclass $\mathscr{F}_{\beta,\beta^{-1}\varepsilon,P}(\mathcal{F}_{L,J,S,M,\bm{d}}^{\Delta^{d}\textup{-NN}})$.
4:  Define $\widehat{g}_{n}^{\textup{LERM}}=(\widehat{g}_{n,1}^{\textup{LERM}},\cdots,\widehat{g}_{n,d_{1}}^{\textup{LERM}})$ as $\widehat{f}_{n}^{\textup{LERM}}=\sum_{i=1}^{d_{1}}\widehat{g}_{n,i}^{\textup{LERM}}v_{i}$.
We are now ready to state the main theorem on the convergence rate of the ERM under the risk $\mathcal{R}(\cdot;P,U_{1}^{n})$ defined in (5).
Theorem 12.
Let $\alpha>0$, $P\in\mathcal{P}_{\alpha}$, $\beta\in(0,D_{{\textup{proj}}})$, $n\in\mathbb{N}\setminus\{1,2\}$ satisfying $\varepsilon_{n}=n^{-\alpha/(\alpha+d_{0}-1)}<2^{-1}$, and let $U_{1},\cdots,U_{n}$ be i.i.d. random variables drawn from $P$. Then, there are constants $C^{*}>0$ and $N\in\mathbb{N}$ independent of $n$, constant $c>0$, and some $L^{*}\in\mathbb{N}$, $J^{*},S^{*},M^{*}\geq 0$, and $\bm{d}^{*}\in\mathbb{N}^{L^{*}+1}$ satisfying
$L^{*}\lesssim\log_{2}{\varepsilon_{n}^{-1}}+1,\;J^{*}\lesssim\varepsilon_{n}^{-c}\vee 1,\;S^{*}\lesssim\varepsilon_{n}^{-(d_{0}-1)/\alpha}(1\vee\log_{2}{\varepsilon_{n}^{-1}}),$
$M^{*}\lesssim|\log(4d_{1}^{-2}\|p_{X}\|_{L^{\infty}(\mathcal{X})}^{-1}\varepsilon_{n})|\vee 1,\;and\;\bm{d}^{*}=(d_{0},d_{\textup{NN},1}^{*},\cdots,d_{\textup{NN},L^{*}-1}^{*},d_{1}),$
such that for the class $\mathcal{F}^{*}=\mathcal{F}_{L^{*},J^{*},S^{*},M^{*},\bm{d}^{*}}^{\Delta^{d}\textup{-NN}}$, if $n\geq N$, then the $(\beta,\varepsilon_{n},P,n,\mathcal{F}^{*})$-local ERM $\widehat{g}_{n}^{\textup{LERM}}$ satisfies
$\mathcal{R}(\widehat{g}_{n}^{\textup{LERM}};P,U_{1}^{n})\leq C^{*}n^{-\frac{\alpha}{\alpha+d_{0}-1}}\log^{4}{n}.$
The proof is deferred to Appendix A.7. In Section 4.2, we also give an outline and some key points in the proof. In Section 4.4, we show an extension of Theorem 12 in the setting where the global ERM is considered.
It is worth mentioning that in the context of nonparametric estimation of smooth boundaries, the convergence rate $n^{-\frac{\alpha}{\alpha+d_{0}-1}}$ is known to be optimal up to some logarithmic factor in the minimax sense in the literature (Mammen and Tsybakov, , 1999; Tsybakov, , 2004; Meyer, , 2023; Imaizumi and Fukumizu, , 2019, 2022). To the best of our knowledge, this is the first time that the rate $n^{-\frac{\alpha}{\alpha+d_{0}-1}}\log^{4}{n}$ is attained when a pairwise binary classification setting is considered. In Section 4.3 and Section 4.4, we discuss the differences to the previous work in detail.
4.2 Highlight of the Proof of Theorem 12
We shall give an overview of the proof of Theorem 12. The proof consists of several steps. In Section 4.2.1 we present the foundations of the proof. In Section 4.2.2 we show an estimation bound. Finally, in Section 4.2.3 we comment on the remained parts to complete the proof. The complete proofs are deferred to Appendix A.
4.2.1 Outline
Recall that for a given estimator $\widehat{f}_{n}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{F}_{0}$, the estimator $\widehat{g}_{n}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{G}_{0}$ of the indicator functions is uniquely defined as
$\widehat{f}_{n}=\sum_{i=1}^{d_{1}}\widehat{g}_{n,i}v_{i},$
where the sequence $\mathscr{S}_{P}$ has $d_{1}$ elements. In our analysis we wish to develop a bound measuring the gap between the estimator $\widehat{f}_{n}$ and contrastive function $f^{*}$ of some $P\in\mathcal{P}$, namely the risk $\mathbb{E}[\|\widehat{f}_{n}-f^{*}\|_{\mathcal{X},P_{X},2}^{2}]$. Indeed, once obtaining such a bound, by Proposition 6 one can directly transfer it to the bound of the risk $\mathcal{R}(\widehat{g}_{n};P,U_{1}^{n})$ in (5).
It is shown by Lecué, (2007, Proposition 1) that under the Tsybakov noise condition (Mammen and Tsybakov, , 1999; Tsybakov, , 2004), the excess risk of hinge loss gives an upper bound of the $L^{1}$-risk between classifiers. For completeness, we shall recall the fact in detail. To this end, we slightly abuse the notation $\ell$ to represent the standard loss function in supervised learning. Namely, for $\mathbb{R}$-valued function $g$ on $\mathcal{X}$ we write
$\ell_{g}(x,y)$ $=\max\{0,1-yg(x)\}.$
Lecué, (2007) proves the following fact:
Lemma 13 (Proposition 1 of Lecué, (2007)).
Let $\mathcal{X}_{0}$ be a measurable space. Let $g:\mathcal{X}_{0}\to[-1,1]$ be a measurable function on $\mathcal{X}_{0}$. Let $P_{0}$ be a probability measure in $\mathcal{X}_{0}\times\mathcal{Y}$ that has a probability density function on $\mathcal{X}_{0}\times\mathcal{Y}$, where the marginal distribution of $P_{0}$ on $\mathcal{X}_{0}$ is denoted by $P_{X_{0}}$. Also, denote by $g^{*}(x_{0})=\textup{sign}(2p(y=1|x_{0})-1)$, the Bayes classifier for $P_{0}$. Suppose that $P_{0}$ satisfies $\kappa$-(NC) for some $\kappa\geq 1$. Then, there is a positive constant $C_{0}$ such that we have
$\mathbb{E}_{P_{X_{0}}}[|g-g^{*}|]^{\kappa}\leq C_{0}(\mathbb{E}_{P_{0}}[\ell_{g}]-\mathbb{E}_{P_{0}}[\ell_{g^{*}}]).$
In general, the above property is known as Bernstein condition (Bartlett and Mendelson, , 2006) (see also (Steinwart and Scovel, , 2007; Alquier et al., , 2019)). Applying Proposition 1 in (Lecué, , 2007), we obtain an upper bound of the quantity $\mathbb{E}_{P_{X,X^{\prime}}}[|\psi\circ\rho_{f}-\psi\circ\rho_{f^{*}}|]$, $f\in\mathcal{F}_{0}$, where $f^{*}$ is the contrastive function of $P$. Hence, it is enough to show a lower bound of the quantity.
Nevertheless, it is not easy to derive a lower bound using the $L^{2}$-risk. In particular, Example 9 implies that some condition on vector-valued functions in $\mathcal{F}_{0}$ is required to exclude the functions introduced in the example. Hence, we consider a condition on function $f\in\mathcal{F}_{0}$ such that for any $x\in\mathcal{X}$,
$\|f(x)-f^{*}(x)\|_{2}<\beta,$
where $\beta<D_{\textup{proj}}$, and $f^{*}$ is the contrastive function of the given $P\in\mathcal{P}$. However, the above condition is not suitable when target functions are estimated by using neural networks, as we additionally need to consider the approximation errors. To address this issue, we use a probabilistic condition, namely localized subclasses defined in Definition 10.
To show a lower bound, we need to evaluate carefully the gap $|\psi\circ\rho_{f}-\psi\circ\rho_{f^{*}}|$ measured pointwisely in $\mathcal{X}\times\mathcal{X}$. The details are presented in the next subsection.
4.2.2 A General Estimation Bound
Here, we explain the details of the arguments presented in the last paragraph of Section 4.2.1. Specifically, we show that the term $\mathbb{E}_{P_{X,X^{\prime}}}[|\psi\circ\rho_{f}-\psi\circ\rho_{f^{*}}|]$ has some link to the $L^{2}$-risk $\mathbb{E}_{P_{X}}[\|f-f^{*}\|_{2}^{2}]$. For a given $P\in\mathcal{P}$, let us define
$P_{X,X^{\prime}}^{-}(dx,dx^{\prime})=p_{X}(x)p_{X^{\prime}}(x^{\prime})\mu(dx)\mu(dx^{\prime}).$
Lemma 14.
Let $P\in\mathcal{P}$, and let $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ be any sequence in $\mathscr{S}_{P}$. Denote by $f^{*}$, the contrastive function of $P$. Let $\beta\in(0,D_{{\textup{proj}}})$, $\beta_{0}\geq 0$, and $\mathcal{F}\subset\mathcal{F}_{0}$. For any distinct $i,j\in\{1,\cdots,d_{1}\}$ and every $f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$, we have
$P_{X,X^{\prime}}^{-}((\{x\in\mathcal{X}\;|\;\|f(x)-f^{*}(x)\|_{2}\geq\beta\}\cap\mathcal{K}_{i})\times\mathcal{K}_{j})\leq\beta_{0}.$
Proof.
By the definition of localized subclasses (Definition 10), the assertion holds. ∎
Then, we consider the following bound.
Lemma 15.
Let $\beta\in(0,D_{{\textup{proj}}})$, $\beta_{0}\geq 0$, $n\in\mathbb{N}\setminus\{1\}$, $\mathcal{F}\subset\mathcal{F}_{0}$, and $P\in\mathcal{P}$. Let $f^{*}$ denote the contrastive function of $P$, and let $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ be any sequence in $\mathscr{S}_{P}$. Then, there exists a constant $C_{2}>0$ such that for any $f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$ we have
$\mathbb{E}_{P_{X}}[\|f-f^{*}\|_{2}^{2}]$
$\leq C_{2}\left((D_{\Delta^{d}}-\beta)\beta_{0}+\sum_{i\neq j}\sum_{w=0}^{\lfloor\log_{2}{n}\rfloor}\left(\frac{1}{2}\right)^{2w+1}\beta^{2}P_{i,j}^{-}(2^{-(w+1)}\beta)+\frac{\beta^{2}}{n}\right),$ (6)
where $P_{i,j}^{-}(r)=P_{X,X^{\prime}}^{-}((\{x\in\mathcal{X}\;|\;\|f(x)-f^{*}(x)\|_{2}>r\}\cap\mathcal{K}_{i})\times\mathcal{K}_{j})$.
The proof is deferred to Appendix A.4. In the right-hand side of (15) we truncate the infinite series to prevent it from diverging. This argument makes it possible to proceed the subsequent analysis at the cost of additional factor $\log{n}$ in the final convergence rate.
Now, it remains to show an upper bound of $P_{i,j}^{-}(2^{-(w+1)}\beta)$, $i\neq j$ in (15). For any $r\in(0,2^{-1}\beta]$ we decompose the subset $\{x\in\mathcal{X}\;|\;\|f(x)-f^{*}(x)\|_{2}>r\}$ in
$\{x\in\mathcal{X}\;|\;r<\|f(x)-f^{*}(x)\|_{2}<\beta\},\textup{ and }$
$\{x\in\mathcal{X}\;|\;\beta\leq\|f(x)-f^{*}(x)\|_{2}\leq D_{\Delta^{d}}\}.$
The probability of the latter subset is evaluated by Lemma 14. Thus, we investigate the former subset.
Lemma 16.
Let $\beta\in(0,D_{{\textup{proj}}})$, $\beta_{0}\geq 0$, $\mathcal{F}\subset\mathcal{F}_{0}$, and $P\in\mathcal{P}$. Let $f^{*}$ be the contrastive function of $P$. Let $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ be any sequence in $\mathscr{S}_{P}$. For every $i,j\in\{1,\cdots,d_{1}\}$ such that $i\neq j$, any $r\in(0,2^{-1}\beta]$, and any $f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$, there exists a constant $C_{i,j}>0$ independent of $r$ such that
$P_{X,X^{\prime}}^{-}((\{x\in\mathcal{X}\;|\;r<\|f(x)-f^{*}(x)\|_{2}\leq\beta\}\cap\mathcal{K}_{i})\times\mathcal{K}_{j})$
$\leq C_{i,j}(r\wedge(D_{\Delta^{d}}(1-\beta/D_{{\textup{proj}}})))^{-2}\mathbb{E}_{P_{X,X^{\prime}}^{-}}[|\psi\circ\rho_{f}-\psi\circ\rho_{f^{*}}|]+\beta_{0}.$
The proof can be found in Appendix A.5. In the proof, the condition that $\|f(x)-f^{*}(x)\|_{2}<\beta$ with probability at least $1-\beta_{0}$ (Definition 10) is utilized in order to show lower bounds of the quantity $|\psi\circ\rho_{f}(x,x^{\prime})-\psi\circ\rho_{f^{*}}(x,x^{\prime})|$. This is the technical reason why we need the local condition introduced in Definition 10.
Incorporating the above lemmas in the approach shown in Section 4.2.1, we obtain the estimation bound of the risk $\mathcal{R}(\widehat{g}_{n};P,U_{1}^{n})$ in (5) for a general estimator. The complete proof is deferred to Appendix A.6.
Theorem 17.
Let $\mathcal{F}\subset\mathcal{F}_{0}$, $\beta\in(0,D_{{\textup{proj}}})$, $\beta_{0}\geq 0$, and $n\in\mathbb{N}\setminus\{1\}$. Let $P\in\mathcal{P}$, and let $U_{1},\cdots,U_{n}$ be i.i.d. random variables following $P$. Consider an arbitrary estimator $\widehat{f}_{n}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$, where $\widehat{f}_{n}=\sum_{i=1}^{d_{1}}\widehat{g}_{i}v_{i}$ for the uniquely determined function $(\widehat{g}_{1},\cdots,\widehat{g}_{d_{1}}):(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{G}_{0}$. Then, there are positive constants $C$, $C^{\prime}$, and $C^{\prime\prime}$ that are independent of $n$ such that we have
$\mathcal{R}(\widehat{g}_{n};P,U_{1}^{n})\leq C(\log{n})\mathbb{E}[\mathcal{E}(\widehat{f}_{n}(U_{1}^{n}))]+C^{\prime}\beta_{0}+\frac{C^{\prime\prime}}{n},$ (7)
where $\mathcal{E}(f)=\mathbb{E}_{P}[\ell_{f}]-\mathbb{E}_{P}[\ell_{f^{*}}]$ for any $f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$ denotes the excess risk for the contrastive function $f^{*}=\sum_{i=1}^{d_{1}}g_{i}^{*}v_{i}$ of $P$.
4.2.3 Further Analyses
The proof builds on Theorem 17 presented in Section 4.2.2, where we additionally need to analyze the excess risk. We show that the analyses developed in (Park, , 2009; Kim et al., , 2021) are applicable to a pairwise binary classification setting. In the proof, we also need to evaluate the approximation errors, where we follow similar arguments to the approximation theorems on deep neural networks with softmax activation developed by Bos and Schmidt-Hieber, (2022).
A remained technical issue is that we need to investigate to what extent one can reduce the value of $\beta_{0}$ in (7). In fact, the approximation theorems in (Petersen and Voigtlaender, , 2018; Imaizumi and Fukumizu, , 2019, 2022; Bos and Schmidt-Hieber, , 2022) are developed for the class $\mathcal{F}_{L,J,S,M,\bm{d}}^{\textup{NN}}$, while in our analysis we consider the approximation property of the localized subclass $\mathscr{F}_{\beta,\beta^{-1}\varepsilon_{n},P}(\mathcal{F}_{L,J,S,M,\bm{d}}^{\Delta^{d}\textup{-NN}})$. In Appendix A.7.3, we show that this issue can be resolved: specifically, we prove that some function approximating the true function within a small error belongs to a localized subclass (Proposition 40).
Combining all the steps, one can show the claim of Theorem 12. The complete proof can be found in Appendix A.7.
4.3 Discussion
We discuss the main theorem by comparing it to the results shown in related work.
4.3.1 Related Work on Statistical Learnability of Smooth Boundaries Using DNN
The statistical learnability of smooth boundaries using Deep Neural Networks (DNN), where Hölder continuous functions are used to define the smoothness of boundaries, is studied in (Kim et al., , 2021; Meyer, , 2023; Imaizumi and Fukumizu, , 2019, 2022). The main differences can be seen in five points, namely methods, types of data (either single-covariate data $(X,Z)$ or paired-covariate data $(X,X^{\prime},Y)$), how boundary estimation is carried out (namely the true function $f^{*}$), the risk function, and the convergence rates of upper bounds.
Table 1: A summary of comparison to the previous work on statistical learnability of smooth boundaries using deep neural networks, where the smoothness of boundaries in subsets $\{\mathcal{A}_{i}\}_{i}$ and $\mathcal{A}$ is defined by $\alpha$-Hölder continuous functions for some $\alpha\geq 0$. In each row, we summarize the information from the corresponding reference. “BC”, “MC”, “LS”, “BE” and “CL” are the abbreviations of “Binary Classification”, “Multiclass Classification”, “Least-Squares method”, “Bayes Estimation”, and “Contrastive Learning”, respectively. Regarding the column “Convergence rate”, the notation $\zeta=(d_{0}-1)/\alpha$ is used for convenience. The value $\kappa\geq 1$ is the parameter of the Tsybakov noise condition (Mammen and Tsybakov, , 1999; Tsybakov, , 2004), and Meyer, (2023) also shows results for $L^{2}$-risk under the condition (see also (Tsybakov, , 2004, Proposition 1)). In (Kim et al., , 2021), the true $f^{*}$ is defined with an affine transformation $A$. In (Imaizumi and Fukumizu, , 2019, 2022), they assume that $g_{i}\in\mathcal{C}^{\gamma,d_{0}}$. Note that Imaizumi and Fukumizu, (2019) show the convergence rates of both empirical risk minimizer and Bayes estimator, where they use deep ReLU networks. On the other hand, in (Imaizumi and Fukumizu, , 2022) they consider empirical risk minimization with deep neural networks defined with some more general activation functions.
Reference Algorithm Covariate True $f^{*}$ Risk Convergence rate
(Kim et al., , 2021, Thm. 3.2) BC Single $A\circ\mathds{1}_{\mathcal{A}}$ Excess risk $n^{-\frac{\kappa}{\zeta+2\kappa-1}}\log^{\frac{3\kappa}{\zeta+2\kappa-1}}{n}$
(Kim et al., , 2021, Thm. 5.1) MC Single $A\circ\mathds{1}_{\mathcal{A}_{i}},\forall i$ Excess risk $n^{-\frac{\kappa}{\kappa\zeta+2\kappa-1}}\log^{\frac{3\kappa}{\kappa\zeta+2\kappa-1}}{n}$
(Meyer, , 2023, Cor. 3.8) BC Single $\mathds{1}_{\mathcal{A}}$ Excess risk $n^{-\frac{s\kappa}{\zeta+2\kappa-1}}\log^{\frac{2s\kappa}{\zeta}}{n}$, $s\geq 1$
(Imaizumi and Fukumizu, , 2019, Thm. 1,2) LS & BE Single $\sum_{i}g_{i}\mathds{1}_{\mathcal{A}_{i}}$ $L^{2}$-risk $n^{-(\frac{2\gamma}{2\gamma+d_{0}}\wedge\frac{1}{\zeta+1})}\log^{2}{n}$
(Imaizumi and Fukumizu, , 2022, Thm. 7) LS Single $\sum_{i}g_{i}\mathds{1}_{\mathcal{A}_{i}}$ $L^{2}$-risk $n^{-(\frac{2\gamma}{2\gamma+d_{0}}\wedge\frac{1}{\zeta+1})}\log^{2}{n}$
This work CL Paired $\sum_{i}\mathds{1}_{\mathcal{A}_{i}}v_{i}$ $L^{2}$-risk $n^{-\frac{1}{\zeta+1}}\log^{4}{n}$
The summary of comparison is given in Table 1. In short, the convergence rate obtained in Theorem 12 seems to be consistent with the results shown by Kim et al., (2021), Meyer, (2023), and Imaizumi and Fukumizu, (2019, 2022), while the most clear difference is that we use a contrastive learning algorithm that requires data observed in a pairwise binary classification setting.
Regarding the true function $f^{*}$ to be estimated, Imaizumi and Fukumizu, (2019, 2022) investigate a more general setting than ours. Hence, the convergence rate in Theorem 12 is natural since every term $\mathds{1}_{\mathcal{K}_{i}}$ in our target function $f^{*}=\sum_{i=1}^{d_{1}}\mathds{1}_{\mathcal{K}_{i}}v_{i}$ can be regarded as the function $1\cdot\mathds{1}_{\mathcal{K}_{i}}$, where the coefficient $1$ is $\gamma$-Hölder continuous for any $\gamma>0$. Formally, by Proposition 6 and the Jensen inequality, we have
$\left\|\sum_{i=1}^{d_{1}}g_{i}-g^{*}\right\|_{L^{2}(\mathcal{X},P_{X})}^{2}\leq(d_{1}-1)\|f-f^{*}\|_{\mathcal{X},P_{X},2}^{2}=d_{1}\sum_{i=1}^{d_{1}}\|g_{i}-g_{i}^{*}\|_{L^{2}(\mathcal{X},P_{X})}^{2}.$
where $g\in\mathcal{G}_{0}$, $f=\sum_{i=1}^{d_{1}}g_{i}v_{i}\in\mathcal{F}_{0}$, and $g^{*}=\sum_{i=1}^{d_{1}}\mathds{1}_{\mathcal{K}_{i}}$. Although all the functions $g_{1}^{*},\cdots,g_{d_{1}}^{*}$, and $g^{*}$ belong to a class of regression functions studied in (Imaizumi and Fukumizu, , 2022), it is not possible to apply the theorems shown by Imaizumi and Fukumizu, (2019, 2022) since no label $Z$ is available in the pairwise binary classification setting. Instead, Theorem 12 shows that deep ReLU networks trained with such pairwise relation can estimate the smooth boundaries in $\mathcal{X}$ without using any label $Z$.
Several binary classification problems where decision boundaries are represented by some smooth functions, and the statistical error is measured by the excess misclassification risk, are investigated in (Kim et al., , 2021; Meyer, , 2023). The mathematical relation between the excess misclassification risk and the $L^{1}$-risk of classifiers is shown by Tsybakov, (2004, Proposition 1) (see also (Meyer, , 2023, Corollary 3.8)). Hence, it is not surprising that the result in Theorem 12 naturally aligns with the analyses of Kim et al., (2021) and Meyer, (2023). Nevertheless, it is impossible to directly apply their arguments to our problem setting because one needs to do some additional investigation of the relation between the quantities $\|f(x)-f^{*}(x)\|_{2}$ and $|\rho_{f}(x,x^{\prime})-\rho_{f^{*}}(x,x^{\prime})|$, $x,x^{\prime}\in\mathcal{X}$, $f\in\mathcal{F}_{0}$. To overcome the technical difficulty, we introduced a localized subclass in Definition 10 and studied the additional step utilizing a probabilistic condition on ERM. On the other hand, in (Kim et al., , 2021; Meyer, , 2023) a deterministic argument is considered, namely the setting where every function belongs to a closed neighborhood of any Bayes classifier with respect to the uniform norm.
4.3.2 Related Work on Pairwise Binary Classification
Related work on similarity learning.
Theorem 12 shares some similarity to the claim shown by Bao et al., 2022b in the context of similarity learning. Bao et al., 2022b (, Theorem 1) theoretically show that similarity learning can learn the decision boundary of a binary classification problem. Hence, the claim implied from Theorem 12 aligns with the result of Bao et al., 2022b . Furthermore, in (Bao et al., 2022b, ) they consider the setting where the random variables $X$ and $X^{\prime}$ are sampled independently. On the other hand, in our theory we consider the setting that $X$ and $X^{\prime}$ can be dependent, as in (Tsai et al., , 2020). In addition, Theorem 12 newly investigates the nonparametric estimation perspective of the problem, focusing on estimating multiple smooth boundaries.
The theoretical performance of deep neural networks in similarity learning involving nonparametric estimation has recently been investigated in (Zhou et al., , 2024). Zhou et al., (2024) show several upper bounds of the excess risk of similarity learning in the setting where conditional probability functions belong to the Sobolev space. Meanwhile, in our work, we are mainly interested in the $L^{2}$-risk (5). Furthermore, we focus on the smoothness of boundaries, different from (Zhou et al., , 2024). For the mathematical differences on the smoothness of target functions among classification problems, we refer the readers to (Audibert and Tsybakov, , 2007; Kim et al., , 2021).
Related work on contrastive learning.
The theory of contrastive learning is extensively studied in the literature (see, e.g., (Arora et al., , 2019; HaoChen et al., , 2021; Wang et al., , 2022; Bao et al., 2022a, ; Saunshi et al., , 2022; Huang et al., , 2023; Waida et al., , 2023)). In particular, several properties of the population risk minimizers of contrastive learning algorithms are investigated in (Wang and Isola, , 2020; HaoChen et al., , 2021; Awasthi et al., , 2022; Parulekar et al., , 2023; Johnson et al., , 2023; Zhai et al., , 2024; Koromilas et al., , 2024). However, the estimation performance in terms of the $L^{2}$-risk is less studied.
In the context of contrastive learning, the work by Tosh et al., 2021a ; Tosh et al., 2021b is relevant to our analyses. Indeed, in (Tosh et al., 2021a, , Theorem 4) and (Tosh et al., 2021b, , Theorem 11), they show several estimation error bounds of contrastive learning defined with some specific classification losses, where they consider the setting in which the estimation error is measured by downstream regression tasks and derive upper bounds in terms of an excess risk. On the other hand, in Theorem 12 the estimation error is measured by the $L^{2}$-risk for $\Delta^{d}$-valued functions, which enables to study the performance of boundary estimation. Furthermore, in Theorem 12 the convergence rate is shown.
Some remarks on the technical approach.
In the proof of Theorem 12, we show an inequality between two quantities, namely $\|f-f^{*}\|_{\mathcal{X},P_{X},2}$ and $\mathbb{E}_{P_{X,X^{\prime}}}[|\psi\circ\rho_{f}-\psi\circ\rho_{f^{*}}|]$. A similar inequality for a contrastive learning problem has recently been shown in (Ge et al., , 2024). Specifically, Ge et al., (2024) consider the setting where the paired covariates $(X,X^{\prime})$ are independent of each other to show an inequality in (Ge et al., , 2024, Lemma D.2). On the other hand, we need a largely different approach to show our inequality since we consider the problem setting where the covariates $X$ and $X^{\prime}$ are not necessarily independent, following (Tsai et al., , 2020). Since the independence assumption is used in (Ge et al., , 2024, Lemma D.2), our inequality is not shown by their approach. Furthermore, in the field of contrastive learning, it is common to assume that $X$ and $X^{\prime}$ can be dependent, both in practice and in theory (see, e.g, (Arora et al., , 2019; Chen et al., , 2020; HaoChen et al., , 2021)). Therefore, Theorem 12 also gives some insights of how we can address the technical difficulty, which is described with localization of function classes as shown in Section 4.2.2.
4.4 An Extended Analysis for the Global ERM
In this subsection, we show that the arguments developed in Theorem 12 for the local ERM can be extended to the analysis of the global ERM.
Similarly to Definition 11, the global ERM is defined as follows:
Definition 18 ($(n,\mathcal{F})$-global ERM).
Given $n\in\mathbb{N}\setminus\{1,2\}$ and $\mathcal{F}\subset\mathcal{F}_{0}$, define $\widehat{f}_{n}^{\textup{GERM}}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{F}$ as
$\widehat{f}_{n}^{\textup{GERM}}(u_{1},\cdots,u_{n})=\operatorname*{arg\;min}_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\ell_{f}(u_{i}),$
where $u_{1},\cdots,u_{n}\in\mathcal{X}^{2}\times\mathcal{Y}$. Then, the $(n,\mathcal{F})$-global ERM $\widehat{g}_{n}^{\textup{GERM}}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{G}_{0}$ is defined as
$\widehat{g}_{n}^{\textup{GERM}}=(\widehat{g}_{n,1}^{\textup{GERM}},\cdots,\widehat{g}_{n,d_{1}}^{\textup{GERM}})\textup{ such that }\widehat{f}_{n}^{\textup{GERM}}=\sum_{i=1}^{d_{1}}\widehat{g}_{n,i}^{\textup{GERM}}v_{i}.$
To state the result for the global ERM, we need to define a quantity similar to (5): given an estimator $\widehat{g}_{n}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{G}_{0}$, define
^$\widehat{\mathcal{R}}(\widehat{g}_{n};P,U_{1}^{n})=\sum_{i=1}^{d_{1}}\|\widehat{g}_{n,i}(U_{1},\cdots,U_{n})-\mathds{1}_{\mathcal{K}_{i}}\|_{L^{2}(\mathcal{X},P_{X})}^{2}.$
Then, the following result is obtained.
Corollary 19.
Let $\alpha>0$, $\beta\in(0,D_{{\textup{proj}}})$, $P\in\mathcal{P}_{\alpha}$, $\{\varepsilon_{n_{0}}\}_{n_{0}\in\mathbb{N}}$ be a sequence of non-negative real values such that $\varepsilon_{n_{0}}\geq n_{0}^{-\frac{\alpha}{\alpha+d_{0}-1}}$ for all $n_{0}\in\mathbb{N}$, and $n\in\mathbb{N}\setminus\{1,2\}$ be any natural number such that $\varepsilon_{n}<2^{-1}$. Let $U_{1},\cdots,U_{n}$ be any sequence of i.i.d. random variables drawn from the distribution $P$.
Then, there are constants $C^{*},c_{1},c_{2}>0$ independent of $n$ and some parameters $L^{*}\in\mathbb{N}$, $J^{*},S^{*},M^{*}\geq 0$, and $d^{*}=(d_{0},d_{\textup{NN},1}^{*},\cdots,d_{\textup{NN},L^{*}-1}^{*},d_{1})\in\mathbb{N}^{L^{*}+1}$ dependent to $n,\varepsilon_{n}$ such that for the class $\mathcal{F}^{*}=\mathcal{F}_{L^{*},J^{*},S^{*},M^{*},\bm{d}^{*}}^{\Delta^{d}\textup{-NN}}$, the $(n,\mathcal{F}^{*})$-global ERM $\widehat{g}_{n}^{\textup{GERM}}$ and corresponding vector-valued function $\widehat{f}_{n}^{\textup{GERM}}$ satisfy the following inequality: With probability at least
$1-c_{1}\exp(-128c_{2}n\varepsilon_{n}(\log_{2}\varepsilon_{n}^{-1})(\log^{2}{n}))-Q(\widehat{f}_{n}^{\textup{GERM}}(U_{1}^{n})\notin\mathscr{F}_{\beta,\beta^{-1}\varepsilon_{n},P}(\mathcal{F}^{*})),$
we have
^$\widehat{\mathcal{R}}(\widehat{g}_{n}^{\textup{GERM}};P,U_{1}^{n})\leq C^{*}\varepsilon_{n}(\log^{3}{n})(\log\varepsilon_{n}^{-1}).$
The proof is deferred to Appendix A.8. This corollary implies that the convergence rate cannot exceed $n^{-\frac{\alpha}{\alpha+d_{0}-1}}$. This observation is similar to the analysis of Imaizumi and Fukumizu, (2019, 2022). In addition, we can observe that the notion of localized subclasses regularizes the empirical risk minimization procedure so that ERM does not fail to estimate the true function.
We give several consequences of Corollary 19, focusing on additional conditions on the global ERM.
Remark 20.
Suppose that the setting introduced in Corollary 19 is considered. Let $\widehat{f}_{n}^{\textup{LERM}}$ and $\widehat{f}_{n}^{\textup{GERM}}$ be the $(\beta,n^{-\frac{\alpha}{\alpha+d_{0}-1}},P,n,\mathcal{F}^{*})$-local ERM used in Theorem 12 and the $(n,\mathcal{F}^{*})$-global ERM considered in Corollary 19, respectively. If $\widehat{f}_{n}^{\textup{GERM}}(U_{1}^{n})=\widehat{f}_{n}^{\textup{LERM}}(U_{1}^{n})$ holds $Q$-almost surely, then there are constants $C^{*},c_{1},c_{2}>0$ independent of $n$ such that, with probability at least $1-c_{1}\exp(-128c_{2}n\varepsilon_{n}(\log_{2}\varepsilon_{n}^{-1})(\log^{2}{n}))$ we have
^$\widehat{\mathcal{R}}(\widehat{g}_{n}^{\textup{GERM}};P,U_{1}^{n})\leq C^{*}n^{-\frac{\alpha}{\alpha+d_{0}-1}}\log^{4}{n}.$
Note that Theorem 12 is recovered in this case.
Remark 21.
Suppose that the setting in Corollary 19 is considered with the choice $\varepsilon_{n}=n^{-\frac{\alpha}{\alpha+d_{0}-1}}$. If $\limsup_{n\to\infty}Q(\widehat{f}_{n}^{\textup{GERM}}(U_{1}^{n})\notin\mathscr{F}_{\beta,\beta^{-1}\varepsilon_{n},P}(\mathcal{F}^{*}))=0$, then we have
^$\widehat{\mathcal{R}}(\widehat{g}_{n}^{\textup{GERM}};P,U_{1}^{n})=O_{\mathbb{P}}(n^{-\frac{\alpha}{\alpha+d_{0}-1}}\log^{4}{n}),$
where $O_{\mathbb{P}}(\cdot)$ denotes the rate of convergence in probability.
For instance, the condition in Remark 21 is satisfied for any sequence $\{\varepsilon_{n}\}_{n\in\mathbb{N}}$ such that $\varepsilon_{n}\geq n^{-\frac{\alpha}{\alpha+d_{0}-1}}$ if the global ERM $\widehat{f}_{n}^{\textup{GERM}}(U_{1}^{n})$ converges uniformly to the population risk minimizer $f^{*}$ for $Q$-almost every points in $\Omega$, namely, $\limsup_{n\to\infty}\sup_{x\in\mathcal{X}}\|\widehat{f}_{n}^{\textup{GERM}}(U_{1}^{n})(x)-f^{*}(x)\|_{2}=0$ almost surely. While this condition can be relaxed to some extent, doing so may require to consider the sequence $\varepsilon_{n}$ that determines the convergence rate. To attain good convergence rates, it seems that some careful model selection is required (for model selection, see, e.g., (Mohri et al., , 2018, Chapter 4)).
Additional discussion.
In conventional supervised learning settings, the consistency of global empirical risk minimization using deep neural networks is shown in the literature (Kim et al., , 2021; Meyer, , 2023; Imaizumi and Fukumizu, , 2019, 2022). However as shown in Example 9, in the setting where paired covariates are used, it is impossible to evaluate the $L^{2}$-risk in the whole function class $\mathcal{F}_{L,J,S,M,\bm{d}}^{\Delta^{d}\textup{-NN}}$. This gap is due to the fact that our problem setting is different from the conventional settings, which is a key point to use the localized argument presented in Section 4.2.2.
Corollary 19 shares some similarity with the theory of Mendelson, (2015, 2017). In particular, Mendelson, (2017) shows that localization of function classes plays a key role in measuring the statistical performance of ERM. Meanwhile, the mechanism is different: Mendelson, (2015, 2017) consider localized classes to deal with the complexity of models, while in the present work we use a localization argument to derive an inequality between two different types of quantities, that is, $\|f-f^{*}\|_{\mathcal{X},P_{X},2}$ and $\mathbb{E}_{P_{X,X^{\prime}}}[|\psi\circ\rho_{f}-\psi\circ\rho_{f^{*}}|]$ (see Section 4.2.1).
In the field of contrastive learning, the role of structural assumptions is studied in (Saunshi et al., , 2022; HaoChen and Ma, , 2023; Cabannes et al., , 2023; Parulekar et al., , 2023) to elucidate the fundamental principles of contrastive learning. This concept is first studied in (Saunshi et al., , 2022). It is shown by HaoChen and Ma, (2023) that there is a condition on function classes such that the population risk in a downstream task is bounded by a quantity that depends on the condition. Cabannes et al., (2023) study a trade-off in the design of self-supervised learning algorithms. Parulekar et al., (2023) find that a population risk minimizer is characterized by several properties of function classes. On the other hand, a novel insight derived from the main results in the present work is about how the localization of function classes improves the performance of ERM for estimating smooth boundaries in the pairwise binary classification setting.
5 Application to Multiclass Classification
The estimation problem we addressed in the previous sections is general in the sense that multiple subsets are estimated simultaneously. To show the usefulness of our analyses, we investigate the following problem:
Can one utilize the ERM defined in the previous sections to construct a consistent classifier in multiclass nonparametric classification?
In this section, we show a consequence of Theorem 12 for the non-asymptotic analysis of multiclass nonparametric classification.
Problem setting.
We briefly introduce the problem setting of multiclass classification. We consider to predict the class label $z\in\mathcal{Z}_{d_{1}}=\{1,\cdots,d_{1}\}$ of covariate $x$, using data drawn from a given distribution $P\in\mathcal{P}$ characterized by the partition $\mathscr{S}_{P}$.
To do so, we shall recall a condition of Kim et al., (2021), which controls the behavior of the distribution around the boundaries. Indeed, while the Tsybakov noise condition is initially introduced in (Mammen and Tsybakov, , 1999; Tsybakov, , 2004) under the settings where binary variables are considered, some extensions to multiclass cases are also studied in (Tarigan and van de Geer, , 2008; Kim et al., , 2021; Bos and Schmidt-Hieber, , 2022). In this analysis we focus on the version introduced by Kim et al., (2021).
Let $P_{X,Z_{d_{1}}}$ be a Borel probability measure in $\mathcal{X}\times\mathcal{Z}_{d_{1}}$ such that it has the probability density function $p(x,z)$ in $\mathcal{X}\times\mathcal{Z}_{d_{1}}$. With a slight abuse of notation, we write the marginal distribution of $P_{X,Z_{d_{1}}}$ as $P_{X}$. Let $\mathcal{T}_{d_{1}}$ be the class defined as
$\mathcal{T}_{d_{1}}=\{T:\mathcal{X}\to\mathcal{Z}_{d_{1}}\;|\;T\textup{ is measurable}\}.$
It is well known (see, e.g., (Tarigan and van de Geer, , 2008; Kim et al., , 2021; Mohri et al., , 2018)) that the Bayes classifier $z^{*}:\mathcal{X}\to\mathcal{Z}_{d_{1}}$ minimizing the risk $P_{X,Z_{d_{1}}}(T(x)\neq z)$ in the class $\mathcal{T}_{d_{1}}$ satisfies
$z^{*}(x)\in\operatorname*{arg\;max}_{i\in\mathcal{Z}_{d_{1}}}p(z=i|x).$
The condition of Kim et al., (2021) assumes that there are constants $\kappa\geq 1$, $c>0$, and $t_{0}\in(0,1]$ such that for any Bayes classifier $z^{*}$ of $P_{X,Z_{d_{1}}}$ and every $t\in(0,t_{0}]$,
$P_{X}\left(\left\{x\in\mathcal{X}\;\Big{|}\;\min_{i\in\mathcal{Z}_{d_{1}}\setminus\{z^{*}(x)\}}|p(z=i|x)-p(z=z^{*}(x)|x)|\leq t\right\}\right)\leq c\cdot t^{\frac{1}{\kappa-1}}.$ (8)
We say that the distribution $P_{X,Z_{d_{1}}}$ satisfies $\kappa$-(MNC) if it satisfies the noise condition (8) of Kim et al., (2021) with parameter $\kappa$, for convenience.
Classifiers.
Following (Arora et al., , 2019, Definition 2.1), we shall define a classifier using some vector-valued function in two steps: in the first step, for a function $h:\mathcal{X}\to\mathcal{S}^{d-1}$ we consider a classifier
$\Upsilon_{0}(h)(\cdot)=\min\{j\in\mathcal{Z}_{d_{1}}\;|\;\langle h(\cdot),v_{j}\rangle=\max_{i=1,\cdots,d_{1}}\langle h(\cdot),v_{i}\rangle\},$
[…]
$\Upsilon(\widehat{f}_{n})(u_{1},\cdots,u_{n})(x)$
=$=\begin{cases}\Upsilon_{0}(\frac{\widehat{f}_{n}(u_{1},\cdots,u_{n})}{\|\widehat{f}_{n}(u_{1},\cdots,u_{n})(\cdot)\|_{2}})(x)&\quad\textup{if }\widehat{f}_{n}(u_{1},\cdots,u_{n})(x)\neq\bm{0},\\ 1&\quad\textup{if }\widehat{f}_{n}(u_{1},\cdots,u_{n})(x)=\bm{0}.\end{cases}$
Analyses.
Here, for measurable subsets $\mathcal{K},\mathcal{K}^{\prime}\subset\mathcal{X}$, let $D_{P_{X}}(\mathcal{K},\mathcal{K}^{\prime})$ denote the volume of the symmetric difference measured by a probability measure $P_{X}$ in $\mathcal{X}$, namely
$D_{P_{X}}(\mathcal{K},\mathcal{K}^{\prime})=P_{X}((\mathcal{K}\cup\mathcal{K}^{\prime})\setminus(\mathcal{K}\cap\mathcal{K}^{\prime})).$
We note the following fact.
Lemma 22.
For any partition $\mathscr{S}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}\in\mathscr{P}_{\alpha}$, there is a Borel probability measure $P_{X,Z_{d_{1}}}$ in $\mathcal{X}\times\mathcal{Z}_{d_{1}}$ satisfying 1-(MNC) such that for any Bayes classifier $z^{*}$ minimizing $P_{X,Z_{d_{1}}}(T(x)\neq z)$ in $\mathcal{T}_{d_{1}}$, we have $\sum_{i=1}^{d_{1}}D_{P_{X}}((z^{*})^{-1}(i),\mathcal{K}_{i})=0$.
The proof is deferred to Appendix A.9. Based on the lemma above, we define the class
[…]
where the conditions (B1) and (B2) are defined as
  1. (B1)
    $P_{X,Z_{d_{1}}}$ satisfies $1$-(MNC).
  2. (B2)
    There is some $P\in\mathcal{P}_{\alpha}$ such that the marginal distributions of $P$ and $P_{X,Z_{d_{1}}}$ with respect to $\mathcal{X}$ coincide with each other, and for any partition $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ in $\mathscr{S}_{P}$, any Bayes classifier $z^{*}$ of $P_{X,Z_{d_{1}}}$ satisfies $D_{P_{X}}((z^{*})^{-1}(i),\mathcal{K}_{i})=0$.
Then, we have the following inequality between different risk functions.
Lemma 23.
Given a Borel probability measure $P_{X,Z_{d_{1}}}\in\mathscr{U}_{\alpha}$, let $P$ be an element of $\mathcal{P}$ such that it satisfies the condition (B2) for $P_{X,Z_{d_{1}}}$. For any estimator $\widehat{f}_{n}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{F}_{0}$, define $\widehat{z}_{n}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{T}_{d_{1}}$ as
$\widehat{z}_{n}=\Upsilon(\widehat{f}_{n}).$ (9)
Also, let $\widehat{g}_{n}=(\widehat{g}_{n,1},\cdots,\widehat{g}_{n,d_{1}}):(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{G}_{0}$ denote the estimator satisfying $\widehat{f}_{n}=\sum_{i=1}^{d_{1}}\widehat{g}_{n,i}v_{i}$. Then, there is a constant $C>0$ independent of $n$ such that we have
$\mathbb{E}[P_{X,Z_{d_{1}}}(\widehat{z}_{n}(x)\neq z)]-\inf_{z^{*}}P_{X,Z_{d_{1}}}(z^{*}(x)\neq z)\leq C\mathcal{R}(\widehat{g}_{n};P,U_{1}^{n}),$ (10)
where the infimum in the left-hand side is taken over all measurable maps from $\mathcal{X}$ to $\mathcal{Z}_{d_{1}}$, $U_{1}^{n}$ is any sequence of i.i.d. random variables drawn from $P$, and the expectation in the left-hand side of (10) is taken in terms of the distribution of $U_{1}^{n}$.
The proof is deferred to Appendix A.10.
Result.
We now show the result for the multiclass classification problem defined above.
Theorem 24.
Let $\alpha>0$, $\beta\in(0,D_{{\textup{proj}}})$, $\{\varepsilon_{n_{0}}\}_{n_{0}\in\mathbb{N}}$ be a sequence of non-negative real values such that $\varepsilon_{n_{0}}\geq n_{0}^{-\frac{\alpha}{\alpha+d_{0}-1}}$ for any $n_{0}\in\mathbb{N}$, and $n\in\mathbb{N}\setminus\{1,2\}$ be a natural number satisfying $\varepsilon_{n}<1/2$. Given $P_{X,Z_{d_{1}}}\in\mathscr{U}_{\alpha}$, let $P\in\mathcal{P}_{\alpha}$ be any probability measure satisfying the condition (B2). Let $U_{1},\cdots,U_{n}$ be any i.i.d. sequence of random variables drawn from $P$.
Then, there are constants $C^{*},c_{1},c_{2}>0$ independent of $n$, some $L^{*}\in\mathbb{N}$, $J^{*},S^{*},M^{*}\geq 0$, and $\bm{d}^{*}=(d_{0},d_{\textup{NN},1}^{*},\cdots,d_{\textup{NN},L^{*}-1}^{*},d_{1})\in\mathbb{N}^{L^{*}+1}$, such that for the class $\mathcal{F}^{*}=\mathcal{F}_{L^{*},J^{*},S^{*},M^{*},\bm{d}^{*}}^{\Delta^{d}\textup{-NN}}$, classifier $\widehat{z}_{n}^{\textup{GERM}}$ defined as in (9) using the $(n,\mathcal{F}^{*})$-global ERM $\widehat{g}_{n}^{\textup{GERM}}$, and corresponding vector-valued function $\widehat{f}_{n}^{\textup{GERM}}$, we have the following: With probability at least
$1-c_{1}\exp(-128c_{2}n\varepsilon_{n}(\log_{2}\varepsilon_{n}^{-1})(\log^{2}{n}))-Q(\widehat{f}_{n}^{\textup{GERM}}(U_{1}^{n})\notin\mathscr{F}_{\beta,\beta^{-1}\varepsilon_{n},P}(\mathcal{F}^{*})),$
we have
$P_{X,Z_{d_{1}}}(\widehat{z}_{n}^{\textup{GERM}}(x)\neq z)-\inf_{z^{*}}P_{X,Z_{d_{1}}}(z^{*}(x)\neq z)\leq C^{*}\varepsilon_{n}(\log^{3}{n})(\log\varepsilon_{n}^{-1}).$
Furthermore, in the same setting as above, let $\varepsilon_{n}=n^{-\frac{\alpha}{\alpha+d_{0}-1}}$. Then, there are positive constants $C^{*}>0$ and $N\in\mathbb{N}$ independent of $n$ and a class of deep ReLU networks $\mathcal{F}^{*}=\mathcal{F}_{L^{*},J^{*},S^{*},M^{*},\bm{d}^{*}}^{\Delta^{d}\textup{-NN}}$ with parameters $L^{*},J^{*},S^{*},M^{*},\bm{d}^{*}$, such that for the classifier $\widehat{z}_{n}^{\textup{LERM}}$ defined as in (9) with the $(\beta,\varepsilon_{n},P,n,\mathcal{F}^{*})$-local ERM $\widehat{g}_{n}^{\textup{LERM}}$, if $n\geq N$, then we have
$\mathbb{E}[P_{X,Z_{d_{1}}}(\widehat{z}_{n}^{\textup{LERM}}(x)\neq z)]-\inf_{z^{*}}P_{X,Z_{d_{1}}}(z^{*}(x)\neq z)\leq C^{*}n^{-\frac{\alpha}{\alpha+d_{0}-1}}\log^{4}{n}.$
Proof.
Combine Lemma 22, Lemma 23, and Theorem 12 (or Corollary 19). ∎
This result indicates that the general framework of nonparametric boundary estimation introduced in Section 3 can apply to an other learning problem.
Discussion.
In the context of multiclass nonparametric classification, the convergence rates for the excess misclassification risk are established in (Tarigan and van de Geer, , 2008; Kim et al., , 2021; Bos and Schmidt-Hieber, , 2022). The work by Kim et al., (2021) is closely related to our result. Theorem 5.1 in (Kim et al., , 2021) shows that a multiclass classifier defined with deep ReLU networks attains the convergence rate $((\log^{3}{n})/n)^{\frac{\kappa\alpha}{(2\kappa-1)\alpha+(d_{0}-1)\kappa}}$ under the excess risk, where $\kappa\in[1,\infty)$ represents the parameter of their noise condition, and they consider the case that the decision boundary is parameterized by some $\alpha$-Hölder continuous functions. Thus, the result of Kim et al., (2021) implies the convergence rate $n^{-\frac{\alpha}{\alpha+d_{0}-1}}$ up to a logarithmic factor when $\kappa=1$. The main difference is that Kim et al., (2021) consider the conventional supervised learning setting. On the other hand, we consider a pairwise binary classification setting, which requires a substantially different approach. Additionally, we use a different estimator.
We give comparison to the work by HaoChen et al., (2021). As a special case, HaoChen et al., (2021) show the learnability on multiclass classification in the setting where the label is a function of covariate, by using both a vector-valued function trained via contrastive learning and some linear classifier trained in a supervised downstream task, where the best convergence rate is $O_{\mathbb{P}}(n^{-\frac{1}{2}})$ in the result of (HaoChen et al., , 2021). The main differences to the analysis of HaoChen et al., (2021) are summarized in three points. First, we additionally assume the smoothness of the boundaries, following (Imaizumi and Fukumizu, , 2022). In addition, the condition on the probability distribution $P_{X,Z_{d_{1}}}$ used in our analysis is more general. Finally, Theorem 24 implies a faster rate when $\alpha>d_{0}-1$ and $n$ is sufficiently large, if the local ERM is used.
Recently, Duan et al., (2024) investigated transfer learning using contrastive learning and derived a convergence rate for multiclass classification. Our problem setting, analyses, and results are largely different from the ones of their work. In addition, the purpose of the present work is to study the statistical learnability of smooth boundaries, while Duan et al., (2024) study transfer learning.
6 Conclusion and Future Directions
We showed that it is possible to estimate the nonparametric smooth boundaries introduced by Imaizumi and Fukumizu, (2022), using data observed in a setting of pairwise binary classification. Following (Tsai et al., , 2020), we considered a setting where $Y$ is a binary variable representing whether $X$ and $X^{\prime}$ are statistically dependent or not. We developed a contrastive learning algorithm that outputs an empirical risk minimizer possessing some structural condition. The main result (Theorem 12) shows the statistical learnability of smooth boundaries using deep ReLU networks. We revealed that a localized structure of function classes is useful in this setting, which is similar to the statistical theory of Mendelson, (2015, 2017) but has a different mechanism. We also obtained several results for the global ERM under additional conditions (Corollary 19, Remark 20, and Remark 21). We finally applied the results to a multiclass nonparametric classification problem (Theorem 23). Therefore, statistical learning of smooth boundaries based on the pairwise dependence of covariates is meaningful to solve some specific learning problem.
Although the results obtained in this work are mainly theoretical, they can yield several implications to developing practical algorithms for learning smooth boundaries via pairwise binary classification. For instance, Corollary 19 implies that smooth boundaries are learnable with a pairwise binary classification algorithm if the learning procedure is regularized so that the global ERM chooses a good function in a localized class around the oracle (see also Remark 20 and Remark 21). Thus, an important future work is about how to perform model selection effectively.
We conclude this paper by mentioning several future directions.
  1. 1.
    Since the main focus of this work is about statistical learnability of smooth boundaries, we leave the investigation of minimax optimality as a future work (see, e.g., (Tsybakov, , 2009) for the minimax lower bounds). In fact, the convergence rate in Theorem 12 coincides with the minimax rate shown in (Imaizumi and Fukumizu, , 2019, 2022) up to a logarithmic factor, though the assumption on distributions in $\mathcal{P}_{\alpha}$ is different from the ones introduced in the literature (Mammen and Tsybakov, , 1999; Tsybakov, , 2004; Imaizumi and Fukumizu, , 2019, 2022; Meyer, , 2023). Specifically, we mainly consider the setting where the Tsybakov noise condition (Mammen and Tsybakov, , 1999; Tsybakov, , 2004) is imposed to every distribution $P$ in $\mathcal{P}$, while the smoothness condition introduced by Imaizumi and Fukumizu, (2022) is assumed in the space $\mathcal{X}$, not in the product space $\mathcal{X}^{2}$. Filling this gap may require some complicated analyses.
  2. 2.
    In Theorem 24, we do not consider the setting where the marginal distributions of given $P\in\mathcal{P}$ and $P_{X,Z_{d_{1}}}$ with respect to the variable $X$ can be different. As some similar settings are investigated in the context of transfer learning (see, e.g., (Cai and Wei, , 2021; Duan et al., , 2024)), investigation of this point is an interesting future work. For instance, it might be possible to analyze some transfer learning setting, building on the theoretical analysis of HaoChen et al., (2022).
  3. 3.
    In this work we use a contrastive learning algorithm to train the deep ReLU networks so that it learns the statistical dependence between covariates, following (Tsai et al., , 2020). Thus, an interesting topic is to investigate some practical implications to empirical aspects of contrastive learning, which is not the main purpose of this work. In addition, various learning algorithms are investigated in the field of self-supervised learning, including contrastive learning (see, e.g., (Ermolov et al., , 2021; Dufumier et al., , 2023; Huang et al., , 2023; Balestriero et al., , 2023)). In a future work, we aim at developing a general statistical theory that covers some self-supervised learning algorithms.
Appendix A Proofs
A.1 Proof of Proposition 3
Proof.
Denote by $\mathscr{S}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$. Let $P_{X}$ be an arbitrary Borel probability measure in $(\mathcal{X},\mathcal{B}(\mathcal{X}))$ such that $P_{X}$ is absolutely continuous for the Lebesgue measure $\mu$, and the Lebesgue density is continuous and positive at every point in $\mathcal{X}$. Let $B$ and $B^{\prime}$ be i.i.d. random variables drawn from a distribution on the set $\{1,\cdots,d_{1}\}$ such that $Q(B=i)=Q(B^{\prime}=i)=P_{X}(\mathcal{K}_{i})$ for any $i=1,\cdots,d_{1}$. Also, let $V_{1},\cdots,V_{d_{1}}$ and $V_{1}^{\prime},\cdots,V_{d_{1}}^{\prime}$ be i.i.d. random variables on $\Omega$ such that both $V_{i}$ and $V_{i}^{\prime}$ follow the conditional distribution $P_{X}(\cdot|\mathcal{K}_{i})=P_{X}(\cdot\cap\mathcal{K}_{i})/P_{X}(\mathcal{K}_{i})$ for every $i=1,\cdots,d_{1}$. Here, consider a random variable $\Gamma:\Omega\to\{0,1\}$ drawn from the Bernoulli distribution with parameter $c=Q(\Gamma=1)\in(0,1)$ satisfying
$c>\max_{i\in\{1,\cdots,d_{1}\}}\frac{P_{X}(\mathcal{K}_{i})}{1+P_{X}(\mathcal{K}_{i})}.$
Following (Arora et al., , 2019; Awasthi et al., , 2022), we define as
$Y$ $=2\mathds{1}_{\left\{\Gamma=1\right\}}-1,$
$(X,X^{\prime})|Y$ =$=\begin{cases}(\sum_{i=1}^{d_{1}}V_{i}\mathds{1}_{\{B=i\}},\sum_{i=1}^{d_{1}}V_{i}^{\prime}\mathds{1}_{\{B=i\}})&\quad\textup{if }Y=1,\\ (\sum_{i=1}^{d_{1}}V_{i}\mathds{1}_{\{B=i\}},\sum_{i=1}^{d_{1}}V_{i}^{\prime}\mathds{1}_{\{B^{\prime}=i\}})&\quad\textup{if }Y=-1.\end{cases}$
Note that the distribution of $(X,X^{\prime})|Y$ is indeed an example of joint distributions introduced by Arora et al., (2019, equation (1), (2)). In particular, the above construction satisfies Assumption 3.1 in (Awasthi et al., , 2022). Hence, the above example expresses the setting developed in (Arora et al., , 2019; Awasthi et al., , 2022) as a pairwise binary classification problem. Thus, it suffices to show that the example satisfies (A1)–(A4).
By the construction, the distribution of $(X,X^{\prime},Y)$ satisfies both the conditions (A1) and (A3). In addition, for any $i\in\{1,\cdots,d_{1}\}$ and every $(x,x^{\prime})\in\mathcal{K}_{i}\times\mathcal{K}_{i}$, we have
$\eta(x,x^{\prime})=\frac{c}{(1-P_{X}(\mathcal{K}_{i}))c+P_{X}(\mathcal{K}_{i})}>\frac{1}{2}.$
Note that $\eta(x,x^{\prime})=0$ otherwise. Thus, the distribution satisfies the conditions (A2) and (A4). Therefore, the distribution of $(X,X^{\prime},Y)$, denoted by $P$, belongs to $\mathcal{P}_{\alpha}$. By the definition, we have $[\mathscr{S}]=\mathscr{S}_{P}$. ∎
A.2 Proof of Proposition 6
Proof.
Let $f_{1}=\sum_{i=1}^{d_{1}}g_{1,i}v_{i},f_{2}=\sum_{i=1}^{d_{1}}g_{2,i}v_{i}\in\mathcal{F}_{0}$. Conn et al., (2009, Corollary 2.6) show that
$\langle v_{i},v_{j}\rangle=-1/d\textup{ for any }i,j\in\{1,\cdots,d_{1}\}\textup{ such that }i\neq j.$ (11)
Then, we have
$\|f_{1}(x)-f_{2}(x)\|_{2}^{2}$
$=\left\|\sum_{i=1}^{d_{1}}(g_{1,i}(x)-g_{2,i}(x))v_{i}\right\|_{2}^{2}$
$=\sum_{i=1}^{d_{1}}|g_{1,i}(x)-g_{2,i}(x)|^{2}-\frac{1}{d}\sum_{\begin{subarray}{c}i,j=1,\cdots,d_{1},\\ i\neq j\end{subarray}}(g_{1,i}(x)-g_{2,i}(x))(g_{1,j}(x)-g_{2,j}(x)).$
$=\frac{d_{1}}{d}\sum_{i=1}^{d_{1}}|g_{1,i}(x)-g_{2,i}(x)|^{2}-\frac{1}{d}\left(\sum_{i=1}^{d_{1}}g_{1,i}(x)-\sum_{i=1}^{d_{1}}g_{2,i}(x)\right)^{2}$
$=\frac{d_{1}}{d}\sum_{i=1}^{d_{1}}|g_{1,i}(x)-g_{2,i}(x)|^{2}.$
Here, in the second equality we use (11). In the last inequality we use the fact $\sum_{i=1}^{d_{1}}g_{1,i}(x)=\sum_{i=1}^{d_{1}}g_{2,i}(x)=1$, which is due to the definition of $\mathcal{F}_{0}$. ∎
A.3 Proof of Proposition 8
The following condition guarantees that the Bayes classifier $\textup{sign}\circ(2\eta-1)$ is representable by some $f\in\mathcal{F}_{0}$.
Definition 25 ($(\psi_{0},\mathcal{F})$-representability).
Let $\psi_{0}:\mathbb{R}\to\mathbb{R}$ be a measurable function, and let $\mathcal{F}\subset\mathcal{F}_{0}$. A Borel probability measure $P\in\mathcal{P}$ is said as $(\psi_{0},\mathcal{F})$-representable if there is a vector-valued function $f\in\mathcal{F}$ such that $\psi_{0}\circ\rho_{f}=\textup{sign}\circ(2\eta-1)$ holds $P_{X,X^{\prime}}$-almost surely.
Recall that the function $\psi$ is defined in Definition 7.
Lemma 26.
Every probability measure $P$ in $\mathcal{P}$ is $(\psi,\mathcal{F}_{0})$-representable with the contrastive function $f^{*}$ of $P$.
Proof.
By Theorem 3 in (Alexander, , 1977), for the contrastive function $f^{*}$ (Definition 5) we have $\psi\circ\rho_{f^{*}}(x,x^{\prime})\in\{1,-1\}$. Hence, by the condition (A4) in Definition 1, we have
$\psi\circ\rho_{f^{*}}(x,x^{\prime})=\textup{sign}\circ(2\eta-1)(x,x^{\prime}),$ (12)
for any $(x,x^{\prime})\in\mathcal{X}\times\mathcal{X}$. This fact indicates that $P$ is $(\psi,\mathcal{F}_{0})$-representable. ∎
We now provide the proof of Proposition 8. To this end, we recall a well-known fact on hinge loss: Lin, (2002, Lemma 3.1) shows that a real-valued classifier minimizes expected hinge loss if and only if the classifier is the Bayes classifier (see also Section 3.3 in (Zhang, , 2004) and Example 4 in (Bartlett et al., , 2006)). Then, we can show the existence of the minimizers, i.e., Proposition 8.
Proof of Proposition 8.
By Lemma 3.1 in Lin, (2002) and Lemma 26, $\mathbb{E}_{P}[\ell_{f}]$ is minimized at any function $f^{*}$ representing the Bayes classifier in the sense of Definition 25. ∎
A.4 Proof of Lemma 15
Proof.
Noting the fact that $P_{X}(\mathcal{A})=P_{X,X^{\prime}}^{-}(\mathcal{A}\times\mathcal{X})$ for any measurable $\mathcal{A}\subseteq\mathcal{X}$, we have
$\mathbb{E}_{P_{X}}[\|f-f^{*}\|_{2}^{2}]$
[…]
$\leq\int_{0}^{\infty}\sum_{i\neq j}P_{i,j}^{-}(r^{1/2})dr$
$\quad+\int_{0}^{\infty}\sum_{i=1}^{d_{1}}P_{X,X^{\prime}}^{-}((\{x\in\mathcal{X}\;|\;\|f(x)-f^{*}(x)\|_{2}^{2}>r\}\times\mathcal{K}_{i})\times\mathcal{K}_{i})dr$
$\leq 2\int_{0}^{\infty}\sum_{i\neq j}rP_{i,j}^{-}(r)dr+\max_{i\in\{1,\cdots,d_{1}\}}P_{X^{\prime}}(\mathcal{K}_{i})\mathbb{E}_{P_{X}}[\|f-f^{*}\|_{2}^{2}],$ (13)
where the last inequality is obtained as follows: we note that
$\sum_{i=1}^{d_{1}}P_{X,X^{\prime}}^{-}((\{x\in\mathcal{X}\;|\;\|f(x)-f^{*}(x)\|_{2}^{2}>r\}\cap\mathcal{K}_{i})\times\mathcal{K}_{i})$
$=\sum_{i=1}^{d_{1}}P_{X^{\prime}}(\mathcal{K}_{i})P_{X}(\{x\in\mathcal{X}\;|\;\|f(x)-f^{*}(x)\|_{2}^{2}>r\}\cap\mathcal{K}_{i})$
$\leq\sum_{i=1}^{d_{1}}\max_{j\in\{1,\cdots,d_{1}\}}P_{X^{\prime}}(\mathcal{K}_{j})P_{X}(\{x\in\mathcal{X}\;|\;\|f(x)-f^{*}(x)\|_{2}^{2}>r\}\cap\mathcal{K}_{i})$
$=\max_{j\in\{1,\cdots,d_{1}\}}P_{X^{\prime}}(\mathcal{K}_{j})P_{X}(\{x\in\mathcal{X}\;|\;\|f(x)-f^{*}(x)\|_{2}^{2}>r\}),$
where in the last equality we note that $\mathcal{K}_{1},\cdots,\mathcal{K}_{d_{1}}$ are disjoint. Integrating both the sides in the above calculation yields (13). Here, we note that by the conditions (A3) and (A4) in Definition 1, we have $\max_{j\in\{1,\cdots,d_{1}\}}P_{X^{\prime}}(\mathcal{K}_{j})<1$. Combining this fact and (13), we obtain
$\mathbb{E}_{P_{X}}[\|f-f^{*}\|_{2}^{2}]\leq 2\left(1-\max_{j\in\{1,\cdots,d_{1}\}}P_{X^{\prime}}(\mathcal{K}_{j})\right)^{-1}\int_{0}^{\infty}\sum_{i\neq j}rP_{i,j}^{-}(r)dr.$ (14)
Let us set $C_{2}^{\prime}=2\left(1-\max_{j\in\{1,\cdots,d_{1}\}}P_{X^{\prime}}(\mathcal{K}_{j})\right)^{-1}$. We next show an upper bound for the integral in the right-hand side of (14). We can proceed as follows:
[…]
By Lemma 14, the third term in the right-hand side of (A.4) is evaluated as
[…]
For the first term in the right-hand side of (A.4), we have
[…]
Regarding the second term, we note the bound
[…]
$\leq\sum_{w=0}^{\lfloor\log_{2}{n}\rfloor}\left(\frac{1}{2}\right)^{2w+1}\beta^{2}P_{i,j}^{-}(2^{-(w+1)}\beta),$ (16)
where $\lfloor\log_{2}{n}\rfloor$ denotes the maximum integer which is not greater than $\log_{2}{n}$. Here, in (A.4), we used the fact that $P_{i,j}^{-}(r)$ is a non-increasing function of $r$. Define $C_{2}=C_{2}^{\prime}(1\vee(d+1)dD_{\Delta^{d}})$. Combining (14), (A.4), and (A.4), we obtain
$\mathbb{E}_{P_{X}}[\|f-f^{*}\|_{2}^{2}]$
$\leq C_{2}\left((D_{\Delta^{d}}-\beta)\beta_{0}+\sum_{i\neq j}\sum_{w=0}^{\lfloor\log_{2}{n}\rfloor}\left(\frac{1}{2}\right)^{2w+1}\beta^{2}P_{i,j}^{-}(2^{-(w+1)}\beta)+\frac{\beta^{2}}{n}\right),$
which coincides with the inequality we wish to show. ∎
A.5 Proof of Lemma 16
Figure 1: An illustration in the case where $d_{1}=3$, and both $z_{0}=f(x)$ and $z_{4}=f(x^{\prime})$ are in the simplex $\{z=\sum_{h\in\{i,j,h_{0}\}}c_{h}v_{h}\;|\;c_{i},c_{j},c_{h_{0}}\in[0,1],\;c_{i}+c_{j}+c_{h_{0}}=1\}$. Note that in this case we have $z_{1}=z_{2}$, though in general $z_{1}$ is not on the line segment connecting $v_{i}$ and $v_{h_{0}}$. We used NumPy (Harris et al., , 2020) and Matplotlib (Hunter, , 2007) to plot this figure.
Proof.
Let $(x,x^{\prime})\in\mathcal{X}^{2}$ be any point such that
$(x,x^{\prime})\in$ ($(\{x\in\mathcal{X}\;|\;r<\|f(x)-f^{*}(x)\|_{2}\leq\beta\}\cap\mathcal{K}_{i})$
×$\times(\mathcal{K}_{j}\cap\{x^{\prime}\in\mathcal{X}\;|\;\|f(x^{\prime})-f^{*}(x^{\prime})\|_{2}<\beta\}).$ (17)
Let $B_{i}=\{z\in\Delta^{d}\;|\;r<\|z-v_{i}\|_{2}\leq\beta\}$. We note that $f^{*}(x)=v_{i}$, which implies that $f(x)\in B_{i}$. Define $M_{i}^{*}(z^{\prime}):=\sup_{z\in B_{i}}\|z-z^{\prime}\|_{2}$ for every $z^{\prime}\in\Delta^{d}$. Then, we proceed as shown below. See also an illustration in Figure 1.
  • Let $z_{0}\in\Delta^{d}$ be $z_{0}=f(x)$.
  • We define $z_{1}=\sum_{h=1}^{d_{1}}c_{h}^{(1)}v_{h}\in\Delta^{d}$ such that $z_{1}$ satisfies
    $\begin{cases}c_{i}^{(1)}&=c_{i}^{(0)},\\ c_{j}^{(1)}&=0,\\ c_{h}^{(1)}&=c_{h}^{(0)}+\frac{1}{d_{1}-2}c_{j}^{(0)}\textup{ for every }h\in\{1,\cdots,d_{1}\}\setminus\{i,j\}.\end{cases}$
    Note that for any $h\in\{1,\cdots,d_{1}\}\setminus\{i,j\}$, we have
    $c_{h}^{(1)}=\frac{1}{d_{1}-2}(c_{h}^{(0)}+c_{j}^{(0)})+\frac{d_{1}-3}{d_{1}-2}c_{h}^{(0)}\leq 1.$
    Then, we have the following:
    Claim 27.
    We have $\|z_{0}-v_{j}\|_{2}\leq\|z_{1}-v_{j}\|_{2}$.
    Note that the proofs of all the claims are shown in this subsection.
  • Let $z_{2}=\sum_{h=1}^{d_{1}}c_{h}^{(2)}v_{h}\in\Delta^{d}$ be the vector such that
    $\begin{cases}c_{i}^{(2)}&=c_{i}^{(1)},\\ c_{j}^{(2)}&=0,\\ c_{h_{0}}^{(2)}&=1-c_{i}^{(1)},\\ c_{h}^{(2)}&=0\textup{ for any }h\in\{1,\cdots,d_{1}\}\setminus\{i,j,h_{0}\},\end{cases}$
    where $h_{0}\in\{1,\cdots,d_{1}\}\setminus\{i,j\}$ is arbitrary and fixed here. Then, we have the following:
    Claim 28.
    We have $\|z_{1}-v_{j}\|_{2}\leq\|z_{2}-v_{j}\|_{2}$.
  • Let $z_{3}=\sum_{h=1}^{d_{1}}c_{h}^{(3)}v_{h}$ be the vector such that $c_{h}^{(3)}=0$ for every $h\in\{1,\cdots,d_{1}\}\setminus\{i,h_{0}\}$, and
    $\|z_{3}-v_{i}\|_{2}=\begin{cases}r&\quad\textup{if }r\leq D_{\Delta^{d}}(1-\beta/D_{{\textup{proj}}}),\\ \beta&\quad\textup{otherwise}.\end{cases}$
    Then, by Pythagorean theorem applied in terms of the middle point of the line segment connecting $v_{i}$ and $v_{h_{0}}$, we obtain $\|z_{2}-v_{j}\|_{2}\leq\|z_{3}-v_{j}\|_{2}$.
Thus, we have
$\|z_{0}-v_{j}\|_{2}$ $\leq\|z_{3}-v_{j}\|_{2}$
$=M_{i}^{*}(v_{j})$
$=((\sqrt{3}D_{\Delta^{d}}/2)^{2}+((D_{\Delta^{d}}/2-r)\vee(\beta D_{\Delta^{d}}/D_{{\textup{proj}}}-D_{\Delta^{d}}/2))^{2})^{1/2}.$
Besides, let us consider the following argument:
  • Let $z_{4}=f(x^{\prime})$, and let $z_{5}\in\{z\in\Delta^{d}\;|\;\|z-v_{j}\|_{2}\leq\beta\}$ be the intersection of the topological boundary of $\Delta^{d}$ and the line containing both $f(x)$ and $f(x^{\prime})$. Then, we have $\|f(x)-z_{4}\|_{2}\leq\|f(x)-z_{5}\|_{2}$.
  • Consider the same procedure as the definitions of $z_{2}$ and $z_{3}$. Note that it suffices to consider the case that $r=0$. Thus, we have $\|f(x)-z_{4}\|_{2}\leq\|f(x)-v_{j}\|_{2}$.
Therefore, we have $\|f(x)-f(x^{\prime})\|_{2}\leq\|z_{0}-v_{j}\|_{2}\leq M_{i}^{*}(v_{j})$. Here denote by $M_{i,j}^{*}=M_{i}^{*}(v_{j})$ for convenience. Then, we have
$\rho_{f}(x,x^{\prime})\leq M_{i,j}^{*2}\leq(\sqrt{3}D_{\Delta^{d}}/2)^{2}+((D_{\Delta^{d}}/2-r)\vee(\beta D_{\Delta^{d}}/D_{{\textup{proj}}}-D_{\Delta^{d}}/2))^{2}.$ (18)
Let us consider the case that $D_{\Delta^{d}}/2-r\geq\beta D_{\Delta^{d}}/D_{{\textup{proj}}}-D_{\Delta^{d}}/2$. Let $r_{0}=(\sqrt{3}D_{\Delta^{d}}/2)^{2}+(D_{\Delta^{d}}/2-r)^{2}$. Since $\psi$ is a decreasing function, (18) implies $\psi\circ\rho_{f}(x,x^{\prime})\geq\psi(M_{i,j}^{*2})\geq\psi(r_{0})$. Here since $f^{*}(x)=v_{i}$ and $f^{*}(x^{\prime})=v_{j}$, by Theorem 3 in Alexander, (1977) we have $\psi\circ\rho_{f^{*}}(x,x^{\prime})=-1$. Thus, it holds that
$|\psi\circ\rho_{f}(x,x^{\prime})-\psi\circ\rho_{f^{*}}(x,x^{\prime})|\geq|\psi(M_{i,j}^{*2})-\psi\circ\rho_{f^{*}}(x,x^{\prime})|\geq 2-2\widetilde{C}_{i,j}r_{0}$ (19)
on every $(x,x^{\prime})$ satisfying (A.5), where $\widetilde{C}_{i,j}=D_{\Delta^{d}}^{-2}$. Here, the inequality $2-2\widetilde{C}_{i,j}r_{0}\geq\widetilde{C}_{i,j}^{\prime}r^{2}$ is satisfied for every $r\in[0,D_{\Delta^{d}}/2]$, where $\widetilde{C}_{i,j}^{\prime}=2\widetilde{C}_{i,j}$. Hence, we have
$|\psi\circ\rho_{f}(x,x^{\prime})-\psi\circ\rho_{f^{*}}(x,x^{\prime})|\geq\widetilde{C}_{i,j}^{\prime}r^{2}=\widetilde{C}_{i,j}^{\prime}(r\wedge(D_{\Delta^{d}}(1-\beta/D_{{\textup{proj}}})))^{2}.$ (20)
On the other hand, in the case that $D_{\Delta^{d}}/2-r<\beta D_{\Delta^{d}}/D_{{\textup{proj}}}-D_{\Delta^{d}}/2$, by the same arguments as (20), we have
$|\psi\circ\rho_{f}(x,x^{\prime})-\psi\circ\rho_{f^{*}}(x,x^{\prime})|$ $\geq\widetilde{C}_{i,j}^{\prime}(D_{\Delta^{d}}(1-\beta/D_{{\textup{proj}}}))^{2}$
$=\widetilde{C}_{i,j}^{\prime}(r\wedge(D_{\Delta^{d}}(1-\beta/D_{{\textup{proj}}})))^{2},$
where we note that $\beta D_{\Delta^{d}}/D_{\textup{proj}}-D_{\Delta^{d}}/2=D_{\Delta^{d}}/2-(D_{\Delta^{d}}-\beta D_{\Delta^{d}}/D_{{\textup{proj}}})$ and $D_{\Delta^{d}}-\beta D_{\Delta^{d}}/D_{{\textup{proj}}}<r\leq D_{\Delta^{d}}/2$ in this case. Therefore, taking a constant $\widetilde{C}_{i,j}^{\prime}$ to satisfy the above condition, we have
[…]
$\subseteq\{(x,x^{\prime})\in\mathcal{X}^{2}|\widetilde{C}_{i,j}^{\prime}(r\wedge(D_{\Delta^{d}}(1-\frac{\beta}{D_{{\textup{proj}}}})))^{2}\leq|\psi\circ\rho_{f}(x,x^{\prime})-\psi\circ\rho_{f^{*}}(x,x^{\prime})|\}.$ (21)
Therefore, we have
$\mathbb{E}_{P_{X,X^{\prime}}^{-}}[\mathds{1}_{(\{x\in\mathcal{X}\;|\;r<\|f(x)-f^{*}(x)\|_{2}\leq\beta\}\cap\mathcal{K}_{i})\times\mathcal{K}_{j}}]$
$\leq\mathbb{E}_{P_{X,X^{\prime}}^{-}}[\mathds{1}_{(\{x\in\mathcal{X}\;|\;r<\|f(x)-f^{*}(x)\|_{2}\leq\beta\}\cap\mathcal{K}_{i})\times(\mathcal{K}_{j}\cap\{x\in\mathcal{X}\;|\;\|f(x)-f^{*}(x)\|_{2}<\beta\})}]+\beta_{0}$
$\leq\mathbb{E}_{P_{X,X^{\prime}}^{-}}[\mathds{1}_{\{(x,x^{\prime})\in\mathcal{X}^{2}\;|\;\widetilde{C}_{i,j}^{\prime}(r\wedge(D_{\Delta^{d}}(1-\beta/D_{{\textup{proj}}})))^{2}\leq|\psi\circ\rho_{f}(x,x^{\prime})-\psi\circ\rho_{f^{*}}(x,x^{\prime})|\}}]+\beta_{0}$
$\leq\widetilde{C}_{i,j}^{\prime-1}(r\wedge(D_{\Delta^{d}}(1-\beta/D_{{\textup{proj}}})))^{-2}\mathbb{E}_{P_{X,X^{\prime}}^{-}}[|\psi\circ\rho_{f}-\psi\circ\rho_{f^{*}}|]+\beta_{0}$
$=C_{i,j}(r\wedge(D_{\Delta^{d}}(1-\beta/D_{{\textup{proj}}})))^{-2}\mathbb{E}_{P_{X,X^{\prime}}^{-}}[|\psi\circ\rho_{f}-\psi\circ\rho_{f^{*}}|]+\beta_{0},$
where the first inequality follows from Lemma 14, the second inequality is due to (A.5), and in the third inequality we used Markov’s inequality. In the last equality, we set $C_{i,j}=\widetilde{C}_{i,j}^{\prime-1}$. Thus, we obtain the claim. ∎
Proof of Claim 27.
It suffices to show the inequality for $\gamma\geq 0$:
$\quad|c_{i}^{(0)}|^{2}+|1-c_{j}^{(0)}|^{2}+\sum_{h\neq i,j}|c_{h}^{(0)}|^{2}$
$\leq|c_{i}^{(0)}|^{2}+|1-c_{j}^{(0)}+\gamma|^{2}+\sum_{h\neq i,j}|c_{h}^{(0)}+\frac{1}{d_{1}-2}\gamma|^{2}.$ (22)
Then, we obtain the claim by setting $\gamma=c_{j}^{(0)}$ and applying Proposition 6.
To show (A.5), it is enough to note that for any $\gamma\geq 0$, we have
$\gamma^{2}+2(1-c_{j}^{(0)})\gamma+\frac{1}{d_{1}-2}\gamma^{2}+\frac{2}{d_{1}-2}\sum_{h\neq i,j}c_{h}^{(0)}\gamma$
$=\frac{d_{1}-1}{d_{1}-2}\gamma^{2}+2(1-\frac{d_{1}-1}{d_{1}-2}c_{j}^{(0)}+\frac{1}{d_{1}-2}\sum_{h\neq i}c_{h}^{(0)})\gamma$
$=\frac{d_{1}-1}{d_{1}-2}\gamma^{2}+\frac{2}{d_{1}-2}((d_{1}-1)-(d_{1}-1)c_{j}^{(0)}-c_{i}^{(0)})\gamma\geq 0,$
where in the inequality we use $0\leq c_{i}^{(0)}+c_{j}^{(0)}\leq 1$ and $20\leq(d_{1}-2)c_{j}^{(0)}\leq d_{1}-$ 2. ∎
Proof of Claim 28.
By Proposition 6, we have
$\|z_{1}-v_{j}\|_{2}^{2}$ $=d_{1}d^{-1}(|c_{i}^{(1)}|^{2}+|1-c_{j}^{(1)}|^{2}+\sum_{h\neq i,j}|c_{h}^{(1)}|^{2}),$
$\|z_{2}-v_{j}\|_{2}^{2}$ $=d_{1}d^{-1}(|c_{i}^{(2)}|^{2}+|1-c_{j}^{(2)}|^{2}+\sum_{h\neq i,j}|c_{h}^{(2)}|^{2})$
$=d_{1}d^{-1}(|c_{i}^{(1)}|^{2}+|1-c_{j}^{(1)}|^{2}+\sum_{h\neq i,j}|c_{h}^{(2)}|^{2}),$
where the last equality is due to the definitions of $c_{i}^{(2)}$ and $c_{j}^{(2)}$, namely $c_{i}^{(2)}=c_{i}^{(1)}$ and $c_{j}^{(2)}=c_{j}^{(1)}=0$.
Hence, it suffices to show that $\sum_{h\neq i,j}|c_{h}^{(1)}|^{2}\leq\sum_{h\neq i,j}|c_{h}^{(2)}|^{2}$. However, we defined the sequence $\{c_{h}^{(2)}\}_{h\neq i,j}$ as $c_{h_{0}}^{(2)}>0$ and $c_{h}^{(2)}=0$ for any $h\in\{1,\cdots,d_{1}\}\setminus\{i,j,h_{0}\}$, which is an optimal solution of the maximization problem:
$\max_{\bm{c}\in[0,1]^{d_{1}-2}}\|\bm{c}\|_{2}^{2}\quad\textup{subject to}\quad\|\bm{c}\|_{1}=1-c_{i}^{(1)}.$
Therefore, we obtain the claim. ∎
A.6 Proof of Theorem 17
Proof.
Let $C_{2}$ and $C_{i,j}$, $i,j\in\{1,\cdots,d_{1}\}$ such that $i\neq j$, be the constants satisfying the conditions in Lemma 15 and Lemma 16, respectively. Applying Lemma 14 and Lemma 16 to Lemma 15, we have
$\mathbb{E}_{P_{X}}[\|f-f^{*}\|_{2}^{2}]$
$\leq C_{2}(D_{\Delta^{d}}-\beta)\beta_{0}$
$\quad+C_{2}\sum_{i\neq j}C_{i,j}\sum_{w=0}^{\lfloor\log_{2}{n}\rfloor}2^{-(2w+1)}\beta^{2}((2^{-(w+1)}\beta)\wedge(D_{\Delta^{d}}(1-\beta/D_{{\textup{proj}}})))^{-2}$
$\quad\cdot\mathbb{E}_{P_{X,X^{\prime}}^{-}}[|\psi\circ\rho_{f}-\psi\circ\rho_{f^{*}}|]+2C_{2}\sum_{i\neq j}\sum_{w=0}^{\infty}\left(\frac{1}{2}\right)^{2w+1}\beta^{2}\beta_{0}+C_{2}\frac{\beta^{2}}{n}$
$\leq C_{2}^{\prime}(\log{n})C_{2}\sum_{i\neq j}C_{i,j}\cdot\mathbb{E}_{P_{X,X^{\prime}}^{-}}[|\psi\circ\rho_{f}-\psi\circ\rho_{f^{*}}|]$
$\quad+C_{2}(D_{\Delta^{d}}-\beta+\frac{4}{3}d(d+1)\beta^{2})\beta_{0}+C_{2}\frac{\beta^{2}}{n},$
where $C_{2}^{\prime}=(\log{2})^{-1}(4+2\beta^{2}(D_{\Delta^{d}}(1-\beta/D_{{\textup{proj}}}))^{-2})$, and in the last inequality we use the following inequalities valid for any $n\in\mathbb{N}\setminus\{1\}$,
$\sum_{w=0}^{\lfloor\log_{2}{n}\rfloor}2^{-(2w+1)}\beta^{2}((2^{-(w+1)}\beta)\wedge(D_{\Delta^{d}}(1-\beta/D_{{\textup{proj}}})))^{-2}$
$\leq 4\log_{2}{n}+2(\log_{2}{n})\beta^{2}(D_{\Delta^{d}}(1-\beta/D_{{\textup{proj}}}))^{-2}$
$\leq(\log{2})^{-1}(4+2\beta^{2}(D_{\Delta^{d}}(1-\beta/D_{{\textup{proj}}}))^{-2})\log{n}.$
Note that $p_{Y}(-1)>0$ holds by the condition (A3) in Definition 1. Define
$C_{3}$ $=C_{2}^{\prime}C_{2}\sum_{i\neq j}C_{i,j}p_{Y}(-1)^{-1},$
$C_{4}$ $=C_{2}(D_{\Delta^{d}}-\beta+(4/3)d(d+1)\beta^{2}),$
$C_{5}$ $=C_{2}\beta^{2}.$
Then, we have
$\mathbb{E}_{P_{X}}[\|f-f^{*}\|_{2}^{2}]\leq C_{3}(\log{n})\mathbb{E}_{P_{X,X^{\prime}}}[|\psi\circ\rho_{f}-\psi\circ\rho_{f^{*}}|]+C_{4}\beta_{0}+\frac{C_{5}}{n}.$ (23)
Here, by Lemma 26, $\psi\circ\rho_{f^{*}}$ is the Bayes classifier. By Proposition 1 in (Lecué, , 2007) (see also Lemma 13), there is a constant $C_{0}>0$ such that
$\mathbb{E}_{P_{X,X^{\prime}}}[|\psi\circ\rho_{f}-\psi\circ\rho_{f^{*}}|]\leq C_{0}(\mathbb{E}_{P}[\ell_{f}]-\mathbb{E}_{P}[\ell_{f^{*}}]).$ (24)
Applying (24) to (23), we have
$\mathbb{E}_{P_{X}}[\|f-f^{*}\|_{2}^{2}]\leq C_{0}C_{3}(\log{n})(\mathbb{E}_{P}[\ell_{f}]-\mathbb{E}_{P}[\ell_{f^{*}}])+C_{4}\beta_{0}+\frac{C_{5}}{n}.$ (25)
Finally, set $f=\widehat{f}_{n}(U_{1}^{n}(\omega))$ with arbitrary $\omega\in\Omega$, and integrate (25) on $\Omega$. Then, by Proposition 6, we obtain the claim. ∎
A.7 Proof of Theorem 12
In this section, we prove Theorem 12. The proof consists of five steps:
  • The step 0 is presented as Theorem 17 in Appendix A.6.
  • The step 1 is to investigate the approximation error measured by the excess risk of hinge loss, following (Park, , 2009; Kim et al., , 2021) (Appendix A.7.1).
  • The step 2 is to approximate the true function $f^{*}$, accordingly to the notion defined in Step 1. We basically follow the arguments developed by Bos and Schmidt-Hieber, (2022) (Appendix A.7.2).
  • The step 3 is devoted to the analysis of the parameter $\beta_{0}$ in the definition of localized subclasses (Appendix A.7.3).
  • The step 4 completes the proof of Theorem 12 (Appendix A.7.4).
A.7.1 Step 1
We aim at deriving an upper bound of the excess risk. We consider to apply the results shown in (Park, , 2009; Kim et al., , 2021). To this end, we need to investigate the approximation error bounds.
First, we use a weaker notion of $(\psi_{0},\mathcal{F})$-representability:
Definition 29 ($(\psi_{0},\mathcal{F},\varepsilon,V)$-weak representability).
Given $\mathcal{F}\subseteq\mathcal{F}_{0}$, $\varepsilon\geq 0$, $V\in[0,1]$, and a measurable function $\psi_{0}:\mathbb{R}\to\mathbb{R}$, a Borel probability measure $P\in\mathcal{P}$ is said as $(\psi_{0},\mathcal{F},\varepsilon,V)$-weak representable, if there is a function $f\in\mathcal{F}$ such that the following inequality holds:
$\sup\left\{P_{X,X^{\prime}}(\mathcal{A})\;|\;\mathcal{A}\in\mathscr{W}_{f,\varepsilon}\right\}\geq V,$ (26)
where we define
$\mathscr{W}_{f,\varepsilon}=\left\{\mathcal{A}\in\mathcal{B}(\mathcal{X}^{2})\;|\;\|\psi_{0}\circ\rho_{f}-\textup{sign}\circ(2\eta-1)\|_{\mathcal{A},P_{X,X^{\prime}},1}\leq\varepsilon\right\}.$
Hereafter, in this section we define $\mathcal{F}\subseteq\mathcal{F}_{0}$, a measurable function $\psi_{0}:\mathbb{R}\to\mathbb{R}$, $\varepsilon\geq 0$, and $V\in[0,1]$.
Using the notion of the weak representability in Definition 29, we can evaluate the approximation error in terms of the excess risk of hinge loss, which will be used when applying the results of (Park, , 2009; Kim et al., , 2021).
Proposition 30.
Let $\mathcal{H}\subset\mathcal{F}_{0}$, $\varepsilon\geq 0$, $V\in[0,1]$, and $\psi$ be the function defined in Definition 7. For every $(\psi,\mathcal{H},\varepsilon,V)$-weak representable $P\in\mathcal{P}$, if there is a vector-valued function $f^{*}\in\mathcal{F}_{0}$ such that $\psi\circ\rho_{f^{*}}=\textup{sign}\circ(2\eta-1)$, $P_{X,X^{\prime}}$-almost surely, then there is a function $f\in\mathcal{H}$ such that we have
$\mathbb{E}_{P}[\ell_{f}]-\mathbb{E}_{P}[\ell_{f^{*}}]\leq\varepsilon+2(1-V).$
Proof.
Let $f_{0}\in\mathcal{H}$ be a function satisfying the condition of the weak representability,
$\sup\{P_{X,X^{\prime}}(\mathcal{A})\;|\;\mathcal{A}\in\mathscr{W}_{f_{0},\varepsilon}\}\geq V.$
Let $\delta>0$ be arbitrary. Then, there is a subset $\mathcal{A}\in\mathscr{W}_{f_{0},\varepsilon}$ such that $P_{X,X^{\prime}}(\mathcal{A})\geq V-\delta$ and $\|\psi\circ\rho_{f_{0}}-\textup{sign}\circ(2\eta-1)\|_{\mathcal{A},P_{X,X^{\prime}},1}\leq\varepsilon$. We have
$\mathbb{E}_{P}[\ell_{f_{0}}]-\mathbb{E}_{P}[\ell_{f^{*}}]$
[…]
$\leq\varepsilon+2P_{X,X^{\prime}}(\mathcal{X}\times\mathcal{X}\setminus\mathcal{A})$ (29)
$\leq\varepsilon+2(1-V)+2\delta.$
where in (27) we used the well-known fact that hinge loss is Lipschitz continuous (see, e.g., (Steinwart and Christmann, , 2008, Example 2.27)), and in (28) we used both the conditions that $\psi\circ\rho_{f^{*}}=\textup{sign}\circ(2\eta-1)$, $P_{X,X^{\prime}}$-almost surely, and the property that $\|\psi\circ\rho_{f_{0}}-\textup{sign}\circ(2\eta-1)\|_{\mathcal{A},P_{X,X^{\prime}},1}\leq\varepsilon$ holds. In (29), we use the triangle inequality and $|\psi\circ\rho_{f}|\leq 1$ on $\mathcal{X}^{2}$ for every $f\in\mathcal{F}_{0}$. Since $\delta$ is arbitrary, we obtain the claim. ∎
Next, we show several properties of weak representability.
Proposition 31.
If $P\in\mathcal{P}$ is $(\psi_{0},\mathcal{F})$-representable, then $P$ is $(\psi_{0},\mathcal{F},0,V)$-weak representable for any $V\in[0,1]$.
Proof.
By the definition of $(\psi_{0},\mathcal{F})$-representability, there exists a function $f\in\mathcal{F}$ such that the following identity holds:
$\|\psi_{0}\circ\rho_{f}-\textup{sign}\circ(2\eta-1)\|_{\mathcal{X}^{2},P_{X,X^{\prime}},1}=0.$
Using the notation
$\mathscr{W}_{f,\varepsilon,\psi_{0}}=\{\mathcal{A}\in\mathcal{B}(\mathcal{X}^{2})\;|\;\|\psi_{0}\circ\rho_{f}-\textup{sign}\circ(2\eta-1)\|_{\mathcal{A},P_{X,X^{\prime}},1}\leq\varepsilon\},$ (30)
it holds that
$\sup\{P_{X,X^{\prime}}(\mathcal{A})\;|\;\mathcal{A}\in\mathscr{W}_{f,0,\psi_{0}}\}=1\geq V.$
This shows the claim. ∎
We next see several useful properties of the weak representability:
Lemma 32.
Let $\mathcal{F}^{\prime}\subset\mathcal{F}_{0}$, $\varepsilon^{\prime}\geq 0$, and $V^{\prime}\in[0,1]$ be such that $\mathcal{F}^{\prime}\subseteq\mathcal{F}$, $\varepsilon^{\prime}\leq\varepsilon$, and $V^{\prime}\geq V$. Then, any $(\psi_{0},\mathcal{F}^{\prime},\varepsilon^{\prime},V^{\prime})$-weak representable $P\in\mathcal{P}$ is $(\psi_{0},\mathcal{F},\varepsilon,V)$-weak representable.
Proof.
For convenience, we use the notation $\mathscr{W}_{f,\varepsilon,\psi_{0}}$ of (30) in this proof. Let $P$ be an arbitrary $(\psi_{0},\mathcal{F}^{\prime},\varepsilon^{\prime},V^{\prime})$-weak representable distribution in $\mathcal{P}$. Recall that from the definition of $P$, there exists some $f\in\mathcal{F}^{\prime}$ such that
$\sup\{P_{X,X^{\prime}}(\mathcal{A})\;|\;\mathcal{A}\in\mathscr{W}_{f,\varepsilon^{\prime},\psi_{0}}\}\geq V^{\prime}.$
Here notice that from the condition that $\mathcal{F}^{\prime}\subseteq\mathcal{F}$, it holds that $f\in\mathcal{F}$. Besides, if an arbitrary $\mathcal{A}\in\mathcal{B}(\mathcal{X}^{2})$ satisfies the condition $\|\psi_{0}\circ\rho_{f}-\textup{sign}\circ(2\eta-1)\|_{\mathcal{A},P_{X,X^{\prime}},1}\leq\varepsilon^{\prime}$, then by the condition $\varepsilon^{\prime}\leq\varepsilon$ we have $\|\psi_{0}\circ\rho_{f}-\textup{sign}\circ(2\eta-1)\|_{\mathcal{A},P_{X,X^{\prime}},1}\leq\varepsilon$. Hence, we have
$\sup\{P_{X,X^{\prime}}(\mathcal{A})\;|\;\mathcal{A}\in\mathscr{W}_{f,\varepsilon,\psi_{0}}\}\geq\sup\{P_{X,X^{\prime}}(\mathcal{A})\;|\;\mathcal{A}\in\mathscr{W}_{f,\varepsilon^{\prime},\psi_{0}}\}\geq V^{\prime}.$
Since $V^{\prime}\geq V$ from the condition, we obtain
$\sup\{P_{X,X^{\prime}}(\mathcal{A})\;|\;\mathcal{A}\in\mathscr{W}_{f,\varepsilon,\psi_{0}}\}\geq V,$
which implies the assertion. ∎
The following lemma is a fundamental part in the proof of Proposition 34.
Lemma 33.
Let $\mathcal{F}\subseteq\mathcal{F}_{0}$, and let $P\in\mathcal{P}$. Let $\psi$ be the function defined in Definition 7. Suppose that $P$ is $(\psi,\mathcal{F},\varepsilon,V)$-weak representable, and $f\in\mathcal{F}$ satisfies the condition (26). Then for any subset $\mathcal{G}\subseteq\mathcal{F}$, $P$ is $(\psi,\mathcal{G},(2\varepsilon)\vee(32D_{\Delta^{d}}^{-1}\mu(\mathcal{X})\|p_{X,X^{\prime}}\|_{\mathcal{X}^{2},\infty}\varepsilon_{0}),V)$-weak representable, where $\varepsilon_{0}=\inf_{g\in\mathcal{G}}\|f-g\|_{\mathcal{X},1}$.
Proof.
In the proof we use the notation $\mathscr{W}_{f,\varepsilon,\psi}$ in (30). Let $\mathcal{A}\in\mathscr{W}_{f,\varepsilon,\psi}$. Let $g\in\mathcal{G}$ be arbitrary. Then we notice that
$\|\psi\circ\rho_{g}-\textup{sign}\circ(2\eta-1)\|_{\mathcal{A},P_{X,X^{\prime}},1}$
$\leq\|\psi\circ\rho_{f}-\psi\circ\rho_{g}\|_{\mathcal{A},P_{X,X^{\prime}},1}+\|\psi\circ\rho_{f}-\textup{sign}\circ(2\eta-1)\|_{\mathcal{A},P_{X,X^{\prime}},1}$
$\leq\|\psi\circ\rho_{f}-\psi\circ\rho_{g}\|_{\mathcal{A},P_{X,X^{\prime}},1}+\varepsilon,$ (31)
where in the second inequality we used the condition $\mathcal{A}\in\mathscr{W}_{f,\varepsilon,\psi}$. Let us bound the first term in the last line of the above inequality. It holds that
$\|\psi\circ\rho_{f}-\psi\circ\rho_{g}\|_{\mathcal{A},P_{X,X^{\prime}},1}$
$=\int_{\mathcal{A}}|\psi\circ\rho_{f}(x,x^{\prime})-\psi\circ\rho_{g}(x,x^{\prime})|P_{X,X^{\prime}}(dx,dx^{\prime})$
$\leq 2D_{\Delta^{d}}^{-2}\int_{\mathcal{X}\times\mathcal{X}}|\rho_{f}(x,x^{\prime})-\rho_{g}(x,x^{\prime})|P_{X,X^{\prime}}(dx,dx^{\prime})$
$\leq 2D_{\Delta^{d}}^{-2}\|p_{X,X^{\prime}}\|_{\mathcal{X}^{2},\infty}\|\rho_{f}-\rho_{g}\|_{\mathcal{X}^{2},\mu\otimes\mu,1},$
where in the first inequality we used the definition of $\psi$ and $\mathcal{A}\subset\mathcal{X}\times\mathcal{X}$, and in the last inequality we used Hölder’s inequality. Besides, we also notice that for any $(x,x^{\prime})\in\mathcal{X}^{2}$,
$\quad|\rho_{f}(x,x^{\prime})-\rho_{g}(x,x^{\prime})|$
$\leq(\|f(x)-f(x^{\prime})\|_{2}+\|g(x)-g(x^{\prime})\|_{2})\cdot(\|f(x)-g(x)\|_{2}+\|f(x^{\prime})-g(x^{\prime})\|_{2})$ (32)
$\leq 2D_{\Delta^{d}}(\|f(x)-g(x)\|_{2}+\|f(x^{\prime})-g(x^{\prime})\|_{2}),$ (33)
where we use in (A.7.1) and (33) the triangle inequality. Thus,
$\|\rho_{f}-\rho_{g}\|_{\mathcal{X}^{2},\mu\otimes\mu,1}\leq 4D_{\Delta^{d}}\mu(\mathcal{X})\|f-g\|_{\mathcal{X},1}.$
Using this fact, we have
$\|\psi\circ\rho_{f}-\psi\circ\rho_{g}\|_{\mathcal{A},P_{X,X^{\prime}},1}\leq 8D_{\Delta^{d}}^{-1}\mu(\mathcal{X})\|p_{X,X^{\prime}}\|_{\mathcal{X}^{2},\infty}\|f-g\|_{\mathcal{X},1}.$ (34)
Let $g_{0}\in\mathcal{G}$ be a function such that $\|f-g_{0}\|_{\mathcal{X},1}\leq 2\varepsilon_{0}$. By (34), we have
$\|\psi\circ\rho_{f}-\psi\circ\rho_{g_{0}}\|_{\mathcal{A},P_{X,X^{\prime}},1}\leq 16D_{\Delta^{d}}^{-1}\mu(\mathcal{X})\|p_{X,X^{\prime}}\|_{\mathcal{X}^{2},\infty}\varepsilon_{0}.$ (35)
From (31) and (35), we have
$\|\psi\circ\rho_{g_{0}}-\textup{sign}\circ(2\eta-1)\|_{\mathcal{A},P_{X,X^{\prime}},1}\leq(2\varepsilon)\vee(32D_{\Delta^{d}}^{-1}\mu(\mathcal{X})\|p_{X,X^{\prime}}\|_{\mathcal{X}^{2},\infty}\varepsilon_{0})$
Let $\varepsilon_{1}=(2\varepsilon)\vee(32D_{\Delta^{d}}^{-1}\mu(\mathcal{X})\|p_{X,X^{\prime}}\|_{\mathcal{X}^{2},\infty}\varepsilon_{0})$. Therefore, we have
$\sup\{P_{X,X^{\prime}}(\mathcal{A})\;|\;\mathcal{A}\in\mathscr{W}_{g_{0},\varepsilon_{1},\psi}\}\geq\sup\{P_{X,X^{\prime}}(\mathcal{A})\;|\;\mathcal{A}\in\mathscr{W}_{f,\varepsilon,\psi}\}\geq V.$
This indicates that $P$ is $(\psi,\mathcal{G},(2\varepsilon)\vee(32D_{\Delta^{d}}^{-1}\mu(\mathcal{X})\|p_{X,X^{\prime}}\|_{\mathcal{X}^{2},\infty}\varepsilon_{0}),V)$-weak representable. ∎
Proposition 34.
Let $\mathcal{H}\subset\mathcal{F}_{0}$, and $\psi$ be the function defined in Definition 7. For every $P\in\mathcal{P}$, there is some $C_{1}>0$ such that $P$ is $(\psi,\mathcal{H},C_{1}\varepsilon_{\mathcal{H},f^{*}},1)$-weak representable, where $f^{*}=\sum_{i=1}^{d_{1}}g_{i}^{*}v_{i}$ is the contrastive function of $P$, and $\varepsilon_{\mathcal{H},f^{*}}$ is defined as
$\varepsilon_{\mathcal{H},f^{*}}=\inf_{f=\sum_{i=1}^{d_{1}}g_{i}v_{i}\in\mathcal{H}}\max_{i=1,\cdots,d_{1}}\|g_{i}-g_{i}^{*}\|_{L^{1}(\mathcal{X})}.$ (36)
Proof.
By Lemma 26 and Proposition 31, $P$ is $(\psi,\mathcal{F}_{0},0,V)$-weak representable for any $V\in[0,1]$.
Here let $f_{1},f_{2}\in\mathcal{F}_{0}$ be arbitrary, and denote by $f_{1}=\sum_{i=1}^{d_{1}}g_{i}^{(1)}v_{i}$ and $f_{2}=\sum_{i=1}^{d_{1}}g_{i}^{(2)}v_{i}$. Then, we have
$\|f_{1}-f_{2}\|_{\mathcal{X},1}$ $\leq d^{\frac{1}{2}}\|\|f_{1}(x)-f_{2}(x)\|_{2}\|_{L^{1}(\mathcal{X})}$ (37)
$=d^{\frac{1}{2}}\left\|\left\|\sum_{i=1}^{d_{1}}g_{i}^{(1)}(x)v_{i}-\sum_{i=1}^{d_{1}}g_{i}^{(2)}(x)v_{i}\right\|_{2}\right\|_{L^{1}(\mathcal{X})}$
$\leq d^{\frac{1}{2}}\left\|\sum_{i=1}^{d_{1}}|g_{i}^{(1)}(x)-g_{i}^{(2)}(x)|\left\|v_{i}\right\|_{2}\right\|_{L^{1}(\mathcal{X})}$ (38)
$\leq d^{\frac{1}{2}}\sum_{i=1}^{d_{1}}\left\|g_{i}^{(1)}-g_{i}^{(2)}\right\|_{L^{1}(\mathcal{X})},$ (39)
where in (37) we used the Cauchy-Schwarz inequality, and in (38) and (39) we used the triangle inequality.
By Lemma 33, $P$ is $(\psi,\mathcal{H},32D_{\Delta^{d}}^{-1}\mu(\mathcal{X})\|p_{X,X^{\prime}}\|_{\mathcal{X}^{2},\infty}\varepsilon_{0},1)$-weak representable, where $\varepsilon_{0}=\inf_{f\in\mathcal{H}}\|f-f^{*}\|_{\mathcal{X},1}$. Denote by $f^{*}=\sum_{i=1}^{d_{1}}g_{i}^{*}v_{i}$. Then, by (39) we have
$\varepsilon_{0}\leq d^{\frac{1}{2}}d_{1}\inf_{f=\sum_{i=1}^{d_{1}}g_{i}v_{i}\in\mathcal{H}}\max_{i=1,\cdots,d_{1}}\|g_{i}-g_{i}^{*}\|_{L^{1}(\mathcal{X})}=d^{\frac{1}{2}}d_{1}\varepsilon_{\mathcal{H},f^{*}}.$
Here, let $C_{1}=32d^{\frac{1}{2}}d_{1}D_{\Delta^{d}}^{-1}\mu(\mathcal{X})\|p_{X,X^{\prime}}\|_{\mathcal{X}^{2},\infty}$. By Lemma 32, we have that $P$ is $(\psi,\mathcal{H},C_{1}\varepsilon_{\mathcal{H},f^{*}},1)$-weak representable. This shows the claim. ∎
A.7.2 Step 2
We then show the approximation error bound for the indicator functions $g_{i}^{*}=\mathds{1}_{\mathcal{K}_{i}}$, $i=1,\cdots,d_{1}$. Note that in this section, the notation $\lesssim$ abbreviates a coefficient which is a universal constant independent of the error $\varepsilon>0$.
We start by recalling a useful inequality shown by Bos and Schmidt-Hieber, (2022).
Lemma 35 (The inequality of Bos and Schmidt-Hieber, (2022) for $\|\cdot\|_{L^{1}(\mathcal{X})}$).
Let $g=(g_{1},\cdots,g_{d_{1}}):\mathcal{X}\to\mathbb{R}^{d_{1}}$ be a function such that $\|g\|_{\mathcal{X},\infty}\leq M$ for some $M\geq 0$. Let $g^{*}:\mathcal{X}\to\mathbb{R}^{d_{1}}$, $g^{*}=(g_{1}^{*},\cdots,g_{d_{1}}^{*})$, be a function satisfying $\|g^{*}(\cdot)\|_{1}=1$ on $\mathcal{X}$. Then, for any $i=1,\cdots,d_{1}$, it holds that
$\left\|(H\circ g)_{i}-g_{i}^{*}\right\|_{L^{1}(\mathcal{X})}\leq\left\|\exp\circ{g_{i}}-g_{i}^{*}\right\|_{L^{1}(\mathcal{X})}+\sum_{j=1}^{d_{1}}\left\|\exp\circ{g_{j}}-g_{j}^{*}\right\|_{L^{1}(\mathcal{X})},$
where $(H\circ g)_{i}$ denotes the $i$-th component of the vector-valued function $H\circ g$.
Proof.
This inequality is shown in Lemma 4.3 of (Bos and Schmidt-Hieber, , 2022). Note that in (Bos and Schmidt-Hieber, , 2022, p.2741), they derive the bound for the $L^{\infty}(\mathcal{X})$ norm. Hence, we need to replace the $L^{\infty}(\mathcal{X})$ norm with the $L^{1}(\mathcal{X})$ norm in their proof. ∎
The following useful approximation property is shown by Bos and Schmidt-Hieber, (2022):
Lemma 36 (Theorem 4.1 in Bos and Schmidt-Hieber, (2022)).
Let $0<\varepsilon\leq\frac{1}{2}$, and $\alpha>0$. Then, there are constant $c^{\prime}>0$ and $g_{\bm{W}^{*},\bm{b}^{*}}\in\mathcal{F}_{L_{\textup{log}},J_{\textup{log}},S_{\textup{log}},M_{\textup{log}},\bm{d}_{\textup{log}}}^{\textup{NN}}([0,1])$ with parameters $L_{\textup{log}}\lesssim\log_{2}\varepsilon^{-1}$, $J_{\textup{log}}\leq 1$, $S_{\textup{log}}\lesssim\varepsilon^{-1/\alpha}\log_{2}\varepsilon^{-1}$, $M_{\textup{log}}\leq|\log(4\varepsilon)|\vee\log(1+4\varepsilon)$, and $\bm{d}_{\textup{log}}=(1,\lfloor c^{\prime}\varepsilon^{-1/\alpha}\rfloor,\cdots,\lfloor c^{\prime}\varepsilon^{-1/\alpha}\rfloor,1)$ such that for any $s\in[0,1]$,
$|\exp\circ g_{\bm{W}^{*},\bm{b}^{*}}(s)-s|\leq 4\varepsilon.$ (40)
For the approximation of indicator functions, we adapt the analyses of Petersen and Voigtlaender, (2018) and Imaizumi and Fukumizu, (2019, 2022). A similar approach is considered in (Kim et al., , 2021), though we need to deal with the softmax function. This additional step is done by applying the analyses shown by Bos and Schmidt-Hieber, (2022).
Petersen and Voigtlaender, (2018) show the following fact.
Lemma 37 (Lemma 3.4 in Petersen and Voigtlaender, (2018)).
Let $\alpha>0$, and let $\mathcal{C}^{\alpha}([-2^{-1},2^{-1}]^{d_{0}-1})$ be the unit ball of Hölder space defined on the domain $[-2^{-1},2^{-1}]^{d_{0}-1}$. Let $\mathcal{K}\subset[-2^{-1},2^{-1}]^{d_{0}}$ be any set such that there are function $h\in\mathcal{C}^{\alpha}([-2^{-1},2^{-1}]^{d_{0}-1})$ and index $i\in\{1,\cdots,d_{0}\}$ satisfying
$\mathcal{K}=\{x\in[-2^{-1},2^{-1}]^{d_{0}}\;|\;x_{i}\geq h(x_{\setminus i})\}.$
Then, for every $\varepsilon\in(0,2^{-1})$ and $s>0$, there are constant $c\in\mathbb{N}$ independent of $\varepsilon$ and deep ReLU networks $g_{\bm{W},\bm{b}}\in\mathcal{F}_{L,J,S,M,\bm{d}_{1}}^{\textup{-NN}}([-2^{-1},2^{-1}]^{d_{0}})$ with $L\leq L_{0}\in\mathbb{N}$ ($L_{0}$ is a universal constant), $J\leq\varepsilon^{-c}$, $S\lesssim\varepsilon^{-s(d_{0}-1)/\alpha}$, $M=1$, and $\bm{d}_{1}=(d_{0},d_{\textup{NN},1},\cdots,d_{\textup{NN},L-1},1)$, such that
$\|g_{\bm{W},\bm{b}}-\mathds{1}_{\mathcal{K}}\|_{L^{s}([-2^{-1},2^{-1}]^{d_{0}},\mu)}<\varepsilon.$
The analysis of Imaizumi and Fukumizu, (2019) guarantees the following fact.
Lemma 38 (Consequence of the analysis of Imaizumi and Fukumizu, (2019)).
Let $\alpha>0$. Let $\mathcal{K}\subset\mathcal{X}$ be a subset such that there are functions $h_{1},\cdots,h_{d_{*}}\in\mathcal{C}^{\alpha,d_{0}-1}$, $j_{1},\cdots,j_{d_{*}}\in\{1,\cdots,d_{0}\}$, and $s_{1},\cdots,s_{d_{*}}\in\{1,-1\}$ satisfying
$\mathcal{K}=\bigcap_{i=1}^{d_{*}}\{x\in\mathcal{X}\;|\;s_{i}x_{j_{i}}\geq s_{i}h_{i}(x_{\setminus j_{i}})\}.$
Then, for any $\varepsilon\in(0,2^{-1})$, there are constant $c>0$, universal constant $L_{0}\in\mathbb{N}$, and $g_{\bm{W},\bm{b}}\in\mathcal{F}_{L,J,S,M,\bm{d}_{1}}^{\textup{NN}}$ with $L\leq L_{0}$, $J\leq\varepsilon^{-c}$, $S\lesssim\varepsilon^{-(d_{0}-1)/\alpha}$, $M\lesssim 1$, and $\bm{d}_{1}=(d_{0},d_{\textup{NN},1},\cdots,d_{\textup{NN},L-1},1)$, such that
$\|g_{\bm{W},\bm{b}}-\mathds{1}_{\mathcal{K}}\|_{L^{1}(\mathcal{X})}<\varepsilon.$
Proof.
This fact is shown by Imaizumi and Fukumizu, (2019). Note that Imaizumi and Fukumizu, (2019, Appendix B.1) prove the inequality for the $L^{2}(\mathcal{X})$-norm, using several approximation theorems shown in (Yarotsky, , 2017; Petersen and Voigtlaender, , 2018). In particular, Imaizumi and Fukumizu, (2019, Appendix B.1) show the estimate by applying Lemma 3.4 in (Petersen and Voigtlaender, , 2018) for the $L^{2}(\mathcal{X})$-norm. Hence, the statement above is proven by applying Lemma 3.4 in (Petersen and Voigtlaender, , 2018) with $s=1$ instead of $s=2$ in the proof of Imaizumi and Fukumizu, (2019). ∎
Then, we obtain the following approximation error bounds:
Proposition 39.
Let $0<\varepsilon<2^{-1}$. Let $P\in\mathcal{P}_{\alpha}$, and let $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ be any sequence in $\mathscr{S}_{P}$. Denote by $g_{i}^{*}=\mathds{1}_{\mathcal{K}_{i}}$, $i=1,\cdots,d_{1}$. Then, there are ReLU networks $g_{\bm{W}^{*},\bm{b}^{*}}\in\mathcal{F}_{L^{*},J^{*},S^{*},M^{*},\bm{d}^{*}}^{\textup{NN}}(\mathcal{X})$ with constant $c\geq 0$ and parameters $L^{*}\lesssim\log_{2}\varepsilon^{-1}+1$, $J^{*}\lesssim\varepsilon^{-c}\vee 1$, $S^{*}\lesssim\varepsilon^{-(d_{0}-1)/\alpha}(1\vee\log_{2}\varepsilon^{-1})$, $M^{*}\lesssim|\log(4\varepsilon)|\vee 1$, and $\bm{d}^{*}=(d_{0},d_{\textup{NN},1}^{*},\cdots,d_{\textup{NN},L^{*}-1}^{*},d_{1})$ such that for every $i=1,\cdots,d_{1}$,
$\|(H\circ g_{\bm{W}^{*},\bm{b}^{*}})_{i}-g_{i}^{*}\|_{L^{1}(\mathcal{X})}\leq\varepsilon.$
Proof.
We basically follow the proof of Lemma 4.3 in (Bos and Schmidt-Hieber, , 2022).
For some neural networks $g_{\bm{W}_{1},\bm{b}_{1},i}:\mathbb{R}^{d_{0}}\to\mathbb{R}$ and $g_{\bm{W}_{2},\bm{b}_{2},i}:\mathbb{R}\to\mathbb{R}$ for $i=1,\cdots,d_{1}$, define $g_{\bm{W},\bm{b},i}=g_{\bm{W}_{2},\bm{b}_{2},i}\circ\sigma\circ g_{\bm{W}_{1},\bm{b}_{1},i}$. Following the approach of Lemma 4.3 in (Bos and Schmidt-Hieber, , 2022), we evaluate every quantity $\|\exp\circ g_{\bm{W},\bm{b},i}-g_{i}^{*}\|_{L^{1}(\mathcal{X})}$, $i=1,\cdots,d_{1}$, as follows:
$\left\|\exp\circ g_{\bm{W},\bm{b},i}-g_{i}^{*}\right\|_{L^{1}(\mathcal{X})}$
$\leq\left\|\exp\circ g_{\bm{W},\bm{b},i}-\sigma\circ g_{\bm{W}_{1},\bm{b}_{1},i}\right\|_{L^{1}(\mathcal{X})}+\left\|\sigma\circ g_{\bm{W}_{1},\bm{b}_{1},i}-g_{i}^{*}\right\|_{L^{1}(\mathcal{X})}.$ (41)
Hence, Lemma 35 and (A.7.2) assert that
$\left\|(H\circ g_{\bm{W},\bm{b}})_{i}-g_{i}^{*}\right\|_{L^{1}(\mathcal{X})}$
$\leq 2\sum_{j=1}^{d_{1}}\left\|\exp\circ g_{\bm{W},\bm{b},j}-g_{j}^{*}\right\|_{L^{1}(\mathcal{X})}$ (42)
$\leq 2\sum_{j=1}^{d_{1}}(\left\|\exp\circ g_{\bm{W},\bm{b},j}-\sigma\circ g_{\bm{W}_{1},\bm{b}_{1},j}\right\|_{L^{1}(\mathcal{X})}+\left\|g_{\bm{W}_{1},\bm{b}_{1},j}-g_{j}^{*}\right\|_{L^{1}(\mathcal{X})}),$ (43)
where (A.7.2) follows from Lemma 35. (43) is due to (A.7.2) with the inequality $\|\sigma\circ g_{\bm{W}_{1},\bm{b}_{1},j}-g_{j}^{*}\|_{L^{1}(\mathcal{X})}\leq\|g_{\bm{W}_{1},\bm{b}_{1},j}-g_{j}^{*}\|_{L^{1}(\mathcal{X})}$.
Regarding the second term in (43), by Lemma 38, Lemma 5 in (Imaizumi and Fukumizu, , 2022), and the condition (P2), there are some $c\in\mathbb{N}$ and deep ReLU networks $g_{\bm{W}_{1}^{*},\bm{b}_{1}^{*},j}\in\mathcal{F}_{L_{1},J_{1},S_{1},M_{1},\bm{d}_{1,1}}^{\textup{NN}}$ with $L_{1},J_{1},S_{1},M_{1},\bm{d}_{1,1}$ satisfying the conditions in Lemma 38 such that for every $j=1,\cdots,d_{1}$ we have
$\|g_{\bm{W}_{1}^{*},\bm{b}_{1}^{*},j}-g_{j}^{*}\|_{L^{1}(\mathcal{X})}<\frac{\varepsilon}{10d_{1}}.$ (44)
Regarding the first term in (43), we proceed as follows: By Lemma 36, one can take deep ReLU networks $g_{\bm{W}_{2}^{*},\bm{b}_{2}^{*},1},\cdots,g_{\bm{W}_{2}^{*},\bm{b}_{2}^{*},d_{1}}\in\mathcal{F}_{L_{2},J_{2},S_{2},M_{2},\bm{d}_{1,2}}^{\textup{NN}}$ with $L_{2},J_{2},S_{2},M_{2},\bm{d}_{1,2}$ satisfying the conditions in Lemma 36 so that (40) is satisfied for the error $\frac{\varepsilon}{10d_{1}}$. We need to note that $|g_{\bm{W}_{1}^{*},\bm{b}_{1}^{*},i}|\leq 1$ by the condition $M_{1}=1$, which implies that we can compose the functions $\sigma\circ g_{\bm{W}_{1}^{*},\bm{b}_{1}^{*},i}$ and $g_{\bm{W}_{2}^{*},\bm{b}_{2}^{*},i}$. Thus by Lemma 36, it holds that
$\|\exp\circ g_{\bm{W}_{2}^{*},\bm{b}_{2}^{*},j}\circ\sigma\circ g_{\bm{W}_{1}^{*},\bm{b}_{1}^{*},j}-\sigma\circ g_{\bm{W}_{1}^{*},\bm{b}_{1}^{*},j}\|_{L^{1}(\mathcal{X})}\leq\frac{2\varepsilon}{5d_{1}}.$ (45)
Let $g_{\bm{W}^{*},\bm{b}^{*},j}=g_{\bm{W}_{2}^{*},\bm{b}_{2}^{*},j}\circ\sigma\circ g_{\bm{W}_{1}^{*},\bm{b}_{1}^{*},j}$ for each $j=1,\cdots,d_{1}$. Applying the standard arguments for connecting multiple neural networks (see, e.g., (Schmidt-Hieber, , 2020; Nakada and Imaizumi, , 2020; Bos and Schmidt-Hieber, , 2022; Imaizumi and Fukumizu, , 2019, 2022)), we have that $g_{\bm{W}^{*},\bm{b}^{*}}\in\mathcal{F}_{L^{*},J^{*},S^{*},M^{*},\bm{d}^{*}}^{\textup{NN}}(\mathcal{X})$ with parameters $L^{*}\lesssim\log_{2}\varepsilon^{-1}+1$, $J^{*}\lesssim\varepsilon^{-c}\vee 1$, $S^{*}\lesssim\varepsilon^{-(d_{0}-1)/\alpha}(1\vee(\log_{2}\varepsilon^{-1}))$, $M^{*}\lesssim|\log(4\varepsilon)|\vee 1$, and some integers $\bm{d}^{*}=(d_{0},d_{\textup{NN},1}^{*},\cdots,d_{\textup{NN},L^{*}-1}^{*},d_{1})$, where $d_{\textup{NN},1}^{*},\cdots,d_{\textup{NN},L^{*}-1}^{*}$ are sufficiently large so that the condition in Lemma 36 is satisfied. By (43), (44), and (45), we have $\|(H\circ g_{\bm{W}^{*},\bm{b}^{*}})_{i}-g_{i}^{*}\|_{L^{1}(\mathcal{X})}\leq\varepsilon$. ∎
A.7.3 Step 3
Proposition 40.
Let $\beta\in(0,D_{{\textup{proj}}})$, $\varepsilon>0$, and $P\in\mathcal{P}$. Let $f^{*}=\sum_{i=1}^{d_{1}}g_{i}^{*}v_{i}$ be the contrastive function of $P$. If $f=\sum_{i=1}^{d_{1}}g_{i}v_{i}\in\mathcal{F}\subset\mathcal{F}_{0}$ satisfies
$\|g_{i}-g_{i}^{*}\|_{L^{1}(\mathcal{X})}\leq d_{1}^{-2}\|p_{X}\|_{L^{\infty}(\mathcal{X})}^{-1}\varepsilon\quad\forall i\in\{1,\cdots,d_{1}\},$ (46)
then we have $f\in\mathscr{F}_{\beta,\beta^{-1}\varepsilon,P}(\mathcal{F})$.
Proof.
Let $f\in\mathcal{F}$ be a function satisfying (46). Then, we have
$P_{X}(\{x\in\mathcal{X}\;|\;\|f(x)-f^{*}(x)\|_{2}\geq\beta\})$
$\leq P_{X}\left(\bigcup_{i=1}^{d_{1}}\{x\in\mathcal{X}\;|\;|g_{i}(x)-g_{i}^{*}(x)|\geq d_{1}^{-1}\beta\}\right)$ (47)
$\leq\sum_{i=1}^{d_{1}}d_{1}\beta^{-1}\mathbb{E}_{P_{X}}[|g_{i}-g_{i}^{*}|]$ (48)
$\leq d_{1}\beta^{-1}\|p_{X}\|_{L^{\infty}(\mathcal{X})}\sum_{i=1}^{d_{1}}\|g_{i}-g_{i}^{*}\|_{L^{1}(\mathcal{X})}.$ (49)
Here, in (47) we note that for $x\in\{x^{\prime}\in\mathcal{X}\;|\;\|f(x^{\prime})-f^{*}(x^{\prime})\|_{2}\geq\beta\}$, we have
$\beta\leq\|f(x)-f^{*}(x)\|_{2}\leq\sum_{i=1}^{d_{1}}|g_{i}(x)-g_{i}^{*}(x)|\|v_{i}\|_{2}=\sum_{i=1}^{d_{1}}|g_{i}(x)-g_{i}^{*}(x)|.$
Hence, there is at least one index $i\in\{1,\cdots,d_{1}\}$ such that $|g_{i}(x)-g_{i}^{*}(x)|\geq d_{1}^{-1}\beta$ holds. In (48) and (49), we use Markov’s inequality and Hölder’s inequality, respectively. By (46) and (49), we have
$P_{X}(\{x\in\mathcal{X}\;|\;\|f(x)-f^{*}(x)\|_{2}<\beta\})\geq 1-\beta^{-1}\varepsilon,$
which shows the claim. ∎
A.7.4 Step 4
To apply the results in (Park, , 2009; Kim et al., , 2021), we need to evaluate the covering numbers of the class of classifiers. Given a compact subset $\mathcal{A}$ in a measure space equipped with the Lebesgue measure, let $\mathscr{N}(\delta,\mathcal{H},\|\cdot\|_{L^{\infty}(\mathcal{A})})$ denote the covering number of a class $\mathcal{H}$ of real-valued functions with respect to the pseudo-distance $\|\cdot\|_{L^{\infty}(\mathcal{A})}$ (see e.g., (Steinwart and Christmann, , 2008, Definition 6.19) for covering numbers). Besides, given a map $h:\Delta^{d}\to\mathcal{A}$ for a set $\mathcal{A}$, denote by $h\circ\mathcal{F}=\{h\circ f\;|\;f\in\mathcal{F}\}$. Then, the following evaluation holds.
Lemma 41.
Let $\beta>0$, $\beta_{0}\geq 0$ $P\in\mathcal{P}$, $L\in\mathbb{N}$, $J,S,M\geq 0$, and $\bm{d}=(d_{0},d_{\textup{NN},1},\cdots,d_{\textup{NN},L-1},d_{1})$. Consider a function class $\mathcal{F}^{\prime}\subset\mathcal{F}_{L,J,S,M,\bm{d}}^{\Delta^{d}\textup{-NN}}$. Then, for any $\delta\geq 0$, there exists a constant $C_{6}>0$ independent of $n,\delta$ such that we have
$\log\mathscr{N}(\delta,\psi\circ\rho\circ\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
$\leq d_{1}^{2}\log\mathscr{N}(C_{6}\delta,\mathcal{F}_{L,J,S,M,\bm{d}_{1}}^{\textup{NN}},\|\cdot\|_{L^{\infty}(\mathcal{X})}),$
where $\bm{d}_{1}=(d_{0},d_{\textup{NN},1},\cdots,d_{\textup{NN},L-1},1)$. Note that in the above statement, we can take $C_{6}=2^{-4}dd_{1}^{-1}e^{-4M}D_{\Delta^{d}}^{2}$.
Proof.
Note that $\psi$ is Lipschitz continuous. Thus, we have
$\mathscr{N}(\delta,\psi\circ\rho\circ\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
$\leq\mathscr{N}(2^{-1}D_{\Delta^{d}}^{2}\delta,\rho\circ\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})}).$ (50)
Then, we note that
$\mathscr{N}(2^{-1}D_{\Delta^{d}}^{2}\delta,\rho\circ\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
$=\mathscr{N}(2^{-1}D_{\Delta^{d}}^{2}\delta,\{(x,x^{\prime})\mapsto\|f(x)-f(x^{\prime})\|_{2}^{2}\;|\;f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})\},\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})}).$ (51)
We rearrange the right-hand side of (A.7.4). Write as $f(X_{i})=\sum_{j=1}^{d_{1}}c_{j}(f,X_{i})v_{j}$ and $f(X_{i}^{\prime})=\sum_{j=1}^{d_{1}}c_{j}(f,X_{i}^{\prime})v_{j}$. Similarly to the proof of Proposition 6, by Corollary 2.6 in (Conn et al., , 2009), we have
$\|f(X_{i})-f(X_{i}^{\prime})\|_{2}^{2}=\frac{d_{1}}{d}\sum_{j=1}^{d_{1}}(c_{j}(f,X_{i})-c_{j}(f,X_{i}^{\prime}))^{2}.$
Thus, we obtain
$\mathscr{N}(2^{-1}D_{\Delta^{d}}^{2}\delta,\{(x,x^{\prime})\mapsto\|f(x)-f(x^{\prime})\|_{2}^{2}\;|\;f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})\},\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
$\leq$
$\mathscr{N}(c_{6}\delta,\{(x,x^{\prime})\mapsto\|\{c_{j}(f,x)-c_{j}(f,x^{\prime})\}_{j=1}^{d_{1}}\|_{2}^{2}|f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})\},\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})}),$ (52)
where $c_{6}=2^{-1}dd_{1}^{-1}D_{\Delta^{d}}^{2}$. Here, $\{c_{j}(f,\cdot)-c_{j}(f,\cdot)\}_{j=1}^{d_{1}}$ denotes the sequence of functions $(x,x^{\prime})\mapsto c_{j}(f,x)-c_{j}(f,x^{\prime})$, $j=1,\cdots,d_{1}$, and $\|\cdot\|_{2}$ denotes the 2-norm. Since both $|c_{j}(f,\cdot)|$ and $|c_{j}(f,\cdot)|$ are bounded by the definition of $\Delta^{d}$, by the Lipschitz continuity of the function $s\mapsto s^{2}$ on a closed interval, we have
$\mathscr{N}(c_{6}\delta,\{(x,x^{\prime})\mapsto\|\{c_{j}(f,x)-c_{j}(f,x^{\prime})\}_{j=1}^{d_{1}}\|_{2}^{2}\;|\;f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})\},\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
$\leq\prod_{j=1}^{d_{1}}\mathscr{N}(c_{6}^{\prime}\delta,\{(x,x^{\prime})\mapsto c_{j}(f,x)-c_{j}(f,x^{\prime})\;|\;f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})\},\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
$\leq\prod_{j=1}^{d_{1}}\mathscr{N}(2^{-3}dd_{1}^{-2}D_{\Delta^{d}}^{2}\delta,\{x\mapsto c_{j}(f,x)\;|\;f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})\},\|\cdot\|_{L^{\infty}(\mathcal{X})}),$ (53)
where $c_{6}^{\prime}=2^{-1}d_{1}^{-1}c_{6}$. Combining (A.7.4), (A.7.4), (A.7.4), and (A.7.4), we have
$\mathscr{N}(\delta,\psi\circ\rho\circ\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
$\leq\prod_{j=1}^{d_{1}}\mathscr{N}(L_{2}\delta,\{x\mapsto c_{j}(f,x)\;|\;f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})\},\|\cdot\|_{L^{\infty}(\mathcal{X})}),$ (54)
where $L_{2}=2^{-3}dd_{1}^{-2}D_{\Delta^{d}}^{2}$. Here for every $j=1,\cdots,d_{1}$, $c_{j}(f,\cdot)$ is represented as a softmax function by the definition of $\mathcal{F}_{L,J,S,M,\bm{d}}^{\Delta\textup{-NN}}$. Write $c_{j}(f,x)$ as a function of $g_{1}(x),\cdots,g_{d_{1}}(x)$, namely $c_{j}(f,x)=\exp(g_{j}(x))/\sum_{i=1}^{d_{1}}\exp(g_{i}(x))$, where $g_{1},\cdots,g_{d_{1}}\in\mathcal{F}_{L,J,S,M,\bm{d}_{1}}^{\textup{NN}}$. For the softmax function on the domain $[-M,M]^{d_{1}}$ and any $x_{1},\cdots,x_{d_{1}},x_{1}^{\prime},\cdots,x_{d_{1}}^{\prime}\in[-M,M]$, we have
$\left|\frac{e^{x_{j}}}{\sum_{i=1}^{d_{1}}e^{x_{i}}}-\frac{e^{x_{j}^{\prime}}}{\sum_{i=1}^{d_{1}}e^{x_{i}^{\prime}}}\right|$ $\leq d_{1}^{-2}e^{2M}\sum_{i=1}^{d_{1}}|e^{x_{j}+x_{i}^{\prime}}-e^{x_{j}^{\prime}+x_{i}}|$
$\leq 2d_{1}^{-2}e^{4M}\sum_{j=1}^{d_{1}}|x_{j}-x_{j}^{\prime}|.$
Hence, we have
$\log\mathscr{N}(L_{2}\delta,\{x\mapsto c_{j}(f,x)\;|\;f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})\},\|\cdot\|_{L^{\infty}(\mathcal{X})})$
$\leq d_{1}\log\mathscr{N}(2^{-1}d_{1}e^{-4M}L_{2}\delta,\mathcal{F}_{L,J,S,M,\bm{d}_{1}}^{\textup{NN}},\|\cdot\|_{L^{\infty}(\mathcal{X})}).$ (55)
Thus, setting $C_{6}=2^{-1}d_{1}e^{-4M}L_{2}$, by (A.7.4) and (A.7.4), we obtain
$\log\mathscr{N}(\delta,\psi\circ\rho\circ\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
$\leq d_{1}^{2}\log\mathscr{N}(C_{6}\delta,\mathcal{F}_{L,J,S,M,\bm{d}_{1}}^{\textup{NN}},\|\cdot\|_{L^{\infty}(\mathcal{X})}),$ (56)
which shows the assertion. ∎
The following fact is shown by Nakada and Imaizumi, (2020).
Lemma 42 (Lemma 21 in Nakada and Imaizumi, (2020)).
Let $L\in\mathbb{N}$, $J,S,M\geq 0$, and $\bm{d}_{1}=(d_{0},d_{\textup{NN},1},\cdots,d_{\textup{NN},L-1},1)$. Then, for every $s>0$, it holds that
$\log\mathscr{N}(s,\bigcup_{L^{\prime}\leq L}\mathcal{F}_{L^{\prime},J,S,M,\bm{d}_{1}}^{\textup{NN}},\|\cdot\|_{L^{\infty}(\mathcal{X})})\leq S\log(2s^{-1}LJ^{L}(S+1)^{L}).$ (57)
Given a compact subset $\mathcal{A}$ of a measure space equipped with the Lebesgue measure, the bracketing number of a class $\mathcal{H}$ of real-valued functions on $\mathcal{A}$ with respect to $\|\cdot\|_{L^{s}(\mathcal{A})}$ is denoted by
$\mathscr{B}(\delta,\mathcal{H},\|\cdot\|_{L^{s}(\mathcal{A})})$
(see, e.g., (van de Geer, , 2000, Definition 2.2) for bracketing numbers). Now, we recall a useful fact on the excess risk, which is first shown by Park, (2009). In what follows, we recall a simplified version shown by Kim et al., (2021):
Lemma 43 (Theorem A.1 in Kim et al., (2021)).
Let $\mathcal{X}_{0}=[0,1]^{d_{3}}$ for some $d_{3}\in\mathbb{N}$. For every $n\in\mathbb{N}$, consider a constant $M_{n}>0$ and a class $\mathcal{H}_{n}$ of $M_{n}$-uniformly bounded, measurable real-valued functions on $\mathcal{X}_{0}$ (namely, $\sup_{h\in\mathcal{H}_{n}}\|h\|_{\infty}\leq M_{n}$). Let ¯:×$\bar{\ell}:\mathbb{R}\times\mathcal{Y}\to\mathbb{R}$ be a loss function such that there is a $c_{0}$-Lipschitz continuous function $\tilde{\ell}:\mathbb{R}\to\mathbb{R}$ satisfying ¯$\bar{\ell}(z,y)=\tilde{\ell}(yz)$ for any $y\in\mathcal{Y}$ and any $z\in\mathbb{R}$. Let $P_{0}$ be a Borel probability measure in $\mathcal{X}_{0}\times\mathcal{Y}$. Suppose that there is a measurable function $h^{*}:\mathcal{X}_{0}\to\mathbb{R}$ satisfying $\mathbb{E}_{P_{0}}[\bar{\ell}(h^{*}(x_{0}),y)]=\inf_{h}\mathbb{E}_{P_{0}}[\bar{\ell}(h(x_{0}),y)]$ (where the infimum is taken over all the measurable functions on $\mathcal{X}$), and there are some $\kappa_{0}\in(0,1]$, positive constants $c$, $c^{\prime}=c^{\prime}(\kappa_{0},c_{0},c)$, and positive sequences $\{\varepsilon_{n}\}_{n\in\mathbb{N}}$ and $\{\widetilde{\varepsilon}_{n}\}_{n\in\mathbb{N}}$, such that the conditions below are satisfied for an arbitrary $n\in\mathbb{N}$:
  1. (C1)
    There is a function $h\in\mathcal{H}_{n}$ such that $\mathbb{E}_{P_{0}}[\bar{\ell}(h(x_{0}),y)]-\mathbb{E}_{P_{0}}[\bar{\ell}(h^{*}(x_{0}),y)]\leq\varepsilon_{n}$.
  2. (C2)
    The following inequality is satisfied for every $h\in\mathcal{H}_{n}$:
    $\mathbb{E}_{P_{0}}[|\bar{\ell}(h(x_{0}),y)-\bar{\ell}(h^{*}(x_{0}),y)|^{2}]$
    $\leq cM_{n}^{2-\kappa_{0}}(\mathbb{E}_{P_{0}}[\bar{\ell}(h(x_{0}),y)-\bar{\ell}(h^{*}(x_{0}),y)])^{\kappa_{0}}.$
  3. (C3)
    It holds that $\log\mathscr{B}(\widetilde{\varepsilon}_{n},\mathcal{H}_{n},\|\cdot\|_{L^{2}(\mathcal{X}_{0})})\leq c^{\prime}n\widetilde{\varepsilon}_{n}^{2-\kappa_{0}}M_{n}^{-2+\kappa_{0}}$.
Then, for the estimator $\widehat{h}_{n}:(\mathcal{X}_{0}\times\mathcal{Y})^{n}\to\mathcal{H}_{n}$ such that for any sequence of pairs $(x_{0,1},y_{1}),\cdots,(x_{0,n},y_{n})\in\mathcal{X}_{0}\times\mathcal{Y}$ the empirical risk $\frac{1}{n}\sum_{i=1}^{n}\bar{\ell}(h(x_{0,i}),y_{i})$ is minimized at $\widehat{h}_{n}((x_{0,1},y_{1}),\cdots,(x_{0,n},y_{n}))$ in $\mathcal{H}_{n}$, there are positive universal constants $c_{1},c_{2}>0$ such that the following inequality holds:
$\mathbb{E}[\mathds{1}_{\{\mathbb{E}_{P_{0}}[\bar{\ell}(\widehat{h}_{n}(x_{0}),y)-\bar{\ell}(h^{*}(x_{0}),y)]\geq(2\varepsilon_{n})\vee 128c_{0}^{-1}\widetilde{\varepsilon}_{n}\}}]$
$\leq c_{1}\exp(-c_{2}n((2\varepsilon_{n})\vee(128c_{0}^{-1}\widetilde{\varepsilon}_{n}))^{2-\kappa_{0}}M_{n}^{-2+\kappa_{0}}).$
We aim at applying this fact to a pairwise binary classification setting by checking the conditions (C1)–(C3) in Lemma 43.
We now prove Theorem 12.
Proof of Theorem 12.
Let $P\in\mathcal{P}_{\alpha}$. First, note that since $p_{X}$ is continuous on $\mathcal{X}=[0,1]^{d_{0}}$ by the condition (A3) in Definition 1, we have $d_{1}^{-2}\|p_{X}\|_{L^{\infty}(\mathcal{X})}^{-1}\leq 1$. Also, $\varepsilon_{n}<\frac{1}{2}$ by the definition. By Proposition 39, there exist some $c>0$ and some $L^{*}\lesssim\log_{2}\varepsilon_{n}^{-1}+1$, $1\leq J^{*}\lesssim\varepsilon_{n}^{-c}\vee 1$, $S^{*}\lesssim\varepsilon_{n}^{-(d_{0}-1)/\alpha}(1\vee\log_{2}\varepsilon_{n}^{-1})$, $M^{*}\lesssim|\log(4d_{1}^{-2}\|p_{X}\|_{L^{\infty}(\mathcal{X})}^{-1}\varepsilon_{n})|\vee 1$, and $\bm{d}^{*}=(d_{0},d_{\textup{NN},1}^{*},\cdots,d_{\textup{NN},L^{*}-1}^{*},d_{1})$ such that
$\|(H\circ g_{\bm{W}^{*},\bm{b}^{*}})_{i}-g_{i}^{*}\|_{L^{1}(\mathcal{X})}\leq d_{1}^{-2}\|p_{X}\|_{L^{\infty}(\mathcal{X})}^{-1}\varepsilon_{n}.$ (58)
Note that the subclass $\mathcal{F}^{*}=\mathcal{F}_{L^{*},J^{*},S^{*},M^{*},\bm{d}^{*}}^{\Delta^{d}\textup{-NN}}$ is considered in this proof. Denote by $\mathcal{F}_{\textup{local}}^{*}=\mathscr{F}_{\beta,\beta^{-1}\varepsilon_{n},P}(\mathcal{F}^{*})$. Then, the set $\psi\circ\rho\circ\mathcal{F}_{\textup{local}}^{*}$ is $1$-uniformly bounded. Also, it is well known that hinge loss is $1$-Lipschitz continuous in the first argument (see, e.g., (Steinwart and Christmann, , 2008, Example 2.27)). Thus, we can set $c_{0}=1$ and $M_{n}=1$ hereafter.
We consider to apply Lemma 43 to Theorem 17. To this end, we shall check that the conditions (C1) – (C3) in Lemma 43 are satisfied.
Checking the condition (C1) in Lemma 43.
By (58) and Proposition 40, for the function $f_{\bm{W}^{*},\bm{b}^{*}}=\sum_{i=1}^{d_{1}}H_{i}\circ g_{\bm{W}^{*},\bm{b}^{*}}v_{i}$, we have
$f_{\bm{W}^{*},\bm{b}^{*}}\in\mathcal{F}_{\textup{local}}^{*}.$ (59)
Let $\varepsilon_{\mathcal{F}_{\textup{local}}^{*},f^{*}}$ be the quantity defined in (36). By Proposition 34, there is a constant $C_{1}>0$ independent of $n$ such that
[…]
Here, we note that by (58) and (59), we have
$\varepsilon_{\mathcal{F}_{\textup{local}}^{*},f^{*}}\leq d_{1}^{-2}\|p_{X}\|_{L^{\infty}(\mathcal{X})}^{-1}\varepsilon_{n}.$ (61)
Hence, applying Lemma 32 and (61) to (60), we have that
[…]
By (62) and Proposition 30, there is some $f_{\bm{W},\bm{b}}\in\mathcal{F}_{\textup{local}}^{*}$ such that we have
$\mathbb{E}_{P}[\ell_{f_{\bm{W},\bm{b}}}]-\mathbb{E}_{P}[\ell_{f^{*}}]\leq C_{1}d_{1}^{-2}\|p_{X}\|_{L^{\infty}(\mathcal{X})}^{-1}\varepsilon_{n}.$
Since $\psi\circ\rho_{f^{*}}$ is the Bayes classifier by Lemma 26, the condition (C1) in Lemma 43 is satisfied.
Checking the condition (C2) in Lemma 43.
By Proposition 1 in (Lecué, , 2007) (see Lemma 13; see also Lemma 6.1 in (Steinwart and Scovel, , 2007)), the condition (C2) in Lemma 43 is satisfied with $\kappa_{0}=1$.
Checking the condition (C3) in Lemma 43.
By (van de Geer, , 2000, Lemma 2.1) and (Kim et al., , 2021, (A.1)), for any $s>0$, we have
$\log\mathscr{B}(s,\psi\circ\rho\circ\mathcal{F}_{\textup{local}}^{*},\|\cdot\|_{L^{2}(\mathcal{X}^{2})})\leq\log\mathscr{N}(\frac{s}{2},\psi\circ\rho\circ\mathcal{F}_{\textup{local}}^{*},\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})}).$ (63)
By Lemma 41 and Lemma 42, for the constant $C_{6}=2^{-4}dd_{1}^{-1}e^{-4M^{*}}D_{\Delta^{d}}^{2}$, we have
$\log\mathscr{N}(2^{-1}s,\psi\circ\rho\circ\mathcal{F}_{\textup{local}}^{*},\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
$\leq d_{1}^{2}\log\mathscr{N}(2^{-1}C_{6}s,\mathcal{F}_{L^{*},J^{*},S^{*},M^{*},\bm{d}_{1}^{*}}^{\textup{NN}},\|\cdot\|_{L^{\infty}(\mathcal{X})})$
$\leq d_{1}^{2}S^{*}\log(4s^{-1}C_{6}^{-1}L^{*}(J^{*})^{L^{*}}(S^{*}+1)^{L^{*}}),$ (64)
where Lemma 41 is applied in the first inequality, and Lemma 42 is used in the second inequality. Here, $\bm{d}_{1}^{*}=(d_{0},d_{\textup{NN},1}^{*},\cdots,d_{\textup{NN},L^{*}-1}^{*},1)$. Hence, by (63) and (64), the condition (C3) in Lemma 43 is satisfied with any $\widetilde{\varepsilon}_{0,n}$ satisfying
$d_{1}^{2}S^{*}\log(4\widetilde{\varepsilon}_{0,n}^{-1}C_{6}^{-1}L^{*}(J^{*})^{L^{*}}(S^{*}+1)^{L^{*}})\leq c^{\prime}n\widetilde{\varepsilon}_{0,n},$ (65)
where $c^{\prime}$ is some constant independent of $n$. Hence, we define $\widetilde{\varepsilon}_{n}$ as
$\widetilde{\varepsilon}_{n}$ $=n^{-1}\varepsilon_{n}^{-\frac{d_{0}-1}{\alpha}}(1\vee\log_{2}\varepsilon_{n}^{-1})\log^{2}{n}$
$=n^{-\frac{\alpha}{\alpha+d_{0}-1}}(1\vee\log_{2}\varepsilon_{n}^{-1})\log^{2}{n}$
$=\varepsilon_{n}(1\vee\log_{2}\varepsilon_{n}^{-1})\log^{2}{n},$
where in the second equality we used the definition $\varepsilon_{n}=n^{-\frac{\alpha}{\alpha+d_{0}-1}}$. Then, $\widetilde{\varepsilon}_{n}$ satisfies (65).
Therefore, all the conditions (C1) – (C3) in Lemma 43 are satisfied.
Concluding the proof.
By Theorem 17, there are positive constants $C,C^{\prime}$, and $C^{\prime\prime}$ independent of $n$ such that
$\mathcal{R}(\widehat{g}_{n}^{\textup{LERM}};P,U_{1}^{n})\leq C(\log{n})\mathbb{E}[\mathcal{E}(\widehat{f}_{n}^{\textup{LERM}}(U_{1}^{n}))]+\frac{C^{\prime}\varepsilon_{n}}{\beta}+\frac{C^{\prime\prime}}{n}.$ (66)
By Lemma 43, there are positive universal constants $c_{1},c_{2}$ such that we have
$\mathbb{E}[\mathds{1}_{\{\mathcal{E}(\widehat{f}_{n}^{\textup{LERM}}(U_{1}^{n}))\geq 128C_{1}\widetilde{\varepsilon}_{n}\}}]\leq c_{1}e^{-128c_{2}n\widetilde{\varepsilon}_{n}},$ (67)
where $\varepsilon_{n}\leq\widetilde{\varepsilon}_{n}$ since $n>2$ by the definition. Then, we have
[…]
$\leq 128C_{1}\widetilde{\varepsilon}_{n}+2c_{1}e^{-128c_{2}n\widetilde{\varepsilon}_{n}},$ (69)
where in (69), we used (67). By the definition of $\widetilde{\varepsilon}_{n}$, there is a natural number $N$ such that for every $n\geq N$, we have $128C_{1}\widetilde{\varepsilon}_{n}\geq 2c_{1}e^{-128c_{2}n\widetilde{\varepsilon}_{n}}$. Thus, by (66) and (69), for any $n\geq N$, we have
$\mathcal{R}(\widehat{g}_{n}^{\textup{LERM}};P,U_{1}^{n})\leq 256C_{1}C\varepsilon_{n}(1\vee\log_{2}\varepsilon_{n}^{-1})(\log{n})^{3}+\frac{C^{\prime}}{\beta}\varepsilon_{n}+\frac{C^{\prime\prime}}{n}.$
Define
$C^{*}=3\max\{256(\log{2})^{-1}C_{1}C\alpha(\alpha+d_{0}-1)^{-1},C^{\prime}\beta^{-1},C^{\prime\prime}\}.$
Since we have selected as $\varepsilon_{n}=n^{-\alpha/(\alpha+d_{0}-1)}$ such that $\varepsilon_{n}<\frac{1}{2}$, if $n\geq N$, we obtain
$\mathcal{R}(\widehat{g}_{n}^{\textup{LERM}};P,U_{1}^{n})\leq C^{*}n^{-\frac{\alpha}{\alpha+d_{0}-1}}\log^{4}{n}.$ (70)
Therefore, we have shown the claim. ∎
A.8 Proof of Corollary 19
Proof.
The proof is almost the same as that of Theorem 12. Specifically, by Proposition 6 and (25) in the proof of Theorem 17, there are positive constants $C,C^{\prime},C^{\prime\prime}$ independent of $n$ such that in the event $\{\omega\in\Omega\;|\;\widehat{f}_{n}^{\textup{GERM}}(U_{1}^{n}(\omega))\in\mathscr{F}_{\beta,\beta^{-1}\varepsilon_{n},P}(\mathcal{F}^{*})\}$ we have
^$\widehat{\mathcal{R}}(\widehat{g}_{n}^{\textup{GERM}};P,U_{1}^{n}(\omega))$
$\leq C(\log{n})\mathcal{E}(\widehat{f}_{n}^{\textup{GERM}}(U_{1}^{n}(\omega)))+\frac{C^{\prime}\varepsilon_{n}}{\beta}+\frac{C^{\prime\prime}}{n}.$ (71)
Also, note that the fact shown in (Park, , 2009; Kim et al., , 2021) (see also Lemma 43) are applicable to the global ERM since we can check that the conditions (C1) – (C3) are satisfied, in the same way as the proof of Theorem 12. Hence, we apply the same arguments as the proof of Theorem 12: By Lemma 43, there are positive universal constants $c_{1},c_{2}$ such that with probability at least $1-c_{1}\exp(-128c_{2}n\varepsilon_{n}(1\vee\log_{2}\varepsilon_{n}^{-1})(\log^{2}{n}))$, we have
$\mathcal{E}(\widehat{f}_{n}^{\textup{GERM}}(U_{1}^{n}))\leq 128C_{1}\varepsilon_{n}(1\vee\log_{2}\varepsilon_{n}^{-1})(\log^{2}{n}),$
where $C_{1}$ is the constant introduced in Proposition 34.
Therefore, we obtain the following: With probability at least
$1-c_{1}\exp(-128c_{2}n\varepsilon_{n}(\log_{2}\varepsilon_{n}^{-1})(\log^{2}{n}))-Q(\widehat{f}_{n}^{\textup{GERM}}(U_{1}^{n})\notin\mathscr{F}_{\beta,\beta^{-1}\varepsilon_{n},P}(\mathcal{F}^{*})),$
we have
^$\widehat{\mathcal{R}}(\widehat{g}_{n}^{\textup{GERM}};P,U_{1}^{n})$ $\leq C^{*}(\varepsilon_{n}(\log\varepsilon_{n}^{-1})(\log^{3}{n})+\varepsilon_{n}+\frac{1}{n})$
$\leq 3C^{*}\varepsilon_{n}(\log\varepsilon_{n}^{-1})(\log^{3}{n}),$ (72)
where
$C^{*}=\max\{128(\log{2})^{-1}C_{1}C,C^{\prime}\beta^{-1},C^{\prime\prime}\},$
and we used the conditions $\varepsilon_{n}\geq n^{-\frac{\alpha}{\alpha+d_{0}-1}}$ and $\varepsilon_{n}<2^{-1}$. This inequality shows the claim. ∎
A.9 Proof of Lemma 22
Proof.
Given any Borel probability measure $P_{X,Z_{d_{1}}}$ satisfying $1$-(MNC), let $z^{*}$ be any Bayes classifier of $P_{X,Z_{d_{1}}}$. Recall that the condition 1-(MNC) introduced in (Kim et al., , 2021) assumes that there is a constant $t_{0}\in(0,1]$ such that we have
$\min_{i\in\mathcal{Z}_{d_{1}}\setminus\{z^{*}(x)\}}|p(z=i|x)-p(z=z^{*}(x)|x)|>t_{0}\quad\textup{for }P_{X}\textup{-almost all }x\in\mathcal{X}.$
Hence, for $P_{X}$-almost all $x\in\mathcal{X}$, for every $i\in\mathcal{Z}_{d_{1}}\setminus\{z^{*}(x)\}$ we have
$p(z=z^{*}(x)|x)>p(z=i|x)+t_{0}.$
Thus, we have
$\max_{i\in\mathcal{Z}_{d_{1}}\setminus\{z^{*}(x)\}}p(z=i|x)<p(z=z^{*}(x)|x)\quad\textup{for }P_{X}\textup{-almost all }x\in\mathcal{X}.$ (73)
This inequality implies that the cardinality of $\operatorname*{arg\;max}_{i=1,\cdots,d_{1}}p(z=i|x)$ is one for $P_{X}$-almost all $x\in\mathcal{X}$. Therefore, for any Bayes classifiers $z_{1}^{*}$ and $z_{2}^{*}$ of $P_{X,Z_{d_{1}}}$, we have
$\sum_{i=1}^{d_{1}}D_{P_{X}}((z_{1}^{*})^{-1}(i),(z_{2}^{*})^{-1}(i))=0.$ (74)
We now construct a probability distribution. Let $\mathscr{S}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}\in\mathscr{P}_{\alpha}$. Let $P_{X,Z_{d_{1}}}^{(0)}$ be a Borel probability measure in $\mathcal{X}\times\mathcal{Z}_{d_{1}}$ such that it has density $p(x,z)$ in $\mathcal{X}\times\mathcal{Z}_{d_{1}}$ and satisfies
$p(z=i|x)=\begin{cases}1&\quad\textup{if }x\in\mathcal{K}_{i},\\ 0&\quad\textup{otherwise}.\end{cases}$
Then, the function $z_{0}^{*}:\mathcal{X}\to\mathcal{Z}_{d_{1}}$ satisfying $(z_{0}^{*})^{-1}(i)=\mathcal{K}_{i}$ for every $i\in\mathcal{Z}_{d_{1}}$ is a Bayes classifier of $P_{X,Z_{d_{1}}}^{(0)}$. Note that $P_{X,Z_{d_{1}}}^{(0)}$ satisfies $1$-(MNC) by the definition. Thus, by (74), for any Bayes classifier $z^{*}$ of $P_{X,Z_{d_{1}}}$ we have
$\sum_{i=1}^{d_{1}}D_{P_{X}}((z^{*})^{-1}(i),\mathcal{K}_{i})=0.$
We obtain the claim. ∎
A.10 Proof of Lemma 23
Proof.
We introduce the subset
$\mathcal{J}_{i}=\{x\in\mathcal{X}\;|\;p(z=i|x)>\max_{j\in\mathcal{Z}_{d_{1}}\setminus\{i\}}p(z=j|x)\}.$
Let $z^{*}$ be any Bayes classifier of $P_{X,Z_{d_{1}}}$. Let $\mathcal{J}=\bigcup_{i=1}^{d_{1}}\mathcal{J}_{i}\times\{i\}$. Also, define
$\mathcal{J}^{\prime}=\{x\in\mathcal{X}\;|\;p(z=z^{*}(x)|x)\leq\max_{j\in\mathcal{Z}_{d_{1}}\setminus\{z^{*}(x)\}}p(z=j|x)\}.$
By (73), $P_{X}(\mathcal{J}^{\prime})=0$. Note that $\mathcal{X}=(\bigcup_{i=1}^{d_{1}}\mathcal{J}_{i})\cup\mathcal{J}^{\prime}$. Hence, we have
$P_{X,Z_{d_{1}}}((\mathcal{X}\times\mathcal{Z}_{d_{1}})\setminus\mathcal{J})$
$=P_{X,Z_{d_{1}}}\left(\bigcup_{i=1}^{d_{1}}\mathcal{J}^{\prime}\times\{i\}\right)+P_{X,Z_{d_{1}}}\left(\bigcup_{i\neq j}\mathcal{J}_{j}\times\{i\}\right)$
$\leq P_{X}(\mathcal{J}^{\prime})+P_{X,Z_{d_{1}}}(z^{*}(x)\neq z)$
$\leq P_{X,Z_{d_{1}}}(z^{*}(x)\neq z).$
Here, we notice that
$\mathcal{J}\subset\bigcup_{i=1}^{d_{1}}(z^{*})^{-1}(i)\times\{i\}.$
From now on, we fix $\omega\in\Omega$ and abbreviate as $\widehat{z}_{n}=\widehat{z}_{n}(U_{1}^{n}(\omega))$. Then, we have
$P_{X,Z_{d_{1}}}(\widehat{z}_{n}(x)\neq z)$
$\leq\sum_{i=1}^{d_{1}}\int_{(z^{*})^{-1}(i)}\mathds{1}_{\{\widehat{z}_{n}(x)\neq i\}}(x)p(z=i|x)p(x)\mu(dx)+P_{X,Z_{d_{1}}}(z^{*}(x)\neq z).$
Let $\varepsilon\geq 0$ be arbitrary. Let $z_{0}^{*}$ be the Bayes classifier such that we have $P_{X,Z_{d_{1}}}(z_{0}^{*}(x)\neq z)-\inf_{z^{*}}P_{X,Z_{d_{1}}}(z^{*}(x)\neq z)\leq\varepsilon$. Then, we obtain
$P_{X,Z_{d_{1}}}(\widehat{z}_{n}(x)\neq z)-\inf_{z^{*}}P_{X,Z_{d_{1}}}(z^{*}(x)\neq z)$
$\leq\sum_{i=1}^{d_{1}}\int_{(z_{0}^{*})^{-1}(i)}\mathds{1}_{\{\widehat{z}_{n}(x)\neq i\}}(x)P_{X}(dx)+\varepsilon,$ (75)
where in (75) we further use the inequality $p(z=i|x)\leq 1$.
Note that $D_{P_{X}}((z_{0}^{*})^{-1}(i),\mathcal{K}_{i})=0$ by the condition (B2). Also, we have
$(z_{0}^{*})^{-1}(i)\subset\mathcal{K}_{i}\cup((z_{0}^{*})^{-1}(i)\setminus\mathcal{K}_{i})\subset\mathcal{K}_{i}\cup(((z_{0}^{*})^{-1}(i)\cup\mathcal{K}_{i})\setminus((z_{0}^{*})^{-1}(i)\cap\mathcal{K}_{i}))$
Hence, we have
$\int_{(z_{0}^{*})^{-1}(i)}\mathds{1}_{\{\widehat{z}_{n}(x)\neq i\}}(x)P_{X}(dx)$ $\leq\int_{\mathcal{K}_{i}}\mathds{1}_{\{\widehat{z}_{n}(x)\neq i\}}(x)P_{X}(dx)$
$=\int_{\mathcal{K}_{i}}(1-\mathds{1}_{\{\widehat{z}_{n}(x)=i\}}(x))P_{X}(dx).$ (76)
We also abbreviate as $\widehat{f}_{n}=\widehat{f}_{n}(U_{1}^{n}(\omega))$. Here note that
$\mathcal{D}_{i}:=\left\{x\in\mathcal{X}\;|\;\|\widehat{f}_{n}(x)-v_{i}\|_{2}<D_{\Delta^{d}}/2\right\}\subseteq\{x\in\mathcal{X}\;|\;\widehat{z}_{n}(x)=i\}.$ (77)
Let
$\mathcal{D}=\left\{x\in\mathcal{X}\;|\;\|\widehat{f}_{n}(x)-f^{*}(x)\|_{2}<D_{\Delta^{d}}/2\right\}.$
Note that $\mathcal{D}_{i}\cap\mathcal{K}_{i}=\mathcal{D}\cap\mathcal{K}_{i}$ for any $i=1,\cdots,d_{1}$. Then, by (75) and (76) we have
$P_{X,Z_{d_{1}}}(\widehat{z}_{n}(x)\neq z)-\inf_{z^{*}}P_{X,Z_{d_{1}}}(z^{*}(x)\neq z)$
$\leq\sum_{i=1}^{d_{1}}\int_{\mathcal{K}_{i}}(1-\mathds{1}_{\mathcal{D}_{i}}(x))P_{X}(dx)+\varepsilon$ (78)
$=1-\sum_{i=1}^{d_{1}}\int_{\mathcal{K}_{i}}\mathds{1}_{\mathcal{D}}(x)P_{X}(dx)+\varepsilon$ (79)
$=1-P_{X}(\mathcal{D})+\varepsilon$ (80)
$\leq 4D_{\Delta^{d}}^{-2}\mathbb{E}_{P_{X}}[\|\widehat{f}_{n}(x)-f^{*}(x)\|_{2}^{2}]+\varepsilon,$ (81)
where in (A.10) we use (75) and (76), in (79) we note that $f^{*}(x)=v_{i}$ for any $x\in\mathcal{K}_{i}$ by Definition 5, and in (81) we apply Markov’s inequality.
Finally, since $\varepsilon\geq 0$ is arbitrary in (81), by applying Proposition 6 and integrating with respect to $\omega$, we obtain the claim. ∎
Appendix B Notation Lists
We summarize the notations in Table 2 – Table 6.
Table 2: Some general notations.
Notation(s) Definition(s) Section
$\mathcal{B}(\mathcal{A})$ Borel $\sigma$-algebra of a topological space $\mathcal{A}$ Section 2.1
$\|\cdot\|_{L^{s}(\mathcal{A},\nu)}$ $L^{s}$-norm Section 2.1
$\mathcal{L}^{s}(\mathcal{A},\nu)$ $\mathcal{L}^{s}$-space Section 2.1
$\|\cdot\|_{\mathcal{A},\nu,s}$ $\|f\|_{\mathcal{A},\nu,s}=\|\|f\|_{s}\|_{L^{s}(\mathcal{A},\nu)}$, $f$: vector-valued function Section 2.1
sign $\textup{sign}(s)=1$ if $s\geq 0$ and $-1$ if $s<0$ Section 2.1
Table 3: Notations of sets and related notions.
Notation(s) Definition(s) Section
$\mathcal{X},\mu$ $\mathcal{X}=[0,1]^{d_{0}}$, $\mu$: Lebesgue measure Section 1
$\mathcal{L}^{s}(\mathcal{X})$ $\mathcal{L}^{s}(\mathcal{X})=\mathcal{L}^{s}(\mathcal{X},\mu)$ Section 2.1
$\|\cdot\|_{\mathcal{X},s}$ $\|f\|_{\mathcal{X},s}=\|f\|_{\mathcal{X},\mu,s}$ Section 2.1
$\mathcal{Y}$ $\mathcal{Y}=\{1,-1\}$ Section 1
$\mathcal{Z}_{d_{1}}$ $\mathcal{Z}_{d_{1}}=\{1,\cdots,d_{1}\}$ Section 5
$\mathcal{S}^{d-1}$ the unit hypersphere in $\mathbb{R}^{d}$ Section 3.1
$\Delta^{d}$ a regular simplex in $\mathbb{R}^{d}$ Section 3.1
$v_{1},\cdots,v_{d_{1}}$ the vertices of $\Delta^{d}$ Section 3.1
$D_{\Delta^{d}}$ the diameter of the regular simplex $\Delta^{d}$ Section 3.1
$(\Omega,\Sigma,Q)$ a probability space Section 1
$\mathcal{P}_{\alpha},\mathcal{P}$ classes of Borel probability measures in $\mathcal{X}^{2}\times\mathcal{Y}$ Section 2.2
$\widetilde{\mathscr{P}}_{\alpha},\mathscr{P}_{\alpha}$ classes of smooth partitions introduced in (Imaizumi and Fukumizu, , 2022) Section 2.2
Table 4: Notations of probability distributions.
Notation(s) Definition(s) Section
$p(x,x^{\prime},y)$ p.d.f. satisfying (2) introduced in (Tsai et al., , 2020) Section 2.1
$p_{X}(x)$ $p_{X}(x)=\int_{\mathcal{X}}(p(x,x^{\prime},1)+p(x,x^{\prime},-1))\mu(dx^{\prime})$ Section 1.1
$p_{X^{\prime}}(x^{\prime})$ $p_{X^{\prime}}(x^{\prime})=\int_{\mathcal{X}}(p(x,x^{\prime},1)+p(x,x^{\prime},-1))\mu(dx)$ Section 1.1
$p_{X,X^{\prime}}(x,x^{\prime})$ $p_{X,X^{\prime}}(x,x^{\prime})=p(x,x^{\prime},1)+p(x,x^{\prime},-1)$ Section 2.1
$p_{Y}(y)$ $p_{Y}(y)=\int_{\mathcal{X}\times\mathcal{X}}p(x,x^{\prime},y)\mu(dx)\mu(dx^{\prime})$ Section 1.1
$q(x,x^{\prime})$ $q(x,x^{\prime})=p(x,x^{\prime}|y=1)$, following (Tsai et al., , 2020) Section 1.1
$\eta(x,x^{\prime})$ $\eta(x,x^{\prime})=p(y=1|x,x^{\prime})$ Section 2.1
$P$ probability measure whose density is $p(x,x^{\prime},y)$ Section 1
$P_{X,X^{\prime}}$ probability measure whose density is $p_{X,X^{\prime}}$ Section 2.1
$P_{X},P_{X^{\prime}}$ probability measures with densities $p_{X}$ and $p_{X^{\prime}}$ Section 2.1
$P_{X,X^{\prime}}^{-}$ probability measure with density $p_{X}\otimes p_{X^{\prime}}$ Section 4.2.2
$P_{X,Z_{d_{1}}}$ Borel probability measure in $\mathcal{X}\times\mathcal{Z}_{d_{1}}$ Section 5
Table 5: Notations of function classes.
Notation(s) Definition(s) Section
$\mathcal{G}_{0}$ a set of $[0,1]^{d_{1}}$-valued functions on $\mathcal{X}$ Section 2.3
$\mathcal{F},\mathcal{F}_{0}$ sets of $\Delta^{d}$-valued functions on $\mathcal{X}$ Section 3.1
$\mathcal{F}_{L,J,S,M,\bm{d}}^{\textup{NN}}$ ReLU networks class in (Nakada and Imaizumi, , 2020; Imaizumi and Fukumizu, , 2019, 2022) Section 3.5
$\mathcal{F}_{L,J,S,M,\bm{d}}^{\Delta^{d}\textup{-NN}}$ a set of $\Delta^{d}$-valued ReLU networks Section 3.5
$\sigma$ ReLU activation Section 3.5
$H$ softmax function Section 3.5
$\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$ localized subclass (Definition 10) Section 3.3
Table 6: Notations for algorithms.
Notation(s) Definition(s) Section
$\rho_{f}$ squared Euclidean distance Section 3.2
$\psi$ the function $\psi(s)=1-2D_{\Delta^{d}}^{-2}s$ Section 3.2
$\ell_{f}$ hinge loss defined in Definition 7 Section 3.2
$\widehat{g}_{n}^{\textup{LERM}}$ $(\beta,\varepsilon,P,n,\mathcal{F})$-local ERM (Definition 11) Section 3.4
$\widehat{g}_{n}^{\textup{GERM}}$ $(n,\mathcal{F})$-global ERM (Definition 18) Section 4.4
Funding
This work was partially supported by JSPS KAKENHI Grant Number 20H00576, 23H03460, and 24K14849. We also acknowledge partial support by JST BOOST, Japan Grant Number JPMJBS2417 and JPMJBS2430.

### v2

@@
 We next recall the classes of smooth boundaries introduced by Imaizumi and Fukumizu, (2022).
-For a given $\alpha>0$ and $d\in\mathbb{N}$, denote by $\mathcal{C}^{\alpha,d}$, the unit ball of the $\alpha$-Hölder space on $[0,1]^{d}$, namely
-$\mathcal{C}^{\alpha,d}=\left\{g:[0,1]^{d}\to\mathbb{R}\;|\;g\textup{ is }\alpha_{*}\textup{-times differentiable, }\|g\|_{\mathcal{C}^{\alpha,d}}\leq 1\right\},$
+For any $\alpha>0$, $R\geq 0$, and $d\in\mathbb{N}$, denote by $\mathcal{C}^{\alpha,d}$, the ball of the $\alpha$-Hölder space on $[0,1]^{d}$ with radius $R$, namely
+$\mathcal{C}^{\alpha,d}=\left\{g:[0,1]^{d}\to\mathbb{R}\;|\;g\textup{ is }\alpha_{*}\textup{-times differentiable, }\|g\|_{\mathcal{C}^{\alpha,d}}\leq R\right\},$
 where $\alpha_{*}$ is defined as
@@
   1. (P1)
-    Given $\mathscr{S}=\{\mathcal{J}_{i}\}_{i=1}^{d_{1}}$, there are disjoint subsets $\mathcal{I}_{1},\cdots,\mathcal{I}_{d_{1}}\subset\{1,-1\}^{d_{*}}$, functions $h_{1},\cdots,h_{d_{*}}\in\mathcal{C}^{\alpha,d_{0}-1}$, and indices $j_{1},\cdots,j_{d_{*}}\in\{1,\cdots,d_{0}\}$, such that for every $i=1,\cdots,d_{1}$,
+    Given $\mathscr{S}=\{\mathcal{J}_{i}\}_{i=1}^{d_{1}}$, there are disjoint partition $\{\mathcal{I}_{1},\cdots,\mathcal{I}_{d_{1}}\}$ of $\{1,-1\}^{d_{*}}$, functions $h_{1},\cdots,h_{d_{*}}\in\mathcal{C}^{\alpha,d_{0}-1}$, and indices $j_{1},\cdots,j_{d_{*}}\in\{1,\cdots,d_{0}\}$, such that for every $i=1,\cdots,d_{1}$,
     $\mathcal{J}_{i}=\bigcup_{(s_{1},\cdots,s_{d_{*}})\in\mathcal{I}_{i}}\bigcap_{k=1}^{d_{*}}\{x\in\mathcal{X}\;|\;s_{k}x_{j_{k}}\geq s_{k}h_{k}(x_{\setminus j_{k}})\},$
@@
   1. (P1’)
-    Given $\mathscr{S}^{\prime}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$, there are disjoint subsets $\mathcal{I}_{1},\cdots,\mathcal{I}_{d_{1}}\subset\{1,-1\}^{d_{*}}$, functions $h_{1},\cdots,h_{d_{*}}\in\mathcal{C}^{\alpha,d_{0}-1}$, and indices $j_{1},\cdots,j_{d_{*}}\in\{1,\cdots,d_{0}\}$ such that for every $i=1,\cdots,d_{1}$,
+    Given $\mathscr{S}^{\prime}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$, there are disjoint partition $\{\mathcal{I}_{1},\cdots,\mathcal{I}_{d_{1}}\}$ of $\{1,-1\}^{d_{*}}$, functions $h_{1},\cdots,h_{d_{*}}\in\mathcal{C}^{\alpha,d_{0}-1}$, and indices $j_{1},\cdots,j_{d_{*}}\in\{1,\cdots,d_{0}\}$ such that for every $i=1,\cdots,d_{1}$,
     $\mathcal{K}_{i}$ $=\bigcup_{(s_{1},\cdots,s_{d_{*}})\in\mathcal{I}_{i}}\bigcap_{k=1}^{d_{*}}\widetilde{\mathcal{K}}_{s_{k},h_{k},j_{k}},$
@@
 Meanwhile Parulekar et al., (2023) use a general equivalence relation using latent variables introduced by von Kügelgen et al., (2021).
+Definition of the map $\mathscr{S}_{P}$.
 Since any partition $\mathscr{S}$ in $\mathscr{P}_{\alpha}$ is expressed as a sequence, we need to introduce an equivalence relation that identifies the sequences containing the same subsets: namely, given two sequences $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ and $\{\mathcal{K}_{i}^{\prime}\}_{i=1}^{d_{1}}$ of subsets in $\mathcal{X}$, we define the equivalence relation $\sim$ as $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}\sim\{\mathcal{K}_{i}^{\prime}\}_{i=1}^{d_{1}}$ if and only if there is a permutation $\pi$ of the set $\{1,\cdots,d_{1}\}$ such that $\mathcal{K}_{\pi(i)}=\mathcal{K}_{i}^{\prime}$ for any $i\in\{1,\cdots,d_{1}\}$.
-Denote the equivalence class where the sequence $\mathscr{S}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ belongs to, by $[\mathscr{S}]$.
 Definition 2.
-Given $\alpha>0$ and $P\in\mathcal{P}_{\alpha}$, let $\mathscr{S}\in\mathscr{P}_{\alpha}$ be a partition satisfying the condition (A4) in Definition 1.
-Then, we define as $\mathscr{S}_{P}=[\mathscr{S}]$.
+Given $\alpha>0$, we define the map $P\mapsto\mathscr{S}_{P}$ on $\mathcal{P}_{\alpha}$ so that the following condition is satisfied for every $P\in\mathcal{P}_{\alpha}$: for any partition $\mathscr{S}\in\mathscr{P}_{\alpha}$ satisfying the condition (A4) for $P$, we have $\mathscr{S}_{P}\sim\mathscr{S}$.
 Note that by the condition (4), the map $P\mapsto\mathscr{S}_{P}$ on $\mathcal{P}_{\alpha}$ is well-defined.
-Also, given an equivalence class $[\mathscr{S}]\in\mathscr{P}_{\alpha}/\sim$, without loss of generality we can fix an arbitrary partition $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}\in[\mathscr{S}]$ in the proof of the main results of this work, by the definition of the equivalence relation.
+Also in the above definition, without loss of generality we can fix an arbitrary partition $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}\sim\mathscr{S}$ for each $P\in\mathcal{P}_{\alpha}$, by the definition of the equivalence relation.
 2.3 Formal Statement of the Problem
@@
 Given $\alpha>0$, let $P\in\mathcal{P}_{\alpha}$.
-Then, we aim at estimating the sets in the sequence $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}\in\mathscr{S}_{P}$.
+Then, we aim at estimating the sets in the sequence $\mathscr{S}_{P}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$.
 Similarly to Imaizumi and Fukumizu, (2019, 2022), we analyze the convergence rates of estimators in terms of the $L^{2}$-risk.
@@
 Problem.
-Given $P\in\mathcal{P}_{\alpha}$, is there an estimator $\widehat{g}_{n}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{G}_{0}$ such that the $L^{2}$-risk (5) defined with any sequence $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ in $\mathscr{S}_{P}$ tends to zero as $n\to\infty$?
+Given $P\in\mathcal{P}_{\alpha}$, is there an estimator $\widehat{g}_{n}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{G}_{0}$ such that the $L^{2}$-risk (5) defined with the sequence $\mathscr{S}_{P}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ tends to zero as $n\to\infty$?
 In particular:
   1. Q1.
-    Is it reasonable to use the pairwise dependency to estimate smooth boundaries?
+    Is it reasonable to use the pairwise dependence to estimate smooth boundaries?
   2. Q2.
@@
 Proposition 3.
-For every $\mathscr{S}\in\mathscr{P}_{\alpha,+}$, there is a Borel probability measure $P\in\mathcal{P}_{\alpha}$ such that we have $[\mathscr{S}]=\mathscr{S}_{P}$.
+For every $\mathscr{S}\in\mathscr{P}_{\alpha,+}$, there is a Borel probability measure $P\in\mathcal{P}_{\alpha}$ such that we have $\mathscr{S}\sim\mathscr{S}_{P}$.
 The proof is deferred to Appendix A.1.
@@
 Definition 5 (Contrastive function).
-Let $P\in\mathcal{P}$. The contrastive function $f^{*}:\mathcal{X}\to\Delta^{d}$ of $P$ is defined as
+Given $P\in\mathcal{P}$, let $\mathscr{S}_{P}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$. The contrastive function $f^{*}:\mathcal{X}\to\Delta^{d}$ of $P$ is defined as
 $f^{*}(x)=\sum_{i=1}^{d_{1}}\mathds{1}_{\mathcal{K}_{i}}(x)v_{i}.$
@@
 The function $\psi\circ\rho_{f}$ is a binary classifier, since $\psi$ controls the range of $\rho_{f}$.
-In fact, the classifier returns the prediction equal to $+1$ if the covariates $x$ and $x^{\prime}$ coincide.
+In fact, the classifier returns the prediction equal to $+1$ if the vectors $f(x)$ and $f(x^{\prime})$ coincide.
 In addition, by Theorem 3 in (Alexander, , 1977), the prediction equals $-1$ if and only if these points are embedded in different vertices on $\Delta^{d}$.
-Note that a similar classifier defined with the Mahalanobis distance is proposed in (Jin et al., , 2009).
+We give several comments on Definition 7.
+  • A similar classifier defined with the Mahalanobis distance is proposed in (Jin et al., , 2009).
 On the other hand, in Definition 7, we employ the Euclidean distance and additionally embed the variables $x$ and $x^{\prime}$ in the regular simplex.
@@
 Meanwhile, we use the statistical dependence to define the distribution of $(X,X^{\prime},Y)$, as in (Tsai et al., , 2020).
-Hinge loss is often studied in the literature on statistical learning (e.g., (Boser et al., , 1992; Lin, , 2002; Zhang, , 2004; Bartlett et al., , 2006; Lecué, , 2007; Kim et al., , 2021)).
+  • Hinge loss is often studied in the literature on statistical learning (e.g., (Boser et al., , 1992; Lin, , 2002; Zhang, , 2004; Bartlett et al., , 2006; Lecué, , 2007; Steinwart and Christmann, , 2008; Kim et al., , 2021)).
 In particular, in a setting of conventional binary classification, Kim et al., (2021, Theorem 3.2) study nonparametric estimation of smooth boundaries using empirical risk minimization with hinge loss.
@@
   • This proposition slightly generalizes some part of Theorem 3.8 in (Awasthi et al., , 2022) for hinge loss since we investigate in a more general setting.
-  • Note that the relationship between regular simplices and empirical risk minimizers in representation learning is shown in (Liu et al., , 2021; Lee et al., , 2024; Koromilas et al., , 2024).
-On the other hand, in the proposition above we are mainly interested in the population risk minimizers.
-  • Zhou et al., (2024, Theorem 1) show that the minimizer of the population risk with respect to a hinge loss defined in a similarity learning setting embeds covariates in the probability simplex.
+  • The relationship between regular simplices and empirical risk minimizers in representation learning is shown in (Liu et al., , 2021; Lee et al., , 2024; Koromilas et al., , 2024).
+Meanwhile, in the proposition above we are interested in the population risk minimizers.
+  • Zhou et al., (2024, Theorem 1) show that the vector-valued function minimizing the population risk with respect to a hinge loss defined in a similarity learning setting embeds covariates in the probability simplex.
 Since Proposition 8 deals with a different problem setting, the assumptions used in the present work are distinct from the ones considered in (Zhou et al., , 2024).
@@
 Let $\beta>0$, $\beta_{0}\geq 0$, and $P\in\mathcal{P}$.
+Denote by $f^{*}$, the contrastive function of $P$.
 Then, the $(\beta,\beta_{0},P)$-localized subclass of a set $\mathcal{F}\subset\mathcal{F}_{0}$, which is denoted by $\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$, is defined as
@@
 In other words, the localized subclass excludes functions that map $x\in\mathcal{K}_{i}$ to some wrong vertex $v_{j}$ with non-small probability.
-Note that some related concepts are introduced in (Schiebinger et al., , 2015; Trillos et al., , 2021) in the context of clustering theory.
+In several different contexts, some similar conditions are studied:
+  • In the theory of clustering, some related concepts are introduced in (Schiebinger et al., , 2015; Trillos et al., , 2021).
 Indeed, if $f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$ for some $\mathcal{F}\subseteq\mathcal{F}_{0}$, $\beta>0$, $\beta_{0}\geq 0$, and $P\in\mathcal{P}$, then it holds that
@@
 Lemma 14.
-Let $P\in\mathcal{P}$, and let $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ be any sequence in $\mathscr{S}_{P}$.
+Let $P\in\mathcal{P}$, and let $\mathscr{S}_{P}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$.
 Denote by $f^{*}$, the contrastive function of $P$.
@@
 Let $\beta\in(0,D_{{\textup{proj}}})$, $\beta_{0}\geq 0$, $n\in\mathbb{N}\setminus\{1\}$, $\mathcal{F}\subset\mathcal{F}_{0}$, and $P\in\mathcal{P}$.
-Let $f^{*}$ denote the contrastive function of $P$, and let $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ be any sequence in $\mathscr{S}_{P}$.
+Let $f^{*}$ denote the contrastive function of $P$, and let $\mathscr{S}_{P}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$.
 Then, there exists a constant $C_{2}>0$ such that for any $f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$ we have
@@
 Let $f^{*}$ be the contrastive function of $P$.
-Let $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ be any sequence in $\mathscr{S}_{P}$.
+Let $\mathscr{S}_{P}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$.
 For every $i,j\in\{1,\cdots,d_{1}\}$ such that $i\neq j$, any $r\in(0,2^{-1}\beta]$, and any $f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$, there exists a constant $C_{i,j}>0$ independent of $r$ such that
@@
 Some remarks on the technical approach.
-In the proof of Theorem 12, we show an inequality between two quantities, namely $\|f-f^{*}\|_{\mathcal{X},P_{X},2}$ and $\mathbb{E}_{P_{X,X^{\prime}}}[|\psi\circ\rho_{f}-\psi\circ\rho_{f^{*}}|]$.
+In the proof of Theorem 12, we show an inequality between $\|f-f^{*}\|_{\mathcal{X},P_{X},2}^{2}$ and $\mathbb{E}_{P_{X,X^{\prime}}}[|\psi\circ\rho_{f}-\psi\circ\rho_{f^{*}}|]$.
 A similar inequality for a contrastive learning problem has recently been shown in (Ge et al., , 2024).
-Specifically, Ge et al., (2024) consider the setting where the paired covariates $(X,X^{\prime})$ are independent of each other to show an inequality in (Ge et al., , 2024, Lemma D.2).
-On the other hand, we need a largely different approach to show our inequality since we consider the problem setting where the covariates $X$ and $X^{\prime}$ are not necessarily independent, following (Tsai et al., , 2020).
-Since the independence assumption is used in (Ge et al., , 2024, Lemma D.2), our inequality is not shown by their approach.
-Furthermore, in the field of contrastive learning, it is common to assume that $X$ and $X^{\prime}$ can be dependent, both in practice and in theory (see, e.g, (Arora et al., , 2019; Chen et al., , 2020; HaoChen et al., , 2021)).
-Therefore, Theorem 12 also gives some insights of how we can address the technical difficulty, which is described with localization of function classes as shown in Section 4.2.2.
+Specifically, in Lemma D.2 of (Ge et al., , 2024), they prove an upper bound of $\|f-f^{\prime}\|_{\mathcal{X},P_{X},2}^{2}$ based on the assumption that $\mathbb{E}_{P_{X}}[f(x)f^{\prime}(x)^{\top}]$ is a symmetric matrix for some given vector-valued functions $f,f^{\prime}$.
+However, this is a strong condition in the setting where only pairwise binary classification data is observed.
+Also, Ge et al., (2024) consider the setting where the paired covariates $X,X^{\prime}$ are independent.
+On the other hand, in our setting, the covariates $X$ and $X^{\prime}$ are not necessarily independent, following (Tsai et al., , 2020).
+Note that, in the field of contrastive learning, it is common to assume that $X$ and $X^{\prime}$ can be dependent, both in practice and in theory (see, e.g, (Arora et al., , 2019; Chen et al., , 2020; HaoChen et al., , 2021)).
+Therefore, Theorem 12 also gives some insights of how we can address the technical difficulties, which is described with localization of function classes as shown in Section 4.2.2.
 4.4 An Extended Analysis for the Global ERM
@@
   2. (B2)
-    There is some $P\in\mathcal{P}_{\alpha}$ such that the marginal distributions of $P$ and $P_{X,Z_{d_{1}}}$ with respect to $\mathcal{X}$ coincide with each other, and for any partition $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ in $\mathscr{S}_{P}$, any Bayes classifier $z^{*}$ of $P_{X,Z_{d_{1}}}$ satisfies $D_{P_{X}}((z^{*})^{-1}(i),\mathcal{K}_{i})=0$.
+    There is some $P\in\mathcal{P}_{\alpha}$ such that the marginal distributions of $P$ and $P_{X,Z_{d_{1}}}$ with respect to $\mathcal{X}$ coincide with each other, and for the partition $\mathscr{S}_{P}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$, any Bayes classifier $z^{*}$ of $P_{X,Z_{d_{1}}}$ satisfies $D_{P_{X}}((z^{*})^{-1}(i),\mathcal{K}_{i})=0$.
 Then, we have the following inequality between different risk functions.
@@
 Therefore, the distribution of $(X,X^{\prime},Y)$, denoted by $P$, belongs to $\mathcal{P}_{\alpha}$.
-By the definition, we have $[\mathscr{S}]=\mathscr{S}_{P}$.
+By the definition, we have $\mathscr{S}\sim\mathscr{S}_{P}$.
 ∎
@@
 Lemma 37 (Lemma 3.4 in Petersen and Voigtlaender, (2018)).
-Let $\alpha>0$, and let $\mathcal{C}^{\alpha}([-2^{-1},2^{-1}]^{d_{0}-1})$ be the unit ball of Hölder space defined on the domain $[-2^{-1},2^{-1}]^{d_{0}-1}$.
+Let $\alpha>0$, $R\geq 0$, and let $\mathcal{C}^{\alpha}([-2^{-1},2^{-1}]^{d_{0}-1})$ be the ball of Hölder space defined on the domain $[-2^{-1},2^{-1}]^{d_{0}-1}$ with radius $R$.
 Let $\mathcal{K}\subset[-2^{-1},2^{-1}]^{d_{0}}$ be any set such that there are function $h\in\mathcal{C}^{\alpha}([-2^{-1},2^{-1}]^{d_{0}-1})$ and index $i\in\{1,\cdots,d_{0}\}$ satisfying
@@
 Let $0<\varepsilon<2^{-1}$.
-Let $P\in\mathcal{P}_{\alpha}$, and let $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ be any sequence in $\mathscr{S}_{P}$.
+Let $P\in\mathcal{P}_{\alpha}$, and let $\mathscr{S}_{P}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$.
 Denote by $g_{i}^{*}=\mathds{1}_{\mathcal{K}_{i}}$, $i=1,\cdots,d_{1}$.
@@
 Given a compact subset $\mathcal{A}$ in a measure space equipped with the Lebesgue measure, let $\mathscr{N}(\delta,\mathcal{H},\|\cdot\|_{L^{\infty}(\mathcal{A})})$ denote the covering number of a class $\mathcal{H}$ of real-valued functions with respect to the pseudo-distance $\|\cdot\|_{L^{\infty}(\mathcal{A})}$ (see e.g., (Steinwart and Christmann, , 2008, Definition 6.19) for covering numbers).
-Besides, given a map $h:\Delta^{d}\to\mathcal{A}$ for a set $\mathcal{A}$, denote by $h\circ\mathcal{F}=\{h\circ f\;|\;f\in\mathcal{F}\}$.
+Also, for any $\mathcal{F}\subset\mathcal{F}_{0}$, denote by $\rho(\mathcal{F})=\{\rho_{f}\;|\;f\in\mathcal{F}\}$.
 Then, the following evaluation holds.
@@
 Then, for any $\delta\geq 0$, there exists a constant $C_{6}>0$ independent of $n,\delta$ such that we have
-$\log\mathscr{N}(\delta,\psi\circ\rho\circ\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
+$\log\mathscr{N}(\delta,\psi\circ\rho(\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
 $\leq d_{1}^{2}\log\mathscr{N}(C_{6}\delta,\mathcal{F}_{L,J,S,M,\bm{d}_{1}}^{\textup{NN}},\|\cdot\|_{L^{\infty}(\mathcal{X})}),$
@@
 Note that $\psi$ is Lipschitz continuous. Thus, we have
-$\mathscr{N}(\delta,\psi\circ\rho\circ\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
-$\leq\mathscr{N}(2^{-1}D_{\Delta^{d}}^{2}\delta,\rho\circ\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})}).$ (50)
+$\mathscr{N}(\delta,\psi\circ\rho(\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
+$\leq\mathscr{N}(2^{-1}D_{\Delta^{d}}^{2}\delta,\rho(\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})}).$ (50)
 Then, we note that
-$\mathscr{N}(2^{-1}D_{\Delta^{d}}^{2}\delta,\rho\circ\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
+$\mathscr{N}(2^{-1}D_{\Delta^{d}}^{2}\delta,\rho(\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
 $=\mathscr{N}(2^{-1}D_{\Delta^{d}}^{2}\delta,\{(x,x^{\prime})\mapsto\|f(x)-f(x^{\prime})\|_{2}^{2}\;|\;f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})\},\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})}).$ (51)
@@
 where $c_{6}^{\prime}=2^{-1}d_{1}^{-1}c_{6}$. Combining (A.7.4), (A.7.4), (A.7.4), and (A.7.4), we have
-$\mathscr{N}(\delta,\psi\circ\rho\circ\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
+$\mathscr{N}(\delta,\psi\circ\rho(\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
 $\leq\prod_{j=1}^{d_{1}}\mathscr{N}(L_{2}\delta,\{x\mapsto c_{j}(f,x)\;|\;f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})\},\|\cdot\|_{L^{\infty}(\mathcal{X})}),$ (54)
@@
 Thus, setting $C_{6}=2^{-1}d_{1}e^{-4M}L_{2}$, by (A.7.4) and (A.7.4), we obtain
-$\log\mathscr{N}(\delta,\psi\circ\rho\circ\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
+$\log\mathscr{N}(\delta,\psi\circ\rho(\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F}^{\prime})),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
 $\leq d_{1}^{2}\log\mathscr{N}(C_{6}\delta,\mathcal{F}_{L,J,S,M,\bm{d}_{1}}^{\textup{NN}},\|\cdot\|_{L^{\infty}(\mathcal{X})}),$ (56)
@@
 Denote by $\mathcal{F}_{\textup{local}}^{*}=\mathscr{F}_{\beta,\beta^{-1}\varepsilon_{n},P}(\mathcal{F}^{*})$.
-Then, the set $\psi\circ\rho\circ\mathcal{F}_{\textup{local}}^{*}$ is $1$-uniformly bounded.
+Then, the set $\psi\circ\rho(\mathcal{F}_{\textup{local}}^{*})$ is $1$-uniformly bounded.
 Also, it is well known that hinge loss is $1$-Lipschitz continuous in the first argument (see, e.g., (Steinwart and Christmann, , 2008, Example 2.27)).
@@
 By (van de Geer, , 2000, Lemma 2.1) and (Kim et al., , 2021, (A.1)), for any $s>0$, we have
-$\log\mathscr{B}(s,\psi\circ\rho\circ\mathcal{F}_{\textup{local}}^{*},\|\cdot\|_{L^{2}(\mathcal{X}^{2})})\leq\log\mathscr{N}(\frac{s}{2},\psi\circ\rho\circ\mathcal{F}_{\textup{local}}^{*},\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})}).$ (63)
+$\log\mathscr{B}(s,\psi\circ\rho(\mathcal{F}_{\textup{local}}^{*}),\|\cdot\|_{L^{2}(\mathcal{X}^{2})})\leq\log\mathscr{N}(\frac{s}{2},\psi\circ\rho(\mathcal{F}_{\textup{local}}^{*}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})}).$ (63)
 By Lemma 41 and Lemma 42, for the constant $C_{6}=2^{-4}dd_{1}^{-1}e^{-4M^{*}}D_{\Delta^{d}}^{2}$, we have
-$\log\mathscr{N}(2^{-1}s,\psi\circ\rho\circ\mathcal{F}_{\textup{local}}^{*},\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
+$\log\mathscr{N}(2^{-1}s,\psi\circ\rho(\mathcal{F}_{\textup{local}}^{*}),\|\cdot\|_{L^{\infty}(\mathcal{X}^{2})})$
 $\leq d_{1}^{2}\log\mathscr{N}(2^{-1}C_{6}s,\mathcal{F}_{L^{*},J^{*},S^{*},M^{*},\bm{d}_{1}^{*}}^{\textup{NN}},\|\cdot\|_{L^{\infty}(\mathcal{X})})$